### v1

# Theory and Interpretability of Quantum Extreme Learning Machines: a Pauli-transfer matrix approach

Abstract
Quantum reservoir computers (QRCs) have emerged as a promising approach to quantum machine learning, since they utilize the natural dynamics of quantum systems for data processing and are simple to train. Here, we consider $n$-qubit quantum extreme learning machines (QELMs) with continuous-time reservoir dynamics. QELMs are memoryless QRCs capable of various ML tasks, including image classification and time series forecasting. We apply the Pauli transfer matrix (PTM) formalism to theoretically analyze the influence of encoding, reservoir dynamics, and measurement operations, including temporal multiplexing, on the QELM performance. This formalism makes explicit that the encoding determines the complete set of (nonlinear) features available to the QELM, while the quantum channels linearly transform these features before they are probed by the chosen measurement operators. Optimizing a QELM can therefore be cast as a decoding problem in which one shapes the channel-induced transformations such that task-relevant features become available to the regressor. The PTM formalism allows one to identify the classical representation of a QELM and thereby guide its design towards a given training objective. As a specific application, we focus on learning nonlinear dynamical systems and show that a QELM trained on such trajectories learns a surrogate-approximation to the underlying flow map.
I Introduction
QRCs have enjoyed significant attention since their introduction [1], featuring applications in various domains, such as image classification [2, 3, 4, 5], time series forecasting [1, 6, 7, 8, 9, 10], fluid dynamics [11, 12], or post-processing of quantum experiments [13, 14, 15, 16]. A major advantage of QRCs is their suitability for noisy intermediate-scale quantum (NISQ) hardware, primarily because the training process is restricted to the classical readout layer, thereby rendering the training problem convex and avoiding the barren plateau problem associated with variational quantum algorithms [17]. QRCs are often resilient to, and can even benefit from, environmental noise [18, 19, 20, 21]. The relevance of nonlinear input transformations has been analyzed for QRCs in [22, 23, 24], and universality conditions have been derived in [25, 26, 27]. Notably, QRCs are closely related to quantum kernel methods [28]—a fact that can be utilized to find optimal measurement operators [29] and harness insights from the classical domain [30, 31, 32]. For general overviews on QRCs, we refer to [33, 34].
A specific QRC variant without any coherently evolving internal state is the QELM [35, 33, 34, 2]. Its classical counterpart, the ‘extreme learning machine’, is essentially a feedforward network with fixed internal architecture [36], similarly to a random-feature model [37]. A QELM differs from a conventional quantum neural network with fixed internal parameters in two important ways: first, instead of gate-based architectures, it typically employs the natural dynamics of a physical quantum system (a reservoir). Second, instead of optimizing a measurement output for a certain task, one collects multiple unspecific measurements into a large readout vector and feeds it into a classical regression or classification algorithm. The main task in optimizing QELMs thus consists in tuning the quantum properties of the reservoir to maximize information processing and being able to properly decode relevant features from the measurements [38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 14]. Identifying possible quantum advantages, e.g., related to the exponentially large Hilbert space into which data is encoded, is an active research topic [48, 34, 49, 41, 3, 45]. However, the exponential concentration effect emerging for large qubit numbers is a phenomenon closely related to barren plateaus [17, 50] that can strongly reduce QELM performance [51].
The present work aims at a interpretable framework that can explain why and which failures in QELMs occur and thereby support systematic design of QELMs. To this end, we will also analyze well-known optimization tools like temporal multiplexing [1] or information processing capacity [52, 53] in the light of the PTM formalism. We use chaotic dynamical systems as benchmarks since they have a known ground-truth model and are in principle discoverable from training data [54]. Our findings also hold implications for general QML models.
II Model
We briefly review the formalism and specific implementation of a QELM used in this work [1, 33, 34]. Let $\mathcal{M}(\mathcal{H})$ denote the space of $d^{2}$-dimensional Hermitian matrices on the $d$-dimensional Hilbert space $\mathcal{H}$, where $d=2^{n}$ in terms of the number of qubits $n$. Given classical inputs $\mathbf{x}\in\mathbb{R}^{D}$, a QELM is defined by the output function
$f(\mathbf{x})=\sum_{k=1}^{m}w_{k}\mathrm{tr}[M_{k}\rho(\mathbf{x})]$ (1)
where $\{M_{k}\}_{k=1}^{m}$ are a set of measurement operators in $\mathcal{M}(\mathcal{H})$, and $\rho(\mathbf{x})\in\mathcal{M}(\mathcal{H})$ is the density matrix of the quantum system, obtained after applying encoding and other quantum channels to an initial state, $\rho(\mathbf{x})=\mathcal{E}(\cdots\mathcal{E}^{\text{enc}}_{\mathbf{x}}(\rho_{0})\cdots)$. The trainable weights $w_{k}$ are determined by minimizing the empirical risk:
$\min_{\mathbf{w}\in\mathbb{R}^{m}}\sum_{i=1}^{P}\mathcal{L}\left(f(\mathbf{x}_{i}),\,y_{i}\right)+\frac{\lambda}{2}\|\mathbf{w}\|^{2}_{2}$ (2)
[…]
$\mathbf{w}^{*}=(\boldsymbol{\Psi}\boldsymbol{\Psi}^{\top}+\lambda\mathbbm{1}_{m})^{-1}\boldsymbol{\Psi}\mathbf{y},\qquad\Psi_{ki}=\mathrm{tr}(M_{k}\rho(\mathbf{x}_{i}))$ (3)
where $\mathbf{y}\in\mathbb{R}^{P}$ is the sample vector and $\boldsymbol{\Psi}\in\mathbb{R}^{{m}\times P}$ is the readout design matrix.
When optimizing over a complete basis of measurement operators $\{M_{k}\}_{k=1}^{d}$, the QELM takes the form of a quantum kernel model [28, 29] (see Appendix A for a summary). In the present work, we do not use that formalism, but instead work with a preselected observable set and optimize only over the associated weights $w_{k}$.
II.1 Feature map
We consider pure states, i.e., $\rho(\mathbf{x})=|\psi(\mathbf{x})\rangle\langle\psi(\mathbf{x})|$, with the quantum state $|\psi(\mathbf{x})\rangle\in\mathcal{H}$ given by
$|\psi(\mathbf{x})\rangle=U_{R}(t)S(\mathbf{x})|\Phi_{0}\rangle.$ (4)
Here, $S(\mathbf{x})$ is a data encoding unitary, $U_{R}$ is a reservoir unitary, and the initial state $|\Phi_{0}\rangle$ is taken to be the ground state $|\Phi_{0}\rangle=|0\rangle^{\otimes n}$. The unitary $U_{R}(t)=\exp(-\mathrm{i}tH_{R})$ is defined by time evolution under a reservoir Hamiltonian $H_{R}$. Specifically, we consider Haar-random Hamiltonians $H_{R}$ as well as two variants of a transverse field Ising model (TFIM)
$H_{\text{TFIM}}^{\text{zz-x}}$ $=-J\sum_{j=1}^{n-1}Z_{j}Z_{j+1}-h\sum_{j=1}^{n}X_{j},$ (5a)
$H_{\text{TFIM}}^{\text{xx-z}}$ $=-J\sum_{j=1}^{n-1}X_{j}X_{j+1}-h\sum_{j=1}^{n}Z_{j},$ (5b)
with couplings $J,h\in\mathbb{R}$. Here and in the following, $\{X_{j},Y_{j},Z_{j}\}$ denote single-qubit Pauli operators acting on the $j$-th qubit and $Z_{j}Z_{j+1}$ is a shorthand for the usual $n$-qubit tensor product (with identity operators acting on the remaining qubits). The two TFIM variants in Eq. 5 are related by a global Hadamard transform $U_{\text{H}}=\mathrm{H}^{\otimes n}$:
$U_{\text{H}}X_{j}U_{\text{H}}^{\dagger}=Z_{j},\quad U_{\text{H}}Y_{j}U_{\text{H}}^{\dagger}=-Y_{j},\quad U_{\text{H}}Z_{j}U_{\text{H}}^{\dagger}=X_{j},\quad U_{\text{H}}H_{\text{TFIM}}^{\text{zz-x}}U_{\text{H}}^{\dagger}=H_{\text{TFIM}}^{\text{xx-z}}.$ (6)
II.2 Encoding schemes
We consider (parallel) product state encodings with $D=n$, producing the state
$\rho_{\text{enc}}(\mathbf{x})=S(\mathbf{x})\rho_{0}S(\mathbf{x})^{\dagger}=\bigotimes_{j=1}^{n}|\psi_{\text{enc}}(x_{j})\rangle\langle\psi_{\text{enc}}(x_{j})|,\qquad S(\mathbf{x})=\bigotimes_{j=1}^{n}S_{j}(x_{j})$ (7)
where the unitary $S_{j}$ act on the $j$-th qubit. Specifically, in our numerical experiments we mostly work with amplitude encoding [55]. This scheme comes in two variants, which lead to different expressivities: one encodes data on half a great circle of the Bloch sphere
$|\psi_{\text{enc}}(x)\rangle=\sqrt{x}|0\rangle+\sqrt{1-x}|1\rangle,\qquad x\in[0,1],$ (8)
and one on the full great circle
$|\psi_{\text{enc}}(x)\rangle=x|0\rangle+\sqrt{1-x^{2}}|1\rangle,\qquad x\in[-1,1].$ (9)
In a few cases, we also consider rotational encoding [55], defined by the unitary
$S_{j}(x_{j})=\exp(-i\pi x_{j}\sigma^{p_{j}})=R_{p_{j}}(2\pi x),\qquad x\in[0,1],$ (10)
where $\sigma^{p}\in\{\mathbbm{1},X,Y,Z\}$ are single-qubit Pauli matrices, $p\in\{0,x,y,z\}$, and $R_{p_{j}}$ denotes rotation gates. In order to avoid irrelevant phases or trivial states, one typically uses $R_{Y}$ gates, which yield the single-qubit state
$|\psi_{\text{enc}}(x_{j})\rangle=R_{Y}(2\pi x_{j})|0\rangle=\cos(\pi x_{j})|0\rangle+\sin(\pi x_{j})|1\rangle.$ (11)
III Theory
QML models, including and QELMs, have often been analyzed from the Fourier perspective, which is natural for Hamiltonian-based encoding [48, 51, 49]. In the following, we instead employ the Pauli transfer matrix (PTM) formalism, which allows a more direct identification of the relevant features [33, 43, 56, 57, 58] and has been previously used for the analysis of QRCs [33, 25, 46].
III.1 Pauli-transfer formulation of a QELM
In the PTM formalism, we expand states and operators in the $n$-qubit Pauli basis
$\{P_{k}=\sigma_{a_{1}}\otimes\cdots\otimes\sigma_{a_{n}}:\ a_{i}\in\{0,x,y,z\}\}_{k=0}^{d^{2}-1},$ (12)
which consists of $d^{2}$ Pauli-strings $P_{k}$, with $P_{0}=\mathbb{I}$ and $d=2^{n}$, orthonormalized by $\mathrm{tr}(P_{k}P_{l})=d\delta_{kl}$. Every state admits the expansion
$\rho(\mathbf{x})=\frac{1}{2^{n}}\sum_{k=0}^{d^{2}-1}\phi_{k}(\mathbf{x})P_{k},\qquad\phi_{k}(\mathbf{x})=\mathrm{tr}(P_{k}\rho(\mathbf{x})),$ (13)
giving rise to a feature vector $\boldsymbol{\upphi}(\mathbf{x})\in\mathbb{R}^{d^{2}}$. For product state encodings, the components of $\phi_{k}(x)$ correspond to all possible $d^{2}$ monomial combinations of the single-Pauli features (see Section III.2). Any quantum channel $\mathcal{E}$ acts linearly on this vector via the Pauli transfer matrix:
$(T_{\mathcal{E}})_{kl}=\frac{1}{d}\,\mathrm{tr}\big(P_{k}\,\mathcal{E}(P_{l})\big),$ (14)
transforming the features as $\boldsymbol{\upphi}^{\prime}=T_{\mathcal{E}}\,\boldsymbol{\upphi}$. If the channel $\mathcal{E}$ is completely positive (CPTP), it admits a Kraus representation
$\mathcal{E}(\rho)=\sum_{\ell=1}^{\mu}K_{\ell}\rho K_{\ell}^{\dagger}$ (15)
with $\sum_{\ell=1}^{\mu}K_{\ell}^{\dagger}K_{\ell}=\mathbbm{1}$ and $\mu\leq d^{2}$ (dependent on the channel, e.g., $\mu=1$ for a unitary channel, see Appendix D for further discussion).
Consider the QELM in Eq. 1 with a subset of measured Pauli operators $\{P_{k}\}_{k=1}^{{m}}$, ${m}\leq d^{2}$. This selection can be represented by a matrix $\mathcal{S}\in\mathbb{R}^{{m}\times d^{2}}$ (a subset of rows of the identity matrix) that extracts the corresponding rows from $T_{\mathcal{E}}$. The readout vector $\mathbf{F}(\mathbf{x})\in\mathbb{R}^{{m}}$ contains their expectation values $F_{k}(\mathbf{x})=\mathrm{tr}\bigl[P_{k}\,\mathcal{E}(\rho_{\text{enc}}(\mathbf{x}))\bigr]$. Expanding $\rho_{\text{enc}}$ as in Eq. 13 and using Eq. 14 gives the model output function
$f(\mathbf{x})=\mathbf{w}^{\top}\,\mathbf{F}(\mathbf{x})=\mathbf{w}^{\top}\,T_{\mathcal{E}}^{({m})}\,\boldsymbol{\upphi}(\mathbf{x})=\sum_{k=1}^{m}w_{k}\sum_{j=0}^{d^{2}-1}(\mathcal{S}T_{\mathcal{E}})_{kj}\phi_{j}(\mathbf{x}),$ (16)
where $\mathbf{w}\in\mathbb{R}^{{m}}$ are classical weights and $T_{\mathcal{E}}^{({m})}\equiv\mathcal{S}T_{\mathcal{E}}$ is the ${m}\times d^{2}$ block of $T_{\mathcal{E}}$ restricted to the measured observables. This formulation explicitly shows that $f(\mathbf{x})$ is a linear function of the features $\boldsymbol{\upphi}(\mathbf{x})$ generated solely by the encoding and mixed by the reservoir channel $T_{\mathcal{E}}$.
With the Kraus representation (15), we can write
$(T_{\mathcal{E}}^{({m})})_{kj}=\frac{1}{d}\sum_{\ell=1}^{\mu}\mathrm{tr}\bigl(P_{k}\,K_{\ell}P_{j}K_{\ell}^{\dagger}\bigr)=\mathrm{tr}\bigl(P_{k}^{(H)}P_{j}\bigr)$ (17)
where the last expression introduces the generalized Heisenberg representation of an observable $O$,
$O^{\text{(H)}}=\frac{1}{d}\sum_{\ell=1}^{\mu}K_{\ell}^{\dagger}OK_{\ell}.$ (18)
Specifically, for the unitary channel, we have $K=U(t)$ ($\mu=1$), such that $T_{\mathcal{E}}$ becomes an orthogonal matrix $V(t)\in\mathrm{SO}(d^{2})$ with elements
$V_{kj}(t)=\frac{1}{d}\mathrm{tr}\bigl[P_{k}\,U(t)P_{j}U(t)^{\dagger}\bigr],$ (19)
and $V_{kj}(0)=\delta_{kj}$. For general CPTP maps, $T_{\mathcal{E}}$ can also induce contraction and translation of the feature vector (see Section D.2).
III.2 Feature encoding
The dependence of the feature vector $\boldsymbol{\upphi}(\mathbf{x})$ on the encoding can be made explicit for product-state encodings. Using Eq. 7 and an initial product state, the encoded state before the reservoir unitary is
$\rho_{\text{enc}}(\mathbf{x})=S(\mathbf{x})\,|0\rangle\langle 0|^{\otimes n}\,S(\mathbf{x})^{\dagger}=\bigotimes_{j=1}^{n}\rho_{j}(x_{j}),\qquad\rho_{j}(x_{j})=S_{j}(x_{j})|0\rangle\langle 0|S_{j}(x_{j})^{\dagger}.$ (20)
It is convenient to introduce the single-qubit Pauli features
$\boldsymbol{\upphi}^{(j)}(x_{j})=\bigl(1,\phi_{x}^{(j)}(x_{j}),\phi_{y}^{(j)}(x_{j}),\phi_{z}^{(j)}(x_{j})\bigr)^{\top},\qquad\phi_{a}^{(j)}(x_{j})=\mathrm{tr}\bigl[\sigma^{a}\,\rho_{j}(x_{j})\bigr],$ (21)
with $\phi_{0}^{(j)}=\mathrm{tr}(\rho_{j})=1$ (normalization), $\sigma^{0}\equiv\mathbbm{1}$ and $\sigma^{x,y,z}\equiv X,Y,Z$ as before. The three non-trivial components of $\boldsymbol{\upphi}^{(j)}$ define the Bloch vector $\boldsymbol{\varphi}^{(j)}=\{\phi_{x}^{(j)},\phi_{y}^{(j)},\phi_{z}^{(j)}\}$, which fulfills $\|\boldsymbol{\varphi}^{(j)}\|^{2}=1$ for pure states (mainly considered here) and $\|\boldsymbol{\varphi}^{(j)}\|^{2}\leq 1$ for mixed states. Since both the state and the Pauli basis factorize across qubits, the $n$-qubit feature vector is simply
$\boldsymbol{\upphi}(\mathbf{x})=\bigotimes_{j=1}^{n}\boldsymbol{\upphi}^{(j)}(x_{j}),$ (22)
and its components are given by
$\phi_{a_{1}\ldots a_{n}}(\mathbf{x})=\mathrm{tr}\!\Bigl[\Bigl(\bigotimes_{j=1}^{n}\sigma^{a_{j}}\Bigr)\,\rho_{\text{enc}}(\mathbf{x})\Bigr]=\prod_{j=1}^{n}\phi_{a_{j}}^{(j)}(x_{j}),\qquad a_{j}\in\{0,x,y,z\}.$ (23)
The subscript $a_{1}\ldots a_{n}$ corresponds to the explicit site representation of the index $k$ in Eq. 13. The $\phi_{k}$ thus run over all possible combinations of the single-qubit features in Eq. 21: for each qubit one chooses one of the four factors $\{1,\phi^{(j)}_{x},\phi^{(j)}_{y},\phi^{(j)}_{z}\}$, yielding a total of at most $d^{2}=4^{n}$ features. If the measured observables are Pauli strings of product form $P=\bigotimes_{j=1}^{n}\sigma^{a_{j}}$, the corresponding readout feature factorizes analogously as $\phi^{\prime}_{a_{1}\ldots a_{n}}(\mathbf{x})=\prod_{j=1}^{n}\bigl(\boldsymbol{\upphi}^{{}^{\prime}(j)}(x_{j})\bigr)_{a_{j}}$ with the post-reservoir feature vector $\boldsymbol{\upphi}^{\prime}$. The readout features resulting from some simple encoding schemes are summarized in Table 1. Effects of feature mixing by a unitary channel are discussed in the next section.
Feature $\phi_{a_{k}}^{(k)}(u_{k})$
Factor in observable $P$ Amplitude on $[0,1]$ Rotational
$\cdots\otimes X_{k}\otimes\cdots$ $2\sqrt{(1-u_{k})u_{k}}$ $\sin(\pi u_{k})$
$\cdots\otimes Y_{k}\otimes\cdots$ 0 0
$\cdots\otimes Z_{k}\otimes\cdots$ $2u_{k}-1$ $\cos(\pi u_{k})$
$\cdots\otimes\tfrac{1}{2}(\mathbb{I}+Z_{k})\otimes\cdots$ $u_{k}$ $\cos^{2}(\pi u_{k}/2)$
Table 1: Pauli–basis projections $\langle\psi_{\text{enc}}|P|\psi_{\text{enc}}\rangle$ of an encoded product state $\rho_{\text{enc}}(\mathbf{u})$ for rotational encoding (10) and amplitude encoding (8). Here $P=\bigotimes_{k=1}^{n}\sigma_{a_{k}}$ denotes a Pauli operator ($\sigma_{a_{k}}\in\{\mathbb{I},X,Y,Z\}$) and the classical datum entering qubit $k$ is $u_{k}$. Replacing $u_{k}\!\to\!u_{k}^{2}$ transforms the first column to the amplitude encoding variant (9). For convenience, we have included the projection operator $\frac{1}{2}(\mathbb{I}+Z_{k})$. One observes that each Pauli factor $\sigma_{a_{k}}$ (first column) in an observable $P$ produces the corresponding single-qubit feature $\phi_{a_{k}}^{(k)}(u_{k}):=\mathrm{tr}\!\big(\sigma_{a_{k}}\,\rho_{k}(u_{k})\big)$ (2nd and 3rd column). For instance, the $X_{k}$ row gives $\phi_{x}^{(k)}(u_{k})$ and the $Z_{k}$ row gives $\phi_{z}^{(k)}(u_{k})$.
IV PTM analysis for general quantum channels
For separable quantum channels, the PTM factorizes as $T=\bigotimes_{j=1}^{n}T_{j}$ into single-qubit PTMs $T_{j}$. The action of a local unitary is discussed in Section D.1. To understand and optimize a QELM with general nonlocal quantum channels, we must analyze the full PTM [Eqs. 16 and 17]. The optimization can be split into several subtasks that define the fundamental reconstruction problem in (quantum) reservoir computing [42, 46, 59, 14]:
  1. 1.
    Decide which types of features $\phi_{r}$ are required to learn the dataset. For example, learning the flow map of a Markovian dynamical system typically requires low-order polynomial features. For harmonic time series and non-Markovian data, linear delay features are needed to capture temporal dependencies. Results like in Table 1 can aid systematic reservoir design.
  2. 2.
    Select suitable measurement observables $P_{k}$. Ideally, the Heisenberg-evolved observables $P_{k}^{\text{(H)}}(t)$ have strong overlap with the observable $P_{r}$ generating the desired feature $\phi_{r}$ [see Eq. 17]. One possibility to achieve this is to determine an optimal measurement operator $M^{*}$ via primal or dual optimization, which can then be further expanded into a set of rotated Pauli-Z operators [29]. Another possibility is temporal multiplexing [1, 44, 60, 61], discussed further in the PTM framework below (see Section IV.2).
  3. 3.
    When the available measurement operations are restricted, one must tune the quantum channel $\mathcal{E}$ and the readout layer such that the desired features $\phi_{r}$ become decodable. Strict decodability means that, for each such $\phi_{r}$, there exists a weight vector $\mathbf{w}_{r}$ such that $\mathbf{w}_{r}T_{\mathcal{E}}\boldsymbol{\upphi}\propto\phi_{r}$ (see Section IV.3). In practice, decodability in the statistical sense (as a correlation) and within a certain error tolerance is often sufficient, since the features needed for typical training tasks can usually be approximated on limited domains from the available Pauli features (see Section IV.4). We will show that temporal multiplexing directly enhances decodability by increasing the operator support.
In temporal multiplexing [62, 1], instead of working at a fixed evolution time $t$ and a potentially large set of observables $P_{k}$, one selects a small subset $\mathcal{S}$ of size ${m}=|\mathcal{S}|<d^{2}$ and evaluate them at several times $t_{1},\ldots,t_{L}$. Correspondingly, we stack a subset $\mathcal{S}$ of rows of $V(t)$ into a ${m}{L}\times d^{2}$ observability matrix
$R=\begin{pmatrix}\mathcal{S}V(t_{1})\\ \mathcal{S}V(t_{2})\\ \vdots\\ \mathcal{S}V(t_{L}).\end{pmatrix}$ (24)
The case without temporal multiplexing corresponds to L=1{L}=1. For $\mathcal{S}=\mathbbm{1}_{d^{2}}$ and L=1{L}=1, one then recovers the full unital PTM $R=V(t)$.
IV.1 Classical representation
According to Eq. 16, a QELM has an equivalent classical representation as a PTM-weighted nonlinear function library,
$f(\mathbf{u})=\mathbf{w}^{\top}R\boldsymbol{\upphi}(\mathbf{u})=\mathbf{w}^{\top}\mathbf{F}(\mathbf{u})$ (25)
where $R\in\mathbb{R}^{{B}\times{q}}$ is the effective PTM (24) (including any row-selection and temporal multiplexing) and $F_{k}=R_{kj}\phi_{j}$ are the Pauli features available at the readout. We denote the effective number of features by ${q}\leq d^{2}$, excluding features identically vanishing due to the encoding, while B{B} is the total number of measurements, e.g. B=mL{B}={m}{L}.
If $R$ has full column rank, it is left-invertible  [^1] This allows for the existence of a pseudoinverse such that $R^{+}R=\mathbbm{1}_{q}$. and thus the regressor can always find weights $\mathbf{w}\propto(R^{+})_{r}$ that isolate any desired element $\phi_{r}$. Since one usually trains with arbitrary targets $\mathbf{y}$ instead of $\phi_{r}$, the actual weights [Eq. 3] contain projections from the Pauli feature space to the data space:
$\mathbf{w}^{*}=(GG^{\top}+\lambda\mathbbm{1}_{B})^{-1}G\mathbf{y}=R(CR^{\top}R+\lambda\mathbbm{1}_{q})^{-1}\Phi\mathbf{y},$ (26)
where we introduced the feature design and Gram matrices
$\Phi\equiv[\boldsymbol{\upphi}(\mathbf{u}_{1})\cdots\boldsymbol{\upphi}(\mathbf{u}_{P})]\in\mathbb{R}^{{q}\times P},\quad C=\Phi\Phi^{\top}\in\mathbb{R}^{{q}\times{q}},$ (27)
and the corresponding readout design and Gram matrices
$G\equiv[\mathbf{F}(\mathbf{u}_{1})\cdots\mathbf{F}(\mathbf{u}_{P})]=R\Phi\in\mathbb{R}^{{B}\times P},\quad GG^{\top}=RCR^{\top}\in\mathbb{R}^{{B}\times{B}}$ (28)
The latter are typically computed in a QELM implementation, whereas $R$ and $\Phi$ are theoretical objects. We will analyze in the following subsections the requirements on the encoding and reservoir unitary to improve trainability and avoid expensive hyperparameter search.
The last equality in (26) (which follows from the Woodbury identity) implies that for the full unitary PTM $R=V$ [Eq. 19], the weights are simply rotated versions of the ones without PTM mixing:
$\mathbf{w}^{*}=V(C+\lambda\mathbbm{1}_{q})^{-1}\Phi\mathbf{y}.$ (29)
Moreover, since $V$ is orthogonal, it drops out from the predictor (25), consistent with the absence of a reservoir unitary in the kernel-representation (63). Such a QELM is then simply a regressor acting on the full $d^{2}$-dimensional feature library $\boldsymbol{\upphi}$. In the context of time series processing (especially when using some inputs $u_{j}$ to encode delay states) this is also known as a “next-generation” reservoir computer (NG-RC) [64] or, equivalently, a nonlinear vector autoregressive process (NVAR). When $R$ merely has full column rank, it can be interpreted to induce a regressor operating on the mixed feature set $\boldsymbol{\upphi}^{\prime}=R\boldsymbol{\upphi}$.
A coarse measure for expressivity is the rank of the readout Gram matrix in Eq. 26, which can be bounded as
$\mathrm{rank}(G)=\min(\mathrm{rank}(R),\mathrm{rank}(\Phi))\leq\min({B},P,{q}).$ (30)
For sufficiently diverse data and large sample number $P$, $\Phi$ typically achieves a rank near its maximum $\min({q},P)$. However, it can suffer from severe ill-conditioning since the features $\phi_{k}$ are in general not orthogonal. Anticipating results discussed below (see Fig. 4), for random Hamiltonians $\mathrm{rank}(R)$ typically grows monotonically with the measurement budget B{B} and saturates at $\mathrm{rank}(\Phi)\sim{q}$. By contrast, for structured Hamiltonians like the TFIM, $\mathrm{rank}(R)$ and thus $\mathrm{rank}(G)$ can saturate at a value significantly smaller than q{q}.
IV.2 Krylov growth and Pauli weight spreading
Temporal multiplexing effectively “reverses” the information scrambling caused by quantum evolution. The latter is captured by the Krylov growth of an operator under conjugation [65, 59, 60], which we summarize here in our setting. Specifically, the unitary Heisenberg-evolution [Eq. 18] of a given Pauli observable $P_{k}$ satisfies
$P_{k}^{\text{(H)}}(t)\equiv U(t)^{\dagger}P_{k}U(t)=e^{iHt}P_{k}e^{-iHt}=\sum_{\ell=0}^{\infty}\frac{(it)^{\ell}}{\ell!}\,\mathrm{ad}_{H}^{\ell}(P_{k})$ (31)
where $\mathrm{ad}_{H}(O)\equiv[H,O]$ and $\mathrm{ad}_{H}^{\ell}$ denotes $\ell$ nested commutators with $H$. The time-evolved observable therefore spans the (increasing) Krylov subspace
$\mathcal{K}_{\ell}(P_{k})=\mathrm{span}\bigl\{P_{k},\,\mathrm{ad}_{H}(P_{k}),\,\ldots,\,\mathrm{ad}_{H}^{\ell-1}(P_{k})\bigr\}.$ (32)
While there is no strict one-to-one identification of the integer Krylov order $\ell$ with the continuous time $t$, $\ell$ can be understood as the depth of a truncation of the nested-commutator series. Due to the suppression by the factorial, a rough estimate is $\ell\sim t\,\|H\|$ in case of local Hamiltonians. It is useful to note that for the TFIM models in Eq. 5, the Hadamard equivalence (6) implies that
$(Z_{j_{1}}\cdots Z_{j_{k}})^{\text{H}}(t)\big|_{\text{zz-x}}=U_{\text{H}}^{\dagger}(X_{j_{1}}\cdots X_{j_{k}})^{\text{H}}(t)\big|_{\text{xx-z}}U_{\text{H}}$ (33)
i.e., Pauli-$Z$ observables evolve under $H_{\text{TFIM}}^{\text{zz-x}}$ equivalently to Pauli-$X$ observables under $H_{\text{TFIM}}^{\text{xx-z}}$ up to a global unitary.
Figure 1: Time evolution of the Pauli weight average $\bar{\nu}_{k}$ [Eq. 35] of a single-site Pauli $\sigma_{k}$ in the hhJJ coupling parameter space of the TFIM (5a) for $n=3$ qubits.
In the PTM approach, Krylov growth can be represented in terms of operator spreading in the Pauli basis [66]. Accordingly, expanding the time-evolved observable $P_{k}^{\text{(H)}}(t)$ gives
$P_{k}^{\text{(H)}}(t)=\sum_{j=0}^{d^{2}-1}V_{kj}(t)P_{j}$ (34)
where we used cyclicity of the trace to introduce $V(t)$ from Eq. 19. Thus, for non-Clifford $H$, an initial weight fully localized on $P_{k}$ spreads to other (higher and lower order) Pauli strings as $t$ increases. A convenient diagnostic is the Pauli (Hamming) weight $\nu(P_{j})\in\{0,\ldots,n\}$, i.e., the number of non-identity factors in $P_{j}$, and its coefficient-weighted average
$\bar{\nu}_{k}(t)=\sum_{j=0}^{d^{2}-1}\nu(P_{j})\,V_{kj}(t)^{2}.$ (35)
For local (including nonintegrable/chaotic ones) Hamiltonians, the Lieb–Robinson bound [67, 68, 69, 70, 47] implies that an initially local operator can only develop support within distance $\lesssim v_{\mathrm{LR}}t$ until finite-size saturation:
$\bar{\nu}_{k}(t)\sim\begin{cases}1+\mathrm{const}\cdot v_{\mathrm{LR}}t&\text{(intermediate times)}\\ \mathcal{O}(n)&\text{(late times)}\end{cases}$ (36)
with a typical late-time value $\bar{\nu}_{k}\approx\tfrac{3}{4}n$, corresponding to a uniform distribution of coefficients $V_{kj}^{2}$ over Pauli strings. From the definition (18) and the expansion (34), it straightforwardly follows that $\|P_{k}^{(H)}\|=d=d\sum_{j=0}^{d^{2}-1}\|V_{kj}\|^{2}$, and thus $\sum_{j=0}^{d^{2}-1}V_{kj}^{2}=1$ for all $t$. Consequently, for Haar-random unitaries, one expects a typical late-time magnitude
$|V_{kj}(t\to\infty)|=2^{-n}$ (37)
This is the origin of the exponential concentration effect, where the output of the QELM essentially looses any information about the input for large qubit numbers $n\to\infty$ [51]. Integrable models (such as the TFIM at its integrable points) typically show more structured behavior due to stable quasiparticle modes. An extreme case is a Clifford unitary, for which $V_{kj}=\mathrm{const}.$ By contrast, a Haar-random unitary on the full Hilbert space can generate rapid delocalization [71, 72, 73, 74, 75]. The evolution of the Pauli weight average is illustrated in Fig. 1.
Figure 2: Visualization of the operator spreading (34) via the effective PTM $R$ [Eq. 24] constructed in the temporal multiplexing scheme for the TFIM (a) and random unitary (b). The selected observable set $\mathcal{S}$ are all weight-1 Pauli operators of a $n=3$ qubit system. With increasing evolution time (from top to bottom), these initial operators develop nonzero projection onto the full $d^{2}$-dimensional Pauli basis [Eq. 34]. In the TFIM case, the supporting basis exhibits a sparse pattern, whereas it is dense for a random unitary.
IV.3 Pauli feature decoding
Following Eq. 25, we denote the set of Pauli features available at the readout by $F_{k}=R_{kj}\phi_{j}$ ($k=1,\ldots,{m}$), which emerge by linear transformations of the input features $\phi_{j}$ ($j=1,\ldots,{q}\leq d^{2}$) via the effective PTM $R\in\mathbb{R}^{{m}\times{q}}$ [Eq. 24], possibly including temporal multiplexing. We now ask for conditions that enable the regressor to isolate a given input feature $\phi_{r}(x)$ by finding weights $\mathbf{w}$ such that (approximately)
$\mathbf{w}\cdot\mathbf{F}\propto\phi_{r}\quad\Leftrightarrow\quad\mathbf{w}^{\top}R=\omega\mathbf{e}_{r}^{\top}$ (38)
where $\omega$ is a free parameter and $\mathbf{e}_{r}$ is the $r$-th basis vector of $\mathbb{R}^{{q}}$.
If $R$ corresponds to the full unital channel, then $R=V(t)\in\mathbb{R}^{{q}\times{q}}$ is invertible and any feature can be decoded by
$\mathbf{w}=\omega V(t)\mathbf{e}_{r}.\qquad\text{(unital channel)}$ (39)
If we select subsets of observables, then Eq. 38 can in general not be satisfied for arbitrary $r$ since features get mixed under a quantum channel. The exact condition derived in Appendix B provides a Pauli feature decodability score:
$\gamma_{r}^{2}\equiv(R^{+}R)_{rr},\qquad\sum_{r=1}^{d^{2}}\gamma_{r}^{2}=\mathrm{rank}(R)\leq d^{2}$ (40)
and associated decoding weights
$\mathbf{w}=\omega\left((R^{+})_{r,:}\right)^{\top}.$ (41)
These coincide with the minimum-norm least-squares solution. Exact decodability of the $r$-th feature is indicated by $\gamma_{r}^{2}=1$. A values $\gamma_{r}^{2}<1$ implies that the feature can only be reconstructed up to an error $1-\gamma_{r}^{2}$. We take $\gamma_{r}^{2}\gtrsim 0.5$ as a useful decodability threshold. Note that $\gamma_{r}^{2}$ is a geometric notion of feature decodability and ignores statistical correlations (see Appendix C for further discussion).
Both score and weights preserve tensor product structure of the general PTM $R$: if $R=\otimes_{i=1}^{n}R_{i}$ then $\gamma_{r}^{2}=\prod_{i=1}^{n}\gamma_{r,i}^{2}$ and $\mathbf{w}=\otimes_{i=1}^{n}\mathbf{w}_{i}$. For $R=\mathcal{S}V(t)$, one obtains $\gamma_{r}^{2}=\sum_{k\in\mathcal{S}}V(t)^{2}_{kr}$ showing that feature $r$ is decodable only to the extent that the “spread” of that Pauli under $V(t)$ lands in the measured subset $\mathcal{S}$. [^2] If $R$ contains a $t=0$ block, then since $V(0)=\mathbbm{1}_{d^{2}}$ the corresponding rows of $R$ include $\mathbf{e}_{j}^{\top}$ for all $j\in\mathcal{S}$ (up to basis ordering), hence $\mathbf{e}_{j}\in\mathrm{row}(R)$. Since $R^{+}R$ is the orthogonal projector onto $\mathrm{row}(R)$, $(R^{+}R)\mathbf{e}_{j}=\mathbf{e}_{j}$ and thus $\gamma_{j}^{2}=(R^{+}R)_{jj}=1$. Specifically, for Haar-random unitaries, late-time decodability decays exponentially with qubit number $n$ [see Eq. 37],
$\gamma_{r}^{2}\big|_{t\to\infty}\sim|\mathcal{S}|4^{-n}$ (42)
Noisy channels generically have invertible PTMs [see Section D.2] and hence do not strictly reduce the accessible feature subspace. However, they are contracting and can effectively suppress high-weight Pauli features [see Eq. 84], requiring large weight norms $\|\mathbf{w}\|$ and thereby degrading conditioning. Such non-CPTP maps are generally unstable under finite sampling and thus effectively lead to information loss.
Figure 3: Projector $R^{+}R$ of the effective PTM [Eq. 24] for different unitary types and a complete weight-1 Pauli observable set. Panels (a,b) directly correspond to the $R$ shown in Fig. 2, which are constructed by temporal multiplexing. Panel (c) shows the projector based on $R=V(t=0.5)$ (no temporal multiplexing). The dashed white lines indicate features for which $\gamma_{j}^{2}>0.5$, while the black lines indicate Pauli weight sector boundaries. Due to operator spreading over a time $t=0.5$, decodability in (c) has significantly decreased.
Operator spreading under temporal multiplexing is visualized in Fig. 2, where the effective PTM $R$ is plotted for the case of a TFIM and a random unitary. The related projector matrices $R^{+}R$ are shown in Fig. 3(a,b). Since we include here $t=0$ in the multiplexing scheme, the first $|\mathcal{S}|=9$ rows of $R$ have 1 on the diagonal (no spreading), leading to perfect decodability of the weight-1 Pauli sector. Without temporal multiplexing and nonzero initial time, perfect feature reconstruction is not possible, see Fig. 3(c).
(a)Refer to caption  (b)Refer to caption
(c)Refer to caption  (d)Refer to caption
(e)Refer to caption  (f)Refer to caption (g)Refer to caption
Figure 4: Feature decodability score $\bar{\gamma}_{j}^{2}$ obtained from the effective PTM $R$ [Eq. 40], averaged over each Pauli weight sector, of a $n=3$ qubit model as a function of temporal multiplexing iteration length L{L} [Eq. 24]. The unitary evolution time is given by $t_{L}={L}$ (i.e., initial time $t_{1}=1$). The corresponding rank and number of rows of $R$, representing the measurement budget (B=mL{B}={m}{L}), are shown on the top axis (as a fraction of the total operator space dimension $d^{2}=64$). We consider the TFIM (5a) in (a,c,e), the TFIM (5b) in (g), and random Hamiltonians in (b,d,f). The observable set $\mathcal{S}$ is either only $Z$-Paulis (a,b), $Z$- and ZZZZ-Paulis (c,d,g), or all weight-1 Paulis (e,f).
The evolution with multiplexing steps L{L} of the mean decodability $\bar{\gamma}_{r}^{2}$ of each Pauli sector is further illustrated in Fig. 4. Since we now start with $t=1$, the selected low-weight Paulis are already spread and initial decodability is low on average  [^3] Including $t\approx 0$ in the temporal multiplexing scheme improves decodability at early iterations, but does not affect the final decodability for sufficiently large L{L}.. Compared to structured Hamiltonians, like the TFIM, random Hamiltonians lead to a consistently higher rank of $R$ and a more even decodability over Pauli sectors for large L{L}. In general, we find that temporal multiplexing with Pauli $Z$ and $ZZ$ observables and random unitaries requires ${L}\sim d^{2}/|\mathcal{S}|$ measurements to achieve high decoding scores in all Pauli sectors, along with a saturated rank $\mathrm{rank}(R)\sim|\mathcal{S}|{L}\sim d^{2}$. In practice, however, this exponential measurement budget is usually not needed, since important features can be statistically decoded with much fewer measurements (see Figs. 6 and 7 below).
Summary.
The decodability framework is essentially a reformulation of pseudo-inverse solution of the RC training and provides a handle on the available features [78, 79, 80, 52].The decodability score (40) is a data-independent property of the chosen measurement observables and quantum operations after the encoding. For moderate qubit numbers $n$, full Pauli feature decodability is guaranteed when one either performs a full measurement (${m}=d^{2}$) on an invertible quantum channel, or one employs a temporal multiplexing scheme leading to a full column rank of $R$, again requiring an exponential measurement budget ${m}{L}\sim d^{2}$. Practically sufficient decodability can already be achieved with modest scores and polynomial-in-$n$ measurement budgets, as typically done in applications [1, 44, 60, 61]. Structured unitaries can be advantageous in the large-$n$ limit to avoid exponential concentration issues [51].
IV.4 Statistical decodability of input data
We focused above on the Pauli features $\phi_{j}$, which are generated by the encoding and transform linearly under the action of the QELM. However, they are themselves nonlinear functions of the classical input $u$. For example, with amplitude encoding (8), a single Pauli-$Z$ generates a linear readout $\langle Z\rangle\propto u$, a single Pauli-$X$ can introduce significant nonlinearities since $\langle X\rangle=\sqrt{1-(2u-1)^{2}}=1-\frac{1}{2}(2u-1)^{2}-\frac{1}{8}(2u-1)^{4}+\mathcal{O}(u^{6})$ (cf. Table 1). It is therefore important to understand the actual nonlinearities in $u$ present at the readout. This is a classical problem in reservoir computing, studied initially as memory capacity [80] and later generalized towards information processing capacity [52, 53, 81, 44]. We discuss this, as well as an alternative approach, next.
IV.4.1 Nonlinear degree at readout
In order to estimate the degree of nonlinearity in $u$ available at the readout, one typically calculates the empirical correlation between the readout features $\mathbf{F}(\mathbf{u})$ [see Eq. 16] and certain nonlinear target functions $y(\mathbf{u})$. Raw monomials like $\{u^{j}\}$ are strongly correlated under typical distributions of $u$ and thus do not allow for clear identification of the nonlinear degree within this statistical approach. Following [52, 44], we therefore use orthogonal polynomials $p_{j}(u_{\alpha})$ (degree $j$) to construct orthogonal degree-$k$ targets as $y_{k}(\mathbf{u})=\prod_{i=1}^{n}p_{e_{i}}(\mathbf{u})$, with $\sum_{i=1}^{n}e_{i}=k$. Specifically, for amplitude encoding, we use (shifted) Legendre polynomials, which are orthonormal w.r.t. the uniform distribution [1,1][-1,1] or $[0,1]$; for angle encoding, we use Hermite polynomials, which are orthonormal w.r.t. the Gaussian distribution $\mathcal{N}(0,1)$.
(a)Refer to caption  (b)Refer to caption
Figure 5: Ideal nonlinear capacity score $\mathcal{R}^{2}(k)$ [Eq. 43] for amplitude and rotational encodings [Eqs. 8 and 10], indicating the ability of an ideal QELM $f(\mathbf{u})=\mathbf{w}^{\top}\boldsymbol{\upphi}(\mathbf{u})$ to generate degree-$k$ monomials of the input $u_{\alpha}$ ($\alpha=1,\ldots,n$). Unless indicated, only Pauli-$Z$ features are measured in (a). Nonlinear capacity is significantly enhanced when measuring over all Paulis (b). This is also demonstrated by the score for amplitude encoding with Pauli-$X$ readout included in (a). The dash-dotted curve represents the theoretical score for $\mathcal{R}^{2}_{Z}(k)$ with amplitude encoding and $n=5$ input qubits [Eq. 44], which agrees exactly with the numerical results. $\Sigma$ denotes the integrated capacity over all $k$. The curves drop with increasing $k$ is because the quantum model can only generate square-free monomials of the Pauli features $\phi_{j}$.
We take the coefficient of determination [52, 53]
$\mathcal{R}^{2}(k)=1-\frac{\mathbb{E}_{\mathbf{u}}(y_{k}(\mathbf{u})-\hat{f}(\mathbf{u}))^{2}}{\mathrm{Var}_{\mathbf{u}}y_{k}(\mathbf{u})},$ (43)
to indicate whether the trained predictor $\hat{f}(\mathbf{u})$ can generate nonlinearities of degree $k$ in the input $u$  [^4] In Eq. 43, $\mathrm{Var}_{\mathbf{u}}y(\mathbf{u})$ denotes the variance of $y(\mathbf{u})$ over the distribution of $\mathbf{u}$.. $\mathcal{R}^{2}$ generalizes the geometric score (40) by taking statistical correlations into account (see Appendix C). In fact, for whitened orthonormal targets, one has $\mathcal{R}^{2}(k)=\gamma_{k}^{2}$. In the case of amplitude encoding (8) with Pauli-$Z$ readout, the number of representable degree-$k$ targets is $\binom{n}{k}$, which exponentially decreases compared to the total number of degree-$k$ monomials $\binom{n+k-1}{k}$, giving an expected score
$\mathcal{R}^{2}_{Z}(k)=\begin{cases}\binom{n}{k}/\binom{n+k-1}{k}&\text{for }k\leq n\\ 0&\text{for }k>n\end{cases}$ (44)
In Fig. 5, we first display the nonlinear capacity that can be ideally achieved from the encoded state without effects of feature mixing, i.e., we take $\mathbf{F}=\boldsymbol{\upphi}$ in (25). As shown in panel (a), a Pauli-$X$ readout (black symbols with dotted curves) generates even degree monomials, behaving similarly to angle encoding with Pauli-$Z$ observables (connected open symbols). In Fig. 5(b) we show the nonlinear capacity score $\mathcal{R}^{2}(k)$ obtained for a complete Pauli observable set. Notably, the presence of non-$Z$ Paulis significantly enhances the nonlinearity present in the readout, as becomes clear by comparison to the dash-dotted curve corresponding to $\mathcal{R}^{2}_{Z}(k)$ for $n=5$ [Eq. 44]. Moreover, angle and amplitude encodings show a similar behavior. The decay of $\mathcal{R}^{2}(k)$ with $k$ stems from the fact that the encodings used here can only generate square-free monomials of the Pauli features $\phi_{j}$, while we test against (the practically relevant case of) all possible degree-$k$ targets.
(a)Refer to caption  (b)Refer to caption
(c)Refer to caption  (d)Refer to caption
Figure 6: Nonlinear capacity score $\mathcal{R}^{2}(k)$ [Eq. 43] as a function of the nonlinear order $k$ and the temporal multiplexing iteration length L{L} [Eq. 24]. The unitary evolution time is given by $t_{L}=L$. We consider a $n=3$ qubit model in the same setup as in Fig. 4. The unitary is based on the TFIM (5a) (a,c) or a random Hamiltonian (b,d), while the observable set $\mathcal{S}$ consists of either only $Z$-Paulis (a,b) or $Z$- and ZZZZ-Paulis (c,d).
(a)Refer to caption  (b)Refer to caption
(c)Refer to caption  (d)Refer to caption
Figure 7: Relative nonlinear capacity score $\mathcal{R}^{2}(k)/(\mathcal{R}^{2}_{\text{all}}(k)+\delta)$ [Eq. 43] as a function of the nonlinear order $k$ and the temporal multiplexing iteration length L{L} [Eq. 24]. $\mathcal{R}^{2}_{\text{all}}$ is the nonlinear capacity score obtained for a complete Pauli observable set, and we choose $\delta=0.1$ as a cut-off since scores $\mathcal{R}^{2}_{\text{all}}\lesssim 0.1$ are not considered meaningful. We use the same setup as in Fig. 6. The unitary is based on the TFIM (5a) (a,c) or a random Hamiltonian (b,d), while the observable set $\mathcal{S}$ consists of either only $Z$-Paulis (a,b) or $Z$- and ZZZZ-Paulis (c,d).
Figure 6 illustrates the nonlinear capacity score $\mathcal{R}^{2}(k)$ obtained for a QELM under temporal multiplexing of Pauli-$Z/ZZ$ observables with the TFIM or random Hamiltonians. The score follows a similar pattern as the data-independent Pauli decodability $\gamma_{j}^{2}$ (see Fig. 4) as they both show a delayed growth with increasing nonlinear order $k$ and multiplexing step $L$. However, $\mathcal{R}^{2}(k)$ saturates faster than $\gamma_{j}^{2}$ because the considered nonlinear target functions $y_{k}(\mathbf{u})$ can be generated already from relatively few Pauli features. The decrease of $\mathcal{R}^{2}(k)$ with $k$ stems from two facts: (i) the encodings used here can generate only a small subset of all possible degree-$k$ monomials (see Eq. 44 and Fig. 5). A clearer measure of the effectiveness of temporal multiplexing to generate nonlinearities is thus the score $\mathcal{R}^{2}(k)$ normalized by the score $\mathcal{R}^{2}_{\text{all}}(k)$ obtained when measuring over the complete Pauli observable set (see Fig. 7). (ii) Higher-order observables can be missing for highly structured Hamiltonians or insufficient multiplexing steps. Hamiltonians with a more global and random structure are thus beneficial when measurements are restricted to Pauli-$Z$ observables but higher-order nonlinearities are desired.
IV.4.2 Nonlinear monomials at readout
(a)Refer to caption
(b)Refer to caption  Refer to caption
Figure 8: Ability to construct monomials of the input $\{x,y,z\}$ for amplitude encoding (8) and rotational encoding (10) with $n=D=3$. (a) Decodability score $\kappa_{j}^{2}$ according to Eq. 45. (b) Projector matrix $A^{+}A$, where nonzero off-diagonals correspond to $\kappa_{j}^{2}<1$ and indicate decoding is possible only within a certain accuracy. The underlying Taylor expansion is performed up to order $r=4$ in each $u_{j}=\{x,y,z\}$ and changes are mild upon further increasing $r$. The scores and matrices are averaged over a uniform grid of expansion points $\mathbf{u}_{0}\in[0.1,0.9]^{3}$ for amplitude encoding and $\mathbf{u}_{0}\in[-1,1]^{3}$ for rotational encoding. For readability, monomial labels are incompletely displayed in (a) and replaced by column/row indices in (b).
Raw monomials in $u_{\alpha}$ are often relevant features for practical applications, such as autoregressive learning of dynamical systems. Let thus $\mathcal{B}(\mathbf{u})=\{1,u_{1},\ldots,u_{1}u_{2},\ldots,\mathcal{O}(u^{r})\}$ be the vector of size $b=\binom{D+r}{r}$ of monomials of $\{u_{\alpha}\}_{1\leq\alpha\leq D}$ up to degree $r$. The order-$r$ Taylor expansion of the Pauli features $\{\phi_{j}(\mathbf{u})\}_{1\leq j\leq q}$ within a small region around some operating point $\mathbf{u}_{0}$ (e.g., data mean) can then be written as $\boldsymbol{\upphi}(\mathbf{u})=A\mathcal{B}(\mathbf{u})$ with a coefficient matrix $A\in\mathbb{R}^{q\times b}$. As before, we can define a geometric decodability score
$\kappa_{j}^{2}\equiv(A^{+}A)_{jj},\qquad\sum_{j=1}^{b}\kappa_{j}^{2}=\mathrm{rank}(A)\leq\min(q,b),$ (45)
which measures how accurately a certain monomial $\mathpzc{b}\in\mathcal{B}$ can be constructed from linear combinations of $\phi_{j}$. This score depends on the operating point $\mathbf{u}_{0}$, but is independent of quantum properties of the reservoir, and ranges between 0 and 1 (no/exact construction possible). Intermediate values indicate that construction is possible only within an error $\sim\sqrt{1-\kappa_{j}^{2}}$ due to off-diagonal terms in the $j$-th row/column of the projector $A^{+}A$. A somewhat artificial dependence of $\kappa_{j}^{2}$ on the Taylor order $r$ (owing to the rank constraint) can be mitigated by considering large $r$ and $b$. Indeed, most of the $\kappa_{j}^{2}$ decrease upon increasing $r$, except for a few stable ones. These include monomials that arise directly from the tensor product structure of the encoding state. For amplitude encoding (8), these are all square-free monomials up to order $n$ (e.g., for $n=D=3$: $\mathcal{B}_{\text{stable}}=\{1,x,y,z,xy,xz,yz,xyz\}$).
The typical behavior of $\kappa_{j}^{2}$ and of the projectors $A^{+}A$ (averaged over a range of expansion points) is illustrated in Fig. 8. While the score typically decays with increasing nonlinear degree, rotational encoding tends to have a greater expressivity towards higher-order monomials. This is associated with less leakage into off-diagonal terms of $A^{+}A$ [see Fig. 8(b)], and can explain why rotational encoding often leads to robust performance of QELMs.
IV.5 Summary
Feature decodability is the underlying concept that determines the expressivity of a QELM. Pauli features are the fundamental “building blocks” of the QELM expressivity. Any other nonlinear feature required for the learning task must be constructible from Pauli features. For the encoding types considered here, expressing simple monomials is often only possible up to a certain approximation error ($\kappa_{j}^{2}<1$), which increases with the size of the data domain. Knowing the structure of the readout features allows to design feature engineering strategies to enhance the expressivity at low cost.
V Interpretable learning of dynamical systems
V.1 Time series modeling with QELMs
The classical representation (25) allows us to identify the surrogate model a QELM learns during the training process. Specifically, for the product state encodings considered in Section II.2, the function library $\boldsymbol{\upphi}:\mathbb{R}^{D}\to\mathbb{R}^{{q}}$ consists of ${q}=3^{n}$ monomials of the single-Pauli features $\phi_{k}$ (recall that observables containing Pauli-$Y$ factors have vanishing expectation values, see Table 1). We consider the standard problem of learning to forecast a chaotic dynamical system by training on its sampled trajectory $\mathbf{u}_{t}=\mathbf{u}(t_{0}+t\Delta t)\in\mathbb{R}^{D}$. The continuous and discrete evolutions are given by
[…]
The non-integrable nature and fractal attractors of chaotic systems make them natural benchmarks for system identification, which addresses the question to which extent the underlying data-generating process can be uniquely discovered from the data [83, 54].
If the dynamics is Markovian, it can be modeled solely with a NG-RC or QELM based on an instantaneous feature map [84, 29]. Accordingly, we formulate the training problem of the QELM (25) as
$\text{QELM:}\qquad\mathbf{u}_{t+1}=\mathbf{f}(\mathbf{u}_{t})=W^{\top}R\boldsymbol{\upphi}(\mathbf{u})$ (48)
[…]
We use the Lorenz-63 and Halvorsen models as benchmarks ($D=3$ dimensions) [85]:
$\text{Lorenz-63:}\;\begin{cases}\dot{x}=\sigma(y-x)\\ \dot{y}=x(\rho-z)-y\\ \dot{z}=xy-\beta z\end{cases}\qquad\text{Halvorsen:}\;\begin{cases}\dot{x}=-ax-4y-4z-y^{2}\\ \dot{y}=-ay-4z-4x-z^{2}\\ \dot{z}=-az-4x-4y-x^{2}\end{cases}$ (49)
with parameters $\sigma=10$, $\beta=8/3$, $\rho=28$ for Lorenz-63, and $a=1.4$ for Halvorsen. This setup leads to chaotic behavior with largest Lyapunov exponents $\lambda_{\text{max}}^{\text{Lor63}}\approx 0.89$ and $\lambda_{\text{max}}^{\text{Halv}}\approx 0.76$, respectively. The polynomial form of the vector field in (49) then suggests to use amplitude encoding (8) with $n=3$ qubits, leading to a library of the following 27 terms:
$\boldsymbol{\upphi}_{\text{amp}}^{n=3}(\mathbf{u})=\Bigl\{$ $1,2\sqrt{(1-x)x},2x-1,2\sqrt{(1-y)y},2(2x-1)\sqrt{-((y-1)y)},4\sqrt{(x-1)x(y-1)y},2y-1,$ (50)
$2\sqrt{-((x-1)x)}(2y-1),(2x-1)(2y-1),2\sqrt{(1-z)z},2(2x-1)\sqrt{-((z-1)z)},$ (51)
$2(2y-1)\sqrt{-((z-1)z)},2(2x-1)(2y-1)\sqrt{-((z-1)z)},4\sqrt{(x-1)x(z-1)z},$ (52)
$4(2y-1)\sqrt{(x-1)x(z-1)z},4\sqrt{(y-1)y(z-1)z},4(2x-1)\sqrt{(y-1)y(z-1)z},$ (53)
$8\sqrt{-((x-1)x(y-1)y(z-1)z)},2z-1,2\sqrt{-((x-1)x)}(2z-1),(2x-1)(2z-1),$ (54)
$2\sqrt{-((y-1)y)}(2z-1),2(2x-1)\sqrt{-((y-1)y)}(2z-1),4(2z-1)\sqrt{(x-1)x(y-1)y},$ (55)
$(2y-1)(2z-1),2\sqrt{-((x-1)x)}(2y-1)(2z-1),(2x-1)(2y-1)(2z-1)\Bigr\}$ (56)
We will also consider the amplitude encoding scheme of Eq. 9, which is obtained by replacing in Eq. 56 $\{x,y,z\}\to\{x^{2},y^{2},z^{2}\}$. Depending on the dynamical system, we will observe differences in the performance.
In order to confine the trajectory to the interval $\mathbf{u}\in[0,1]^{3}$ or $1,1]^{3}$, as required by the encoding, we apply the transformation $\mathbf{u}\to\mathbf{u}^{\prime}=\boldsymbol{\alpha}\cdot(\mathbf{u}-\mathbf{m})$ with $\boldsymbol{\alpha},\mathbf{m}\in\mathbb{R}^{3}$ to the state. This leads to the rescaled Lorenz-63 model:
$\dot{u}$ $=-\sigma u+\sigma\frac{\alpha_{x}}{\alpha_{y}}v+\sigma\alpha_{x}(m_{y}-m_{x}),$
$\dot{v}$ $=-v+\frac{\alpha_{y}}{\alpha_{x}}(\rho-m_{z})u-\frac{\alpha_{y}}{\alpha_{z}}m_{x}w-\frac{\alpha_{y}}{\alpha_{x}\alpha_{z}}uw+\alpha_{y}m_{x}(\rho-m_{z})-\alpha_{y}m_{y},$ (57)
$\dot{w}$ $=-\beta w+\frac{\alpha_{z}}{\alpha_{x}\alpha_{y}}uv+\frac{\alpha_{z}}{\alpha_{x}}m_{y}u+\frac{\alpha_{z}}{\alpha_{y}}m_{x}v+\alpha_{z}(m_{x}m_{y}-\beta m_{z}),$
and the rescaled Halvorsen model:
$\dot{u}$ $=-au-4\frac{\alpha_{x}}{\alpha_{y}}v-4\frac{\alpha_{x}}{\alpha_{z}}w-2\frac{\alpha_{x}m_{y}}{\alpha_{y}}v-\frac{\alpha_{x}}{\alpha_{y}^{2}}v^{2}+\alpha_{x}(-am_{x}-4m_{y}-4m_{z}-m_{y}^{2}),$
$\dot{v}$ $=-av-4\frac{\alpha_{y}}{\alpha_{z}}w-4\frac{\alpha_{y}}{\alpha_{x}}u-2\frac{\alpha_{y}m_{z}}{\alpha_{z}}w-\frac{\alpha_{y}}{\alpha_{z}^{2}}w^{2}+\alpha_{y}(-am_{y}-4m_{z}-4m_{x}-m_{z}^{2}),$ (58)
$\dot{w}$ $=-aw-4\frac{\alpha_{z}}{\alpha_{x}}u-4\frac{\alpha_{z}}{\alpha_{y}}v-2\frac{\alpha_{z}m_{x}}{\alpha_{x}}u-\frac{\alpha_{z}}{\alpha_{x}^{2}}u^{2}+\alpha_{z}(-am_{z}-4m_{x}-4m_{y}-m_{x}^{2}).$
This vector field-level transformation allows us to directly compare the learned coefficients with the true ones.
V.2 Numerical results
We use here a complete measurement operator set of $4^{n}$ Pauli observables. In this setup, the reservoir unitary becomes immaterial [see Eq. 29], except for possible numerical conditioning issues. We generate trajectories using a 5th-order Runge-Kutta integrator with a time step of $\Delta t=0.01$. Reported values of the training error and forecast horizon are averaged over multiple initial conditions. After training the QELM, we obtain a classical surrogate model $\hat{\mathbf{f}}(\mathbf{u})$ by inserting into Eq. 25 the learned weights $\hat{\mathbf{w}}$.
The RMS training error
$\epsilon_{\text{train}}=\Big(\frac{1}{P}\sum_{t=1}^{P}\|\mathbf{u}_{t+1}-\hat{\mathbf{f}}(\mathbf{u}_{t})\|^{2}\Big)^{1/2},$ (59)
measures one-step accuracy of the trained predictor $\hat{\mathbf{f}}$. As a measure of multi-step prediction accuracy, the autonomously generated trajectory $\hat{\mathbf{u}}_{t+1}=\hat{\mathbf{f}}(\hat{\mathbf{u}}_{t})$ is compared to the true trajectory via the forecast horizon,
$T_{\text{fch}}=\inf\left\{t:\|\mathbf{u}_{t}-\hat{\mathbf{u}}_{t}\|>\sigma\right\}$ (60)
where $\bar{\mathbf{u}}=\frac{1}{P}\sum_{t=1}^{P}\mathbf{u}_{t}$ and $\sigma^{2}=\frac{1}{P}\sum_{t=1}^{P}\|\bar{\mathbf{u}}-\mathbf{u}_{t}\|^{2}$ are the mean and variance of the trajectory, respectively. We average $T_{\text{fch}}$ over multiple starting points $\mathbf{u}_{0}\in\mathbb{R}^{D}$ and express it in units of the Lyapunov time $T_{L}=1/\lambda_{\text{max}}$.
Figure 9: QELM vs. its classical representation: Forecast obtained from a QELM [Eq. 1, ‘QRC’] with amplitude encoding (8) and $n=3$ qubits, trained on the rescaled Lorenz-63 model. This is macthed exactly by the forecast of the iterated classical representation [Eq. 25, ‘Surrogate’]. Both eventually diverge from the true trajectory due to the Lyapunov instability. Ground-truth trajectories are generated with a resolution of $\Delta t=0.01$. Time is given in model time units. For the scaling transformation in Eq. 57, we use parameters $\boldsymbol{\alpha}=0.004\cdot[1,1,1]$ and $\mathbf{m}=[-30,-30,-5]$.
System type QELM setup $\alpha$ $\lambda$ $\epsilon_{\text{train}}$ $T_{\text{fch}}/T_{L}$
Lorenz-63 std. ampl. enc. $[0,0.3]$, Eq. 8 - $10^{-6}$ $10^{-5}$ $4.3\pm 1.7$
Lorenz-63 resc. ampl. enc. $[0,1]$, Eq. 8 0.005 $10^{-8}$ $10^{-6}$ $6.2\pm 1.4$
Halvorsen std. ampl. enc. $[0.1,0.8]$, Eq. 8 - $10^{-8}$ $10^{-4}$ $2.3\pm 0.8$
Halvorsen std. …+ add squares to readout vector - $10^{-8}$ $10^{-5}$ $5.8\pm 1.7$
Halvorsen resc. ampl. enc. $[0,1]$, Eq. 8 0.027 $10^{-7}$ $10^{-4}$ $3.4\pm 1.0$
Halvorsen resc. …+ add squares to readout vector 0.027 $10^{-7}$ $10^{-6}$ $6.8\pm 1.7$
Halvorsen std. ampl. enc. [0.01,0.01][-0.01,0.01], Eq. 9 - $10^{-15}$ $10^{-7}$ $5.9\pm 1.2$
Halvorsen resc. ampl. enc. [1,1][-1,1], Eq. 9 0.001 $10^{-15}$ $10^{-7}$ $5.9\pm 1.5$
Table 2: Training RMS errors $\epsilon_{\text{train}}$ [Eq. 59] and forecast horizons $T_{\text{fch}}$ [Eq. 60] for the benchmark systems and (unitary-free) QELM setups discussed in the text. We generally use $n=3$ input qubits (no delay states) and a time resolution of $\Delta t=0.01$. $T_{\text{fch}}$ with its standard deviation (not error of the mean) is given in units of the Lyapunov time $T_{L}$. The column ‘type’ indicates whether we use the standard [Eq. 49] or rescaled dynamical system. In the latter case, $\alpha$ gives the rescaling factor in Eqs. 57 and 58 (same for all coordinates).
Results of training a QELM in various setups are summarized in Table 2. As exemplified for the rescaled Lorenz-63 model in Fig. 9, we find exact agreement between the autonomously generated prediction of the trained QELM and the iterated classical model. Both remain close to the true trajectory for a time $T_{\text{fch}}/T_{L}\approx 6$, when the intrinsic Lyapunov instability of the chaotic system leads to a divergence, triggered by errors in the initial condition, discretization, and extra terms in the surrogate model (see below)  [^5] We find that training on increments via $\mathbf{u}_{t+1}-\mathbf{u}_{t}=\mathbf{f}(\mathbf{u}_{t})$ does not significantly improve the forecast horizon, unless one uses small integrator time steps $\Delta t\lesssim 0.001$..
When repeating the experiments with the Halvorsen model (58), we obtain slightly lower forecast horizons of $T_{\text{fch}}/T_{L}\approx 3$ and RMS training errors of $\epsilon_{\text{train}}\approx 10^{-4}$. One reason for this reduced performance compared to the Lorenz-63 model lies in the absence of squares of the state variables $\{x,y,z\}$ in the library (56). These are crucial for the Halvorsen equations (58), whereas the Lorenz-63 model (57) depends only on cross terms of the state variables. Indeed, when enhancing the readout vector by its squares, prediction accuracy for the Halvorsen model improves significantly (see Table 2). Alternatively, a similar performance can also be achieved when using the amplitude encoding variant in (9), which already involves squares of the state variables. In the context of reservoir computing it is, of course, well known that prediction performance can be enhanced by feature engineering [87, 64, 61]. The present analysis exemplifies that, within the PTM framework, this optimization can be systematically phrased as an inverse problem, where the structure of the ground-truth model dictates the required terms in the encoding.
Figure 10: Prediction performance (forecast horizon and RMS training error) under temporal multiplexing of a QELM trained on the Halvorsen model (58). We use amplitude encoding (9) with $n=3$ qubits, single-qubit Pauli-$Z$ observables, and (a) TFIM Hamiltonian (5a), (b) random Hamiltonian. The top horizontal axis shows the numerically obtained rank of the readout matrix $G$ [Eq. 28] relative to the maximum achievable rank $q=3^{n}$ [Eq. 30]. Forecast horizon is given in units of the Lyapunov time $T_{L}$; no additional features are added to the readout.
When employing temporal multiplexing with a QELM, the rank of the readout matrix $G$ [Eqs. 28 and 30] can serve as an indicator of the training accuracy, as illustrated in Fig. 10. This rank is upper bounded by the number of nonzero Pauli features q{q}, which, for the encoding (9) used here, is ${q}=3^{n}$. When $\mathrm{rank}(G)\simeq{q}$, all Pauli features are available at the readout. This can, in principle, improve performance when conditioning is controlled and exponential concentration effects can be avoided [51, 45]. Indeed, the QELM in Fig. 10(b) achieves training error and forecast horizon essentially identical to the ones reported in Table 2 obtained with a complete Pauli observable set. On the other hand, as seen in Fig. 10(a), performance is poor when the readout utilizes only a few effective observables. In the present setting, this is a consequence of using a structured Hamiltonian like the TFIM [Eq. 5a], which spreads observables less effectively over the full Hilbert space (see Fig. 4), such that $\mathrm{rank}(G)$ is capped by the rank of the PTM $R$.
V.3 Interpretability
(a)Refer to caption (b)Refer to caption (c)Refer to caption (d)Refer to caption
[…]
$\boldsymbol{\upphi}(\mathbf{u})=K\mathcal{B}(\mathbf{u}),$ (61)
and insert this into the learned model (25),
[…]
For the dynamical systems considered here, we can simply take $\mathcal{B}=\mathcal{B}_{r}$ to be the set of monomials in $u_{j}$ up to order $r$, of which there are $N={D+r\choose r}$. Accordingly, $K$ follows from Taylor-expanding $\boldsymbol{\upphi}(\mathbf{u})$ around some operating point $\bar{\mathbf{u}}$ (usually the data mean). As shown in Section IV.4.2, low-order monomials are indeed well constructible from $\boldsymbol{\upphi}$. However, due to the tensor product structure of $\boldsymbol{\upphi}$ [see Eq. 22] and the fact, that the derivatives of the single-qubit feature vector lie in a 2-dimensional space (vanishing Pauli-$Y$), $K$ typically has not full column rank when $N<q$. A consequence of this is that the expression of $\boldsymbol{\upphi}$ in terms of $\mathcal{B}$ is not unique. Nevertheless, deviations are small for the settings considered here  [^7] Deviations are small for $N<q$: in practice, for $n=D=3$ (thus $q=27$) and monomials of order $r=3$ (thus $N=19$ excluding the constant term), we get $\mathrm{rank}(K)=16$.
Figure 11 shows the learned flow maps, as approximated in Eq. 62, resulting from training on the (rescaled) Lorenz-63 and Halvorsen models (no feature engineering in the readout). Comparison to the theoretical expectation [Eq. 47] reveals good agreement especially for low-order coefficients. By contrast, deviations can be significant for higher-order coefficients. However, since they are multiplied by increasing powers of $\Delta t$, their influence on the overall training error is small.
VI Summary
We have described a memoryless quantum reservoir computer (quantum extreme learning machine, QELM) in the Pauli transfer matrix (PTM) formalism, building upon [33, 25, 43, 46]. Writing states and channels in the Pauli basis (or any other orthogonal operator basis) turns the QELM into a sequence of real linear maps $T_{\mathcal{E}_{j}}$ acting on the vector of Pauli features $\boldsymbol{\upphi}(\mathbf{u})$ representing the encoded state $\rho_{\text{enc}}(\mathbf{u})$. These features carry the full dependence of the QELM on the encoded input $\mathbf{u}$ and determine the nonlinearities that the model can possibly express. Unitaries appear as orthogonal rotations on the traceless sector of $\boldsymbol{\upphi}(\mathbf{u})$, while general CPTP maps act affinely and typically contract the feature vector. Training a QELM reduces to classical linear regression over the components $\phi_{k}(\mathbf{u})$ and can be framed as a decoding problem of features mixed by quantum channels.
While the PTM formalism has been previously been applied to QRCs [33, 25, 43, 46], we focused here on its implications for time-series prediction and model identification with QELMs. First, this approach allows one to cast the optimization of a QELM as an inverse problem, in which the desired target features imply an ideal encoding scheme and decodability from (preferably low-weight) Pauli features. Second, the PTM formalism provides a classical representation of the QELM [93, 94, 95, 47] and thereby renders it directly interpretable as a nonlinear vector (auto-)regression model (or next-generation RC [64]). These insights can help to construct reduced-order models in cases where the exact data-generating process is unknown and opens the door to further studies of black-box model identification [83, 96]. We also discussed the problem of temporal multiplexing and operator spreading [1, 33, 44, 59, 60] within the PTM framework, which can enable a more precise control over these and similar measurement enhancement strategies. Using the Pauli basis is justified by the measurement process, but not the only possibility. For future work, it could thus be interesting to study other orthogonal operator sets and assess the generality of the present findings.
Figure 12: Pauli weight average (a) of Heisenberg-evolved measurement observables can predict the trainig error (b) and forecast horizon (c) of a QELM. We use a TFIM Hamiltonian (varying coupling parameters $J,h$, fixed $t=6$) with $n=3$ qubits and amplitude encoding (8) (no temporal multiplexing). The Pauli weight average is data-independent, while the QELM in (b,c) is trained on the Lorenz-63 model and employs 1-qubit Pauli-$Z$ observables in the readout, enhanced by monomials up to order 2 in order to increase sensitivity to nonlinear terms.
It has previously been observed [97, 98, 99, 100, 101, 102, 103] that enhancing the quantum nature of a QML model often does not improve its performance on classical datasets. The present study adds further explanations of this finding from the viewpoint of learning dynamical systems with QELMs: first, a reservoir unitary mixes features and reversing this potentially detrimental effect requires complex measurement schemes (either high-order Pauli observables or temporal multiplexing), which, in turn, puts more demand on the linear decoder. We can demonstrate this mechanism directly by comparing the Pauli weight average of single-qubit observables to its training accuracy. As illustrated in Fig. 12 for a minimal QELM without temporal multiplexing, large Pauli weight average strongly correlates with high training error and short forecast horizon. In conventional QML models, instead, controlling this potentially detrimental feature mixing must be achieved by expensive training of variational gates. Second, optimal training error scaling requires learning of the flow map by a structurally matching feature library [88, 84]. Feature libraries provided by usual encoding schemes do not achieve this a priori and lead to surrogate models with significant errors (cf. Fig. 11). The forecast horizon $T_{\text{fch}}$ can be understood here akin to generalization error for static learning tasks: a nonzero $T_{\text{fch}}$ is only possible when the model has learned the actual rule governing the data  [^8] This is in contrast to the one-step training error, which can be decreased even by linear delay features.. Beside optimizing the measurement process [29], the present work thus points to the encoding as a central turning knob for performance of QELMs and, more generally, QML models (see, e.g., [105, 106] for insights along this line). How this surrogate-model argument gets modified when training on datasets without clear ground truths is an interesting avenue for future work [107].
The $4^{n}$-dimensional space of Pauli features  [^9] Due to symmetries, the actual effective dimension is lower, and typically $\sim 3^{n}$ underlies the exponential frequency content found for QML models that employ rotational encoding [109, 49]. These and similar properties stemming from the exponential Hilbert space dimensionality point to a possible quantum advantages, but require specific measurement strategies. Moreover, simply increasing qubit number $n$ does not generally improve QRC performance, as overfitting and exponential concentration effects can become prevalent [51, 12, 45, 47]. In that regime, more sophisticated reservoir design strategies will be required to steer feature propagation through the quantum layers.
Symbol Meaning
$\mathcal{H}$ Hilbert space
$\mathcal{M}(\mathcal{H})$ Space of Hermitian operators on $\mathcal{H}$
$n$ Number of qubits
$d$ Hilbert space dimension ($d=2^{n}$)
$D$ Input data dimension
$P$ Number of training samples
m{m} Number of measurement operators
L{L} Number of temporal multiplexing steps
B{B} Total number of measurements (e.g., B=mL{B}={m}{L})
q{q} Number of nonzero Pauli features (${q}\leq d^{2}$)
$\mathbf{x}$ Input data ($\mathbf{x}\in\mathbb{R}^{D}$)
$y$ Target value ($y\in\mathbb{R}$)
$\rho(\mathbf{x})$ Density matrix of the quantum state (feature map)
$M_{k}$ Measurement operators
$P_{k}$ $n$-qubit Pauli operators ($k=1,\ldots,d^{2}$)
$\sigma_{j}$ single-site Pauli matrices ($\mathbb{I}$, $X$, $Y$, $Z$)
$U$ Unitary evolution operator
$T_{\mathcal{E}}$ Pauli transfer matrix (channel $\mathcal{E}$, $T_{\mathcal{E}}\in\mathbb{R}^{d^{2}\times d^{2}}$)
$V$ Unitary PTM ($V\in\mathrm{SO}(d^{2})$)
$R$ Effective PTM (temporal multiplexing) ($R\in\mathbb{R}^{{B}\times d^{2}}$)
$\mathcal{S}$ Selected observable set ($|\mathcal{S}|={m}$)
$\boldsymbol{\upphi}(\mathbf{x})$ Pauli feature vector ($\boldsymbol{\upphi}(\mathbf{x})\in\mathbb{R}^{d^{2}}$)
$\mathbf{F}(\mathbf{x})$ Readout feature vector ($\mathbf{F}(\mathbf{x})=R\boldsymbol{\upphi}(\mathbf{x})$)
$\mathbf{w}$ Readout weights ($\mathbf{w}\in\mathbb{R}^{B}$)
$\lambda$ Regularization parameter
$f(\mathbf{x})$ QELM output function ($f(\mathbf{x})=\mathbf{w}^{\top}\mathbf{F}(\mathbf{x})$)
$\nu$ Pauli weight
$\gamma$ Noise parameter
Table 3: Summary of notation used in this work.
Appendix A Kernel representation
Training over a complete basis of measurement operators $\{M_{k}\}_{k=1}^{d}$ results in a quantum kernel form of the QELM [29]:
$f(\mathbf{x})=\sum_{j=1}^{P}\alpha_{j}^{*}K(\mathbf{x},\mathbf{x}_{j})=\mathrm{tr}(M^{*}\rho(\mathbf{x}))$ (63)
with coefficients $\alpha_{j}$ and the kernel [28]
$K(\mathbf{x},\mathbf{x}^{\prime})=\mathrm{tr}\big(\rho(\mathbf{x})\rho(\mathbf{x}^{\prime})\big)=\mathrm{tr}\big(\rho_{\text{enc}}(\mathbf{x})\rho_{\text{enc}}(\mathbf{x}^{\prime})\big)$ (64)
Notably, any data-independent unitary such as $U_{R}$ in (4) drops out from Eq. 63, implying that such a QELM is only dependent on the encoding state $\rho_{\text{enc}}(\mathbf{x})$. Furthermore, Eq. 63 allows to identify an optimal measurement operator:
$M^{*}=\sum_{j=1}^{P}\alpha_{j}^{*}\rho(\mathbf{x}_{j}).$ (65)
For the least-squares loss, the optimal parameters have the explicit representation $\boldsymbol{\alpha}^{*}=(\mathbf{K}+\lambda\mathbbm{1})^{-1}\mathbf{y}$, with the Gram matrix $K_{ij}=K(\mathbf{x}_{i},\mathbf{x}_{j})$.
Appendix B Exact feature isolation criterion: null-space condition
Equation 38 is equivalent to requiring $\mathbf{w}R=\omega e_{r}^{\top}$ with some scalar weight $\omega$, i.e., $e_{r}^{\top}\in\mathbb{R}^{q}$ belongs to the row-space of the effective PTM $R\in\mathbb{R}^{{m}\times{q}}$. This, in turn, is equivalent to
$\mathbf{e}_{r}\perp\mathrm{null}(R),\qquad\text{($r$-th feature isolatable)}$ (66)
i.e., the $r$-th component of every (basis) vector of $\mathrm{null}(R)$ must be zero. If Eq. 66 is met, the solution for $\mathbf{w}$ is provided by the pseudoinverse $R^{+}$ via
$\mathbf{w}=\omega\mathbf{e}_{r}^{\top}R^{+}$ (67)
up to a term $\propto h(\mathbbm{1}_{m}-RR^{+})$ with an arbitrary vector $h\in\mathbb{R}^{1\times{m}}$ accounting for the non-uniqueness of the solution if $R$ does not have full row rank. Without this term, Eq. 67 represents the minimum-norm solution.
Figure 13: Null-space criterion (66): Number of isolatable features as a function of the number of observables (=rows of the PTM) m{m}, for (a) the TFIM at various evolution times $t$ and (b) a random unitary, for $n=4$ qubits ($d^{2}=256$). Panel (a) shows the total number of isolatable features, while panel (b) shows the number of isolatable 1-qubit features. In contrast to statistical decodability [cf. Appendix C], isolatability is an exact geometric criterion independent of the data.
Consider the case where $R\in\mathbb{R}^{{m}\times{q}}$ is dense and quasi-random, arising, e.g., for a Haar-random unitary $U$. For m<q{m}<{q}, the null space is necessarily nonzero and in random orientation. Accordingly, the probability that all vectors of $\mathrm{null}(R)$ have the same component vanish is essentially zero, such that no features are isolatable. Once ${m}\geq{q}$, $R$ has full column rank, hence $\mathrm{null}(R)=0$ and the null-space condition in Eq. 66 is trivially fulfilled, implying that all q{q} features become instantly isolatable for a random $R$. In this case one may define $W=R^{+}=R^{\top}$ and assign $\mathbf{w}=\Pi_{r}W$ with a $\Pi_{r}$ that projects onto the $r$-th coordinate. This $W$ fulfills the slightly more general condition of finding a matrix $W\in\mathbb{R}^{{m}\times{m}}$ such that $WR=[\mathbbm{1}_{{m}},0_{{q}-{m}}]$, thereby isolating all m{m} features.  [^10] Writing the PTM as $R=[R_{1},R_{2}]$ with $R_{2}\in\mathbb{R}^{{m}\times({q}-{m})}$, this requires in general $R_{2}=0$, which is not fulfilled for typical (random) unitaries except if symmetries are present and the basis is ordered accordingly (e.g., the parity symmetry of the TFIM would allow a diagonal block structure of $R$ with m=q/2{m}={q}/2).
For a Clifford unitary, which maps between elements of the Pauli group, $\mathrm{row}(R)$ simply consists of all $\mathbf{e}_{j}$, hence $\mathrm{rank}R$ features are isolatable even if $R$ is wide. For unitaries based on structured Hamiltonians or at small evolution times, one may expect Eq. 66 to be fulfilled for a certain set of features, provided m{m} is sufficiently large (but can still be ${m}\ll{q}$). This is confirmed numerically in Fig. 13, where the behavior of the feature isolation criterion (66) is illustrated for various unitaries as a function of the number of observables m{m} (rows of $R$).
Appendix C Decodability scores and coefficient of determination
The decodability score $\gamma_{r}^{2}$ [Eq. 40] tells to which extent a feature can be geometrically reconstructed from the readout. It disregards statistical correlations and effects of noise, i.e., features with small singular values in the effective PTM $R$ could be considered decodable, although it would lead to huge weight norms. We present a refined $\mathcal{R}^{2}$-score below, but reserve a more detailed discussion of noise/regularization effects for a future work.
Consider the regression solution in Section IV.1, assume all quantities are centered and define the target alignment vector
$\mathbf{b}_{k}:=\mathbb{E}\!\left[\boldsymbol{\phi}(\mathbf{u})y_{k}(\mathbf{u})\right]\in\mathbb{R}^{q}.$ (68)
Using Eq. 26, the mean-squared error satisfies
$\mathbb{E}_{\mathbf{u}}\bigl[(y_{k}-f)^{2}\bigr]=\mathrm{Var}_{\mathbf{u}}(y_{k})-(R\mathbf{b}_{k})^{\top}\mathbf{w}^{*}.$ (69)
Plugging this into the definition in Eq. 43 gives the general expression for the coefficient of determination
$\mathcal{R}^{2}(k)=\frac{\mathbf{b}_{k}^{\top}R^{\top}(RCR^{\top})^{+}R\mathbf{b}_{k}}{\mathrm{Var}(y_{k})}.$ (70)
All quantities herein are to be averaged over the input distribution of $\mathbf{u}$. Notably, it is possible to have $\gamma_{k}^{2}=0$ (feature direction not present in row-space of $R$), yet still $\mathcal{R}^{2}(k)>0$ if $\phi_{r}$ is statistically correlated with other $\phi_{j}$ that are present. Hence, the above $\mathcal{R}^{2}$ generally measures predictability from correlations and not just isolatability of a feature.
The coefficient of determination $\mathcal{R}^{2}(k)$ can be related to the decodability score $\gamma_{r}^{2}$ [Eq. 40]: In the special where the targets are $y_{k}(\mathbf{u})=\phi_{k}(\mathbf{u})$, and under the additional assumption of orthonormality,
$\mathbb{E}[y_{k}y_{k^{\prime}}]=\delta_{kk^{\prime}}\implies C=\mathbbm{1}_{q},\quad\mathbf{b}_{k}=\mathbf{e}_{k},\quad\mathrm{Var}(y_{k})=1$ (71)
Eq. 70 reduces to the decodability score
$\mathcal{R}^{2}(k)=\mathbf{b}_{k}^{\top}R^{\top}(RR^{\top})^{+}R\mathbf{b}_{k}=\mathbf{e}_{k}^{\top}(R^{+}R)\mathbf{e}_{k}=\gamma_{k}^{2}.$ (72)
In the generic non-orthogonal case, the exact relation breaks down and $\mathcal{R}^{2}$ depends on the full projector $R^{+}R$ and not just on its diagonal elements. Then, $\mathcal{R}^{2}$ measures the extent to which the target lies in the decodable subspace.
Appendix D Pauli-Transfer Matrix Formalism: Further Details
We provide here some further details on the Pauli-Transfer Matrix (PTM) formalism for QELMs [33, 25, 46], including some geometrical intuitions [111] and discussion of noisy channels.
When ordering the basis with identity first, we have $\phi=(\phi_{0},\mathbf{s})$, where $\phi_{0}=\mathrm{tr}(\rho)=1$ and $\mathbf{s}\in\mathbb{R}^{d^{2}-1}$ is the traceless (Bloch) part. A maximally mixed state has $\mathbf{s}=0$, i.e., is in the center of the generalized Bloch sphere. By contrast, a pure state ($\mathrm{tr}(\rho^{2})=1$) has $|\mathbf{s}|^{2}=d-1$ and thus lies on the hypersphere of radius $\sqrt{d-1}$ (although not all points on it are valid quantum states).
D.1 Unitary channels
For a unitary channel $\mathcal{U}(\rho)=U\rho U^{\dagger}$, one has
$T_{U,\alpha\beta}=\frac{1}{2^{n}}\,\mathrm{tr}\big(P_{\alpha}\,UP_{\beta}U^{\dagger}\big),$ (73)
Trace preservation and unitality imply the form
$T_{U}=\begin{pmatrix}1&\mathbf{0}^{\top}\\ \mathbf{0}&\Omega_{U}\end{pmatrix}$, (74)
i.e., the traceless coefficients undergo a real orthogonal rotation with $\Omega_{U}\in\mathrm{SO}(d^{2}-1)$. For Clifford unitaries, $T_{U}$ is a signed permutation of Pauli strings.
Consider a separable reservoir Hamiltonian $H_{R}=\sum_{j=1}^{n}h_{R}^{(j)}$, such that the reservoir unitary factorizes as $U(t)=\bigotimes_{j=1}^{n}U_{j}(t)$. Such a structure emerges, e.g., in the limit of large transverse field $h/J\gg 1$ of the TFIM (5) This implies a Kronecker-product structure of the unitary PTM (73) (and similarly for any other separable quantum channel),
$T_{U}(t)=V^{(1)}(t)\otimes\cdots\otimes V^{(n)}(t),$ (75)
where each $V^{(j)}(t)\in\mathrm{SO}(4)$ acts on the single-qubit feature vector $\boldsymbol{\upphi}^{(j)}$. Consequently, the evolved features remain factorized:
$\boldsymbol{\upphi}^{\prime}(\mathbf{x})=T_{U}(t)\boldsymbol{\upphi}(\mathbf{x})=\bigl(V^{(1)}(t)\boldsymbol{\upphi}^{(1)}(x_{1})\bigr)\otimes\cdots\otimes\bigl(V^{(n)}(t)\boldsymbol{\upphi}^{(n)}(x_{n})\bigr).$ (76)
Thus, each Pauli-string feature $\phi_{a_{1}\ldots a_{n}}$ evolves only by local mixing among the single-qubit labels $a_{j}\in\{0,x,y,z\}$, without coupling different qubits and without operator spreading.
Each local unitary $U_{k}(t)=e^{-ith_{k}}$ rotates the local Pauli components, so $Y$-strings can become nonzero only through this local mixing. More explicitly, writing the local Pauli feature vector as $\tilde{\boldsymbol{\upphi}}^{(k)}\equiv(\phi_{x}^{(k)},\phi_{y}^{(k)},\phi_{z}^{(k)})^{\top}$, any single-qubit unitary induces a $\mathrm{SO}(3)$ rotation
$\tilde{\boldsymbol{\upphi}}^{{}^{\prime}(k)}(u_{k})=R^{(k)}(t)\,\tilde{\boldsymbol{\upphi}}^{(k)}(u_{k}),\qquad\phi_{0}^{{}^{\prime}(k)}=\phi_{0}^{(k)}=1,$ (77)
i.e., $\phi_{a}^{{}^{\prime}(k)}=\sum_{b\in\{x,y,z\}}R^{(k)}_{ab}(t)\,\phi_{b}^{(k)}$, with $R^{(k)}$ the adjoint representation of $U_{k}(t)$ on the Pauli vector. For example, for a local $Z$-field $h_{k}=Z$,
$U_{k}(t)^{\dagger}XU_{k}(t)=\cos(2t)\,X+\sin(2t)\,Y,\qquad U_{k}(t)^{\dagger}YU_{k}(t)=\cos(2t)\,Y-\sin(2t)\,X,$ (78)
so that $\phi_{y}^{{}^{\prime}(k)}(u_{k})=-\sin(2t)\,\phi_{x}^{(k)}(u_{k})$ and, correspondingly, $\phi^{\prime}_{\ldots y_{k}\ldots}(\mathbf{x})=-\sin(2t)\,\phi_{\ldots x_{k}\ldots}(\mathbf{x})$, while $Z$-features remain invariant. This shows how unitaries act to make specific features accessible to the readout when the latter uses a restricted set of observables.
D.2 General quantum channels: affine action on the Bloch part
Consider a channel $\mathcal{E}$ that is completely positive and trace-preserving (CPTP) and has associated Kraus operators $\{K_{\ell}\}$. Then
$T_{\mathcal{E},\alpha\beta}=\frac{1}{2^{n}}\sum_{\ell}\mathrm{tr}\big(P_{\alpha}K_{\ell}P_{\beta}K_{\ell}^{\dagger}\big),$ (79)
which corresponds in block form (compatible with $\phi=(1,\mathbf{s})$) to
$T_{\mathcal{E}}=\begin{pmatrix}1&\mathbf{0}^{\top}\\ \mathbf{t}&\Omega\end{pmatrix},\qquad\mathbf{s}^{\prime}=\mathbf{t}+\Omega\mathbf{s},$ (80)
where $\mathbf{t}=\mathbf{0}$ for unital channels and $\Omega$ is a contracting map (singular values $\leq 1$). Such a map can improve numerical conditioning by damping high-weight Pauli components that are often noise sensitive. Non-unitality produces a translation $\mathbf{t}$ of the Bloch part, e.g., a maximally mixed state would be shifted away from the center of the Bloch sphere. This becomes a learnable bias once combined with the identity component of an observable.
Consider the following single-qubit PTMs for non-unitary noise channels [112, 113, 114]:
  • Isotropic depolarization $\mathcal{D}_{\lambda}(\rho)=\lambda\rho+(1-\lambda)\mathbb{I}/2$:
    $T_{\mathrm{dep}}(\lambda)=\mathrm{diag}(1,\lambda,\lambda,\lambda).$ (81)
    For an $n$-qubit isotropic depolarizing channel on dimension $d=2^{n}$, $T_{\mathrm{dep}}^{(n)}(\lambda)=\mathrm{diag}\big(1,\lambda I_{d^{2}-1}\big)$.
  • $Z$-dephasing with probability $p$:
    $T_{\mathrm{dephZ}}(p)=\mathrm{diag}(1,1-2p,1-2p,1).$ (82)
  • Amplitude damping with parameter $\gamma\in[0,1]$ and polarization $z_{*}\in[-1,1]$ (with $z_{*}=1$ corresponding to the zero-temperature fixed point $|0\rangle\langle 0|$):
    $T_{\mathrm{AD}}(\gamma)=\begin{pmatrix}1&0&0&0\\ 0&\sqrt{1-\gamma}&0&0\\ 0&0&\sqrt{1-\gamma}&0\\ z_{*}\gamma&0&0&1-\gamma\end{pmatrix}$. (83)
Modeling $n$-qubit noise as tensor products of single-qubit channels $\mathcal{E}^{(n)}=\bigotimes_{i=1}^{n}\mathcal{E}^{(i)}$ is often used as a baseline in QRC models [115, 18, 21, 115, 46]. Since the PTMs (81) and (82) are Pauli-diagonal, they do not lead to Pauli spreading and hence do not enlarge the accessible feature subspace. Amplitude damping [Eq. 83] produces Pauli-I/Z-mixing, which causes operator weight decay.
A particularly transparent characterization emerges if the noise channel is approximately diagonal. For instance, for the $n$-qubit depolarizing channel $\mathcal{D}^{(n)}=\otimes_{i=1}^{n}\mathcal{D}_{\lambda}$ [see Eq. 81], any Pauli string $P$ of weight $\nu(P)$ is an eigenoperator with
$\mathcal{D}^{(n)}(P)=\lambda^{\nu(P)}P$ (84)
Accordingly, high-weight Pauli strings are exponentially suppressed by near-diagonal noise channels. In conjunction with unitary-induced operator spreading, this can increase information loss under finite sampling. On the other hand, amplitude damping noise was observed to improve QRC performance as the associated superoperator spreads the state’s probability weight into more Pauli directions, thereby enriching the feature map [18]. This has been related to the ’coherence influx’ present for amplitude damping but not depolarization [43].

### v2

@@
 Quantum reservoir computers (QRCs) have emerged as a promising approach to quantum machine learning, since they utilize the natural dynamics of quantum systems for data processing and are simple to train.
-Here, we consider $n$-qubit quantum extreme learning machines (QELMs) with continuous-time reservoir dynamics.
-QELMs are memoryless QRCs capable of various ML tasks, including image classification and time series forecasting.
-We apply the Pauli transfer matrix (PTM) formalism to theoretically analyze the influence of encoding, reservoir dynamics, and measurement operations, including temporal multiplexing, on the QELM performance.
-This formalism makes explicit that the encoding determines the complete set of (nonlinear) features available to the QELM, while the quantum channels linearly transform these features before they are probed by the chosen measurement operators.
-Optimizing a QELM can therefore be cast as a decoding problem in which one shapes the channel-induced transformations such that task-relevant features become available to the regressor.
-The PTM formalism allows one to identify the classical representation of a QELM and thereby guide its design towards a given training objective.
+Here, we consider $n$-qubit quantum extreme learning machines (QELMs) with initial-state encoding and continuous-time reservoir dynamics.
+QELMs are memoryless QRCs capable of various ML tasks, such as image classification and time series forecasting.
+We apply the Pauli transfer matrix (PTM) formalism to theoretically analyze the influence of encoding, reservoir dynamics, and measurement operations (including temporal multiplexing) on the QELM performance.
+This formalism makes explicit that the encoding determines the complete set of (nonlinear) features available to this type of QELM, while the quantum channels linearly transform these features before they are probed by the chosen measurement operators.
+Optimizing such a QELM can therefore be cast as a decoding problem in which one shapes the channel-induced transformations such that task-relevant features become available to the regressor, effectively reversing the information scrambling of a unitary.
+We show how operator spreading under unitary evolution determines feature decodability, which underlies the nonlinear processing capacity of the reservoir.
+The PTM formalism yields a nonlinear vector (auto-)regression model as the classical representation of a QELM.
 As a specific application, we focus on learning nonlinear dynamical systems and show that a QELM trained on such trajectories learns a surrogate-approximation to the underlying flow map.
@@
 QRCs have enjoyed significant attention since their introduction [1], featuring applications in various domains, such as image classification [2, 3, 4, 5], time series forecasting [1, 6, 7, 8, 9, 10], fluid dynamics [11, 12], or post-processing of quantum experiments [13, 14, 15, 16].
-A major advantage of QRCs is their suitability for noisy intermediate-scale quantum (NISQ) hardware, primarily because the training process is restricted to the classical readout layer, thereby rendering the training problem convex and avoiding the barren plateau problem associated with variational quantum algorithms [17].
+A major advantage of QRCs is their suitability for noisy intermediate-scale quantum (NISQ) hardware, primarily because the training process is restricted to the classical readout layer, thereby rendering the training problem convex and avoiding the barren plateau problem associated with variational quantum algorithms [17, 18].
 QRCs are often resilient to, and can even benefit from, environmental noise [19, 20, 21, 22].
-The relevance of nonlinear input transformations has been analyzed for QRCs in [22, 23, 24], and universality conditions have been derived in [25, 26, 27].
+The relevance of nonlinear input transformations has been analyzed for QRCs in [23, 24, 25], and universality conditions have been derived in [26, 6, 27, 28, 29, 30].
 Notably, QRCs are closely related to quantum kernel methods [31]—a fact that can be utilized to find optimal measurement operators [32] and harness insights from the classical domain [33, 34, 35].
@@
 A QELM differs from a conventional quantum neural network with fixed internal parameters in two important ways: first, instead of gate-based architectures, it typically employs the natural dynamics of a physical quantum system (a reservoir).
-Second, instead of optimizing a measurement output for a certain task, one collects multiple unspecific measurements into a large readout vector and feeds it into a classical regression or classification algorithm.
+Second, instead of optimizing a measurement output for a certain task, one usually collects multiple unspecific measurements into a large readout vector and feeds it into a classical regression or classification algorithm.
 The main task in optimizing QELMs thus consists in tuning the quantum properties of the reservoir to maximize information processing and being able to properly decode relevant features from the measurements [41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 14].
 Identifying possible quantum advantages, e.g., related to the exponentially large Hilbert space into which data is encoded, is an active research topic [51, 37, 52, 44, 3, 48].
-However, the exponential concentration effect emerging for large qubit numbers is a phenomenon closely related to barren plateaus [17, 50] that can strongly reduce QELM performance [51].
-The present work aims at a interpretable framework that can explain why and which failures in QELMs occur and thereby support systematic design of QELMs.
+However, the exponential concentration effect emerging for large qubit numbers is a phenomenon conceptually related to barren plateaus and can strongly reduce QELM performance [53].
+The present study aims at an interpretable framework that can explain why and which failures in QELMs occur and thereby support their systematic design.
 To this end, we will also analyze well-known optimization tools like temporal multiplexing [1] or information processing capacity [54, 55] in the light of the PTM formalism.
 We use chaotic dynamical systems as benchmarks since they have a known ground-truth model and are in principle discoverable from training data [56].
-Our findings also hold implications for general QML models.
+The developed formalism is restricted to $n$-qubit QELMs in which the encoding is time-independent and separated from reservoir evolution, hence excluding models with dynamical input injection [57, 14] or generic continuous-variable (infinite-dimensional) encodings [58, 59, 24].
+Our findings also hold implications for other QML models that fall into this category.
 II Model
-We briefly review the formalism and specific implementation of a QELM used in this work [1, 33, 34].
-Let $\mathcal{M}(\mathcal{H})$ denote the space of $d^{2}$-dimensional Hermitian matrices on the $d$-dimensional Hilbert space $\mathcal{H}$, where $d=2^{n}$ in terms of the number of qubits $n$.
+We briefly present the formalism of QELMs and the specific implementation used in this work [1, 36, 37].
+Let $\mathcal{M}(\mathcal{H})$ denote the space of $d\times d$ Hermitian matrices (real vector space of dimension $d^{2}$) on the $d$-dimensional Hilbert space $\mathcal{H}$, where $d=2^{n}$ in terms of the number of qubits $n$.
 Given classical inputs $\mathbf{x}\in\mathbb{R}^{D}$, a QELM is defined by the output function
-$f(\mathbf{x})=\sum_{k=1}^{m}w_{k}\mathrm{tr}[M_{k}\rho(\mathbf{x})]$ (1)
+$f(\mathbf{x})=\sum_{k=1}^{m}w_{k}\mathrm{tr}[M_{k}\rho(\mathbf{x})],$ (1)
 where $\{M_{k}\}_{k=1}^{m}$ are a set of measurement operators in $\mathcal{M}(\mathcal{H})$, and $\rho(\mathbf{x})\in\mathcal{M}(\mathcal{H})$ is the density matrix of the quantum system, obtained after applying encoding and other quantum channels to an initial state, $\rho(\mathbf{x})=\mathcal{E}(\cdots\mathcal{E}^{\text{enc}}_{\mathbf{x}}(\rho_{0})\cdots)$.
 The trainable weights $w_{k}$ are determined by minimizing the empirical risk:
-$\min_{\mathbf{w}\in\mathbb{R}^{m}}\sum_{i=1}^{P}\mathcal{L}\left(f(\mathbf{x}_{i}),\,y_{i}\right)+\frac{\lambda}{2}\|\mathbf{w}\|^{2}_{2}$ (2)
+$\min_{\mathbf{w}\in\mathbb{R}^{m}}\sum_{i=1}^{P}\mathcal{L}\left(f(\mathbf{x}_{i}),\,y_{i}\right)+\frac{\lambda}{2}\|\mathbf{w}\|^{2}_{2},$ (2)
 […]
-$\mathbf{w}^{*}=(\boldsymbol{\Psi}\boldsymbol{\Psi}^{\top}+\lambda\mathbbm{1}_{m})^{-1}\boldsymbol{\Psi}\mathbf{y},\qquad\Psi_{ki}=\mathrm{tr}(M_{k}\rho(\mathbf{x}_{i}))$ (3)
+$\hat{\mathbf{w}}=(\boldsymbol{\Psi}\boldsymbol{\Psi}^{\top}+\lambda\mathbbm{1}_{m})^{-1}\boldsymbol{\Psi}\mathbf{y},\qquad\Psi_{ki}=\mathrm{tr}(M_{k}\rho(\mathbf{x}_{i})),$ (3)
 where $\mathbf{y}\in\mathbb{R}^{P}$ is the sample vector and $\boldsymbol{\Psi}\in\mathbb{R}^{{m}\times P}$ is the readout design matrix.
-When optimizing over a complete basis of measurement operators $\{M_{k}\}_{k=1}^{d}$, the QELM takes the form of a quantum kernel model [28, 29] (see Appendix A for a summary).
+When optimizing over a complete basis of measurement operators $\{M_{k}\}_{k=1}^{d^{2}}$, the QELM takes the form of a quantum kernel model [31, 32] (see Appendix A for a summary).
 In the present work, we do not use that formalism, but instead work with a preselected observable set and optimize only over the associated weights $w_{k}$.
@@
 $|\psi(\mathbf{x})\rangle=U_{R}(t)S(\mathbf{x})|\Phi_{0}\rangle.$ (4)
-Here, $S(\mathbf{x})$ is a data encoding unitary, $U_{R}$ is a reservoir unitary, and the initial state $|\Phi_{0}\rangle$ is taken to be the ground state $|\Phi_{0}\rangle=|0\rangle^{\otimes n}$.
+Here, $S(\mathbf{x})$ is a data encoding unitary, $U_{R}$ is a reservoir unitary, and we take the initial state as $|\Phi_{0}\rangle=|0\rangle^{\otimes n}$.
 The unitary $U_{R}(t)=\exp(-\mathrm{i}tH_{R})$ is defined by time evolution under a reservoir Hamiltonian $H_{R}$.
-Specifically, we consider Haar-random Hamiltonians $H_{R}$ as well as two variants of a transverse field Ising model (TFIM)
+Specifically, we consider random Hamiltonians $H_{R}$ as well as two variants of a transverse field Ising model (TFIM)
 $H_{\text{TFIM}}^{\text{zz-x}}$ $=-J\sum_{j=1}^{n-1}Z_{j}Z_{j+1}-h\sum_{j=1}^{n}X_{j},$ (5a)
@@
 II.2 Encoding schemes
-We consider (parallel) product state encodings with $D=n$, producing the state
-$\rho_{\text{enc}}(\mathbf{x})=S(\mathbf{x})\rho_{0}S(\mathbf{x})^{\dagger}=\bigotimes_{j=1}^{n}|\psi_{\text{enc}}(x_{j})\rangle\langle\psi_{\text{enc}}(x_{j})|,\qquad S(\mathbf{x})=\bigotimes_{j=1}^{n}S_{j}(x_{j})$ (7)
+We consider time-independent (parallel) product state encodings with $D=n$, producing the state
+$\rho_{\text{enc}}(\mathbf{x})=S(\mathbf{x})\rho_{0}S(\mathbf{x})^{\dagger}=\bigotimes_{j=1}^{n}|\psi_{\text{enc}}(x_{j})\rangle\langle\psi_{\text{enc}}(x_{j})|,\qquad S(\mathbf{x})=\bigotimes_{j=1}^{n}S_{j}(x_{j}),$ (7)
 where the unitary $S_{j}$ act on the $j$-th qubit.
-Specifically, in our numerical experiments we mostly work with amplitude encoding [55].
-This scheme comes in two variants, which lead to different expressivities: one encodes data on half a great circle of the Bloch sphere
+In our numerical experiments we employ amplitude and rotational encoding schemes [60].
+The former comes in two variants, where one either encodes data on half a great circle of the Bloch sphere:
 $|\psi_{\text{enc}}(x)\rangle=\sqrt{x}|0\rangle+\sqrt{1-x}|1\rangle,\qquad x\in[0,1],$ (8)
-and one on the full great circle
+or on the full great circle:
 $|\psi_{\text{enc}}(x)\rangle=x|0\rangle+\sqrt{1-x^{2}}|1\rangle,\qquad x\in[-1,1].$ (9)
-In a few cases, we also consider rotational encoding [55], defined by the unitary
-$S_{j}(x_{j})=\exp(-i\pi x_{j}\sigma^{p_{j}})=R_{p_{j}}(2\pi x),\qquad x\in[0,1],$ (10)
+Rotational encoding is defined by the unitary
+$S_{j}(x_{j})=\exp(-i\pi x_{j}\sigma^{p_{j}})=R_{p_{j}}(2\pi x_{j}),\qquad x\in[0,1],$ (10)
 where $\sigma^{p}\in\{\mathbbm{1},X,Y,Z\}$ are single-qubit Pauli matrices, $p\in\{0,x,y,z\}$, and $R_{p_{j}}$ denotes rotation gates.
@@
 III Theory
-QML models, including and QELMs, have often been analyzed from the Fourier perspective, which is natural for Hamiltonian-based encoding [48, 51, 49].
-In the following, we instead employ the Pauli transfer matrix (PTM) formalism, which allows a more direct identification of the relevant features [33, 43, 56, 57, 58] and has been previously used for the analysis of QRCs [33, 25, 46].
+QML models, including QELMs, have often been analyzed from the Fourier perspective, which is natural for Hamiltonian-based encoding [51, 53, 52].
+In the following, we instead employ the Pauli transfer matrix (PTM) formalism, which allows a more direct identification of the relevant features for models with initial-state encoding [36, 46, 61, 62, 63] and has been previously used for the analysis of QRCs [36, 28, 49].
 III.1 Pauli-transfer formulation of a QELM
@@
 giving rise to a feature vector $\boldsymbol{\upphi}(\mathbf{x})\in\mathbb{R}^{d^{2}}$.
-For product state encodings, the components of $\phi_{k}(x)$ correspond to all possible $d^{2}$ monomial combinations of the single-Pauli features (see Section III.2).
+For product state encodings, the components of $\phi_{k}(\mathbf{x})$ correspond to all possible $d^{2}$ monomial combinations of the single-Pauli features (see Section III.2).
 Any quantum channel $\mathcal{E}$ acts linearly on this vector via the Pauli transfer matrix:
@@
 Consider the QELM in Eq.
-1 with a subset of measured Pauli operators $\{P_{k}\}_{k=1}^{{m}}$, ${m}\leq d^{2}$.
-This selection can be represented by a matrix $\mathcal{S}\in\mathbb{R}^{{m}\times d^{2}}$ (a subset of rows of the identity matrix) that extracts the corresponding rows from $T_{\mathcal{E}}$.
+1 with a subset $\mathcal{S}$ of measured Pauli operators $\{P_{k}\}_{k\in\mathcal{S}}$, ${m}\equiv|\mathcal{S}|\leq d^{2}$.
+This selection can be represented by a matrix $S\in\mathbb{R}^{{m}\times d^{2}}$ (a subset of rows of the identity matrix) that extracts the corresponding rows from $T_{\mathcal{E}}$.
 The readout vector $\mathbf{F}(\mathbf{x})\in\mathbb{R}^{{m}}$ contains their expectation values $F_{k}(\mathbf{x})=\mathrm{tr}\bigl[P_{k}\,\mathcal{E}(\rho_{\text{enc}}(\mathbf{x}))\bigr]$.
@@
 14 gives the model output function
-$f(\mathbf{x})=\mathbf{w}^{\top}\,\mathbf{F}(\mathbf{x})=\mathbf{w}^{\top}\,T_{\mathcal{E}}^{({m})}\,\boldsymbol{\upphi}(\mathbf{x})=\sum_{k=1}^{m}w_{k}\sum_{j=0}^{d^{2}-1}(\mathcal{S}T_{\mathcal{E}})_{kj}\phi_{j}(\mathbf{x}),$ (16)
-where $\mathbf{w}\in\mathbb{R}^{{m}}$ are classical weights and $T_{\mathcal{E}}^{({m})}\equiv\mathcal{S}T_{\mathcal{E}}$ is the ${m}\times d^{2}$ block of $T_{\mathcal{E}}$ restricted to the measured observables.
+$f(\mathbf{x})=\mathbf{w}^{\top}\,\mathbf{F}(\mathbf{x})=\mathbf{w}^{\top}\,T_{\mathcal{E}}^{({m})}\,\boldsymbol{\upphi}(\mathbf{x})=\sum_{k\in\mathcal{S}}w_{k}\sum_{j=0}^{d^{2}-1}(ST_{\mathcal{E}})_{kj}\phi_{j}(\mathbf{x}),$ (16)
+where $\mathbf{w}\in\mathbb{R}^{{m}}$ are classical weights and $T_{\mathcal{E}}^{({m})}\equiv ST_{\mathcal{E}}$ is the ${m}\times d^{2}$ block of $T_{\mathcal{E}}$ restricted to the measured observables.
 This formulation explicitly shows that $f(\mathbf{x})$ is a linear function of the features $\boldsymbol{\upphi}(\mathbf{x})$ generated solely by the encoding and mixed by the reservoir channel $T_{\mathcal{E}}$.
 With the Kraus representation (15), we can write
-$(T_{\mathcal{E}}^{({m})})_{kj}=\frac{1}{d}\sum_{\ell=1}^{\mu}\mathrm{tr}\bigl(P_{k}\,K_{\ell}P_{j}K_{\ell}^{\dagger}\bigr)=\mathrm{tr}\bigl(P_{k}^{(H)}P_{j}\bigr)$ (17)
+$(T_{\mathcal{E}}^{({m})})_{kj}=\frac{1}{d}\sum_{\ell=1}^{\mu}\mathrm{tr}\bigl(P_{k}\,K_{\ell}P_{j}K_{\ell}^{\dagger}\bigr)=\frac{1}{d}\mathrm{tr}\bigl(P_{k}^{(H)}P_{j}\bigr),$ (17)
 where the last expression introduces the generalized Heisenberg representation of an observable $O$,
-$O^{\text{(H)}}=\frac{1}{d}\sum_{\ell=1}^{\mu}K_{\ell}^{\dagger}OK_{\ell}.$ (18)
+$O^{\text{(H)}}=\sum_{\ell=1}^{\mu}K_{\ell}^{\dagger}OK_{\ell}.$ (18)
 Specifically, for the unitary channel, we have $K=U(t)$ ($\mu=1$), such that $T_{\mathcal{E}}$ becomes an orthogonal matrix $V(t)\in\mathrm{SO}(d^{2})$ with elements
@@
 with $\phi_{0}^{(j)}=\mathrm{tr}(\rho_{j})=1$ (normalization), $\sigma^{0}\equiv\mathbbm{1}$ and $\sigma^{x,y,z}\equiv X,Y,Z$ as before.
-The three non-trivial components of $\boldsymbol{\upphi}^{(j)}$ define the Bloch vector $\boldsymbol{\varphi}^{(j)}=\{\phi_{x}^{(j)},\phi_{y}^{(j)},\phi_{z}^{(j)}\}$, which fulfills $\|\boldsymbol{\varphi}^{(j)}\|^{2}=1$ for pure states (mainly considered here) and $\|\boldsymbol{\varphi}^{(j)}\|^{2}\leq 1$ for mixed states.
-Since both the state and the Pauli basis factorize across qubits, the $n$-qubit feature vector is simply
+The three non-trivial components of $\boldsymbol{\upphi}^{(j)}$ define the Bloch vector, which has unit magnitude for pure states (see Appendix D for further discussion).
+Since both the state and the Pauli basis factorize across qubits, the $n$-qubit feature vector is
 $\boldsymbol{\upphi}(\mathbf{x})=\bigotimes_{j=1}^{n}\boldsymbol{\upphi}^{(j)}(x_{j}),$ (22)
@@
 21: for each qubit one chooses one of the four factors $\{1,\phi^{(j)}_{x},\phi^{(j)}_{y},\phi^{(j)}_{z}\}$, yielding a total of at most $d^{2}=4^{n}$ features.
-If the measured observables are Pauli strings of product form $P=\bigotimes_{j=1}^{n}\sigma^{a_{j}}$, the corresponding readout feature factorizes analogously as $\phi^{\prime}_{a_{1}\ldots a_{n}}(\mathbf{x})=\prod_{j=1}^{n}\bigl(\boldsymbol{\upphi}^{{}^{\prime}(j)}(x_{j})\bigr)_{a_{j}}$ with the post-reservoir feature vector $\boldsymbol{\upphi}^{\prime}$.
 The readout features resulting from some simple encoding schemes are summarized in Table 1.
@@
 Feature $\phi_{a_{k}}^{(k)}(u_{k})$
-Factor in observable $P$ Amplitude on $[0,1]$ Rotational
-$\cdots\otimes X_{k}\otimes\cdots$ $2\sqrt{(1-u_{k})u_{k}}$ $\sin(\pi u_{k})$
-$\cdots\otimes Y_{k}\otimes\cdots$ 0 0
-$\cdots\otimes Z_{k}\otimes\cdots$ $2u_{k}-1$ $\cos(\pi u_{k})$
-$\cdots\otimes\tfrac{1}{2}(\mathbb{I}+Z_{k})\otimes\cdots$ $u_{k}$ $\cos^{2}(\pi u_{k}/2)$
-Table 1: Pauli–basis projections $\langle\psi_{\text{enc}}|P|\psi_{\text{enc}}\rangle$ of an encoded product state $\rho_{\text{enc}}(\mathbf{u})$ for rotational encoding (10) and amplitude encoding (8).
+Factor in observable $P$ Amplitude on $[0,1]$ Amplitude on [1,1][-1,1] Rotational
+$\cdots\otimes X_{k}\otimes\cdots$ $2\sqrt{(1-u_{k})u_{k}}$ $2u_{k}\sqrt{1-u_{k}^{2}}$ $\sin(2\pi u_{k})$
+$\cdots\otimes Y_{k}\otimes\cdots$ 0 0 0
+$\cdots\otimes Z_{k}\otimes\cdots$ $2u_{k}-1$ $2u_{k}^{2}-1$ $\cos(2\pi u_{k})$
+$\cdots\otimes\tfrac{1}{2}(\mathbb{I}+Z_{k})\otimes\cdots$ $u_{k}$ $u_{k}^{2}$ $\cos^{2}(\pi u_{k})$
+Table 1: Pauli–basis projections $\langle\psi_{\text{enc}}|P|\psi_{\text{enc}}\rangle$ of an encoded product state $\rho_{\text{enc}}(\mathbf{u})$ for amplitude encoding (8) or (9) as well as rotational encoding (10).
 Here $P=\bigotimes_{k=1}^{n}\sigma_{a_{k}}$ denotes a Pauli operator ($\sigma_{a_{k}}\in\{\mathbb{I},X,Y,Z\}$) and the classical datum entering qubit $k$ is $u_{k}$.
-Replacing $u_{k}\!\to\!u_{k}^{2}$ transforms the first column to the amplitude encoding variant (9).
 For convenience, we have included the projection operator $\frac{1}{2}(\mathbb{I}+Z_{k})$.
-One observes that each Pauli factor $\sigma_{a_{k}}$ (first column) in an observable $P$ produces the corresponding single-qubit feature $\phi_{a_{k}}^{(k)}(u_{k}):=\mathrm{tr}\!\big(\sigma_{a_{k}}\,\rho_{k}(u_{k})\big)$ (2nd and 3rd column).
+One observes that each Pauli factor $\sigma_{a_{k}}$ (1st column) in an observable $P$ produces the corresponding single-qubit feature $\phi_{a_{k}}^{(k)}(u_{k}):=\mathrm{tr}\!\big(\sigma_{a_{k}}\,\rho_{k}(u_{k})\big)$ (2nd–4th columns).
 For instance, the $X_{k}$ row gives $\phi_{x}^{(k)}(u_{k})$ and the $Z_{k}$ row gives $\phi_{z}^{(k)}(u_{k})$.
-IV PTM analysis for general quantum channels
-For separable quantum channels, the PTM factorizes as $T=\bigotimes_{j=1}^{n}T_{j}$ into single-qubit PTMs $T_{j}$.
+IV PTM analysis for a unitary quantum channel
+For separable (product) quantum channels, the PTM factorizes as $T=\bigotimes_{j=1}^{n}T_{j}$ into single-qubit PTMs $T_{j}$.
 The action of a local unitary is discussed in Section D.1.
@@
 16 and 17].
-The optimization can be split into several subtasks that define the fundamental reconstruction problem in (quantum) reservoir computing [42, 46, 59, 14]:
+We focus in the following on the unitary case.
+The optimization can be split into several subtasks that define the fundamental reconstruction problem in (quantum) reservoir computing with explicit feature maps [45, 49, 64, 14]:
   1. 1.
-    Decide which types of features $\phi_{r}$ are required to learn the dataset.
-For example, learning the flow map of a Markovian dynamical system typically requires low-order polynomial features.
-For harmonic time series and non-Markovian data, linear delay features are needed to capture temporal dependencies.
-Results like in Table 1 can aid systematic reservoir design.
+    Decide which types of features $\{\phi_{r}\}$ are best suited to learn the dataset.
+For example, non-Markovian time series (including harmonic signals) require delay coordinates [65, 66], whereas Markovian dynamical systems can already be modeled by instantaneous monomial features.
+While commonly used encodings support universal function approximation [26, 6, 27, 28, 29, 30], they can differ in their approximation errors, generalization behavior, or robustness during autoregressive rollout.
   2. 2.
-    Select suitable measurement observables $P_{k}$.
+    Select suitable measurement observables $\{P_{k}\}$.
 Ideally, the Heisenberg-evolved observables $P_{k}^{\text{(H)}}(t)$ have strong overlap with the observable $P_{r}$ generating the desired feature $\phi_{r}$ [see Eq.
@@
 We will show that temporal multiplexing directly enhances decodability by increasing the operator support.
-In temporal multiplexing [62, 1], instead of working at a fixed evolution time $t$ and a potentially large set of observables $P_{k}$, one selects a small subset $\mathcal{S}$ of size ${m}=|\mathcal{S}|<d^{2}$ and evaluate them at several times $t_{1},\ldots,t_{L}$.
-Correspondingly, we stack a subset $\mathcal{S}$ of rows of $V(t)$ into a ${m}{L}\times d^{2}$ observability matrix
-$R=\begin{pmatrix}\mathcal{S}V(t_{1})\\ \mathcal{S}V(t_{2})\\ \vdots\\ \mathcal{S}V(t_{L}).\end{pmatrix}$ (24)
-The case without temporal multiplexing corresponds to L=1{L}=1. For $\mathcal{S}=\mathbbm{1}_{d^{2}}$ and L=1{L}=1, one then recovers the full unital PTM $R=V(t)$.
+In temporal multiplexing [69, 1], instead of working at a fixed evolution time $t$ and a potentially large set of observables $P_{k}$, one selects a small subset $\mathcal{S}$ of size m{m} and evaluates them at several times $t_{1},\ldots,t_{L}$ [cf.
+Eq.
+16].
+One then stacks the corresponding rows of $V(t)$ into a ${m}{L}\times d^{2}$ observability matrix
+$R=\begin{pmatrix}SV(t_{1})\\ SV(t_{2})\\ \vdots\\ SV(t_{L})\end{pmatrix}$. (24)
+The case without temporal multiplexing corresponds to L=1{L}=1. The full unitary PTM $R=V(t)$ is then recovered for $S=\mathbbm{1}_{d^{2}}$.
 IV.1 Classical representation
-According to Eq. 16, a QELM has an equivalent classical representation as a PTM-weighted nonlinear function library,
-$f(\mathbf{u})=\mathbf{w}^{\top}R\boldsymbol{\upphi}(\mathbf{u})=\mathbf{w}^{\top}\mathbf{F}(\mathbf{u})$ (25)
+Let ${q}\leq d^{2}$ be the effective number of features (excluding features identically vanishing due to the encoding), and B{B} be the total number of measurements, e.g.
+B=mL{B}={m}{L}.
+According to Eq.
+16, a QELM has an equivalent classical representation as a PTM-weighted nonlinear function library,
+$f(\mathbf{u})=\mathbf{w}^{\top}\mathbf{F}(\mathbf{u}),\qquad\mathbf{F}(\mathbf{u})\equiv R\boldsymbol{\upphi}(\mathbf{u}),$ (25)
 where $R\in\mathbb{R}^{{B}\times{q}}$ is the effective PTM (24) (including any row-selection and temporal multiplexing) and $F_{k}=R_{kj}\phi_{j}$ are the Pauli features available at the readout.
-We denote the effective number of features by ${q}\leq d^{2}$, excluding features identically vanishing due to the encoding, while B{B} is the total number of measurements, e.g.
-B=mL{B}={m}{L}.
 If $R$ has full column rank, it is left-invertible  [^1] This allows for the existence of a pseudoinverse such that $R^{+}R=\mathbbm{1}_{q}$.
@@
 3] contain projections from the Pauli feature space to the data space:
-$\mathbf{w}^{*}=(GG^{\top}+\lambda\mathbbm{1}_{B})^{-1}G\mathbf{y}=R(CR^{\top}R+\lambda\mathbbm{1}_{q})^{-1}\Phi\mathbf{y},$ (26)
+$\hat{\mathbf{w}}=(GG^{\top}+\lambda\mathbbm{1}_{B})^{-1}G\mathbf{y}=R(CR^{\top}R+\lambda\mathbbm{1}_{q})^{-1}\Phi\mathbf{y},$ (26)
 where we introduced the feature design and Gram matrices
@@
 and the corresponding readout design and Gram matrices
-$G\equiv[\mathbf{F}(\mathbf{u}_{1})\cdots\mathbf{F}(\mathbf{u}_{P})]=R\Phi\in\mathbb{R}^{{B}\times P},\quad GG^{\top}=RCR^{\top}\in\mathbb{R}^{{B}\times{B}}$ (28)
+$G\equiv[\mathbf{F}(\mathbf{u}_{1})\cdots\mathbf{F}(\mathbf{u}_{P})]=R\Phi\in\mathbb{R}^{{B}\times P},\quad GG^{\top}=RCR^{\top}\in\mathbb{R}^{{B}\times{B}}.$ (28)
 The latter are typically computed in a QELM implementation, whereas $R$ and $\Phi$ are theoretical objects.
@@
 The last equality in (26) (which follows from the Woodbury identity) implies that for the full unitary PTM $R=V$ [Eq. 19], the weights are simply rotated versions of the ones without PTM mixing:
-$\mathbf{w}^{*}=V(C+\lambda\mathbbm{1}_{q})^{-1}\Phi\mathbf{y}.$ (29)
+$\hat{\mathbf{w}}=V(C+\lambda\mathbbm{1}_{q})^{-1}\Phi\mathbf{y}.$ (29)
 Moreover, since $V$ is orthogonal, it drops out from the predictor (25), consistent with the absence of a reservoir unitary in the kernel-representation (59).
 Such a QELM is then simply a regressor acting on the full $d^{2}$-dimensional feature library $\boldsymbol{\upphi}$.
-In the context of time series processing (especially when using some inputs $u_{j}$ to encode delay states) this is also known as a “next-generation” reservoir computer (NG-RC) [64] or, equivalently, a nonlinear vector autoregressive process (NVAR).
+In the context of time series processing (especially when using some inputs $u_{j}$ to encode delay states) this is reduces to a “next-generation” reservoir computer (NG-RC) [71] or, equivalently, a nonlinear vector autoregressive process (NVAR).
 When $R$ merely has full column rank, it can be interpreted to induce a regressor operating on the mixed feature set $\boldsymbol{\upphi}^{\prime}=R\boldsymbol{\upphi}$.
 A coarse measure for expressivity is the rank of the readout Gram matrix in Eq. 26, which can be bounded as
-$\mathrm{rank}(G)=\min(\mathrm{rank}(R),\mathrm{rank}(\Phi))\leq\min({B},P,{q}).$ (30)
-For sufficiently diverse data and large sample number $P$, $\Phi$ typically achieves a rank near its maximum $\min({q},P)$.
-However, it can suffer from severe ill-conditioning since the features $\phi_{k}$ are in general not orthogonal.
-Anticipating results discussed below (see Fig.
-4), for random Hamiltonians $\mathrm{rank}(R)$ typically grows monotonically with the measurement budget B{B} and saturates at $\mathrm{rank}(\Phi)\sim{q}$.
-By contrast, for structured Hamiltonians like the TFIM, $\mathrm{rank}(R)$ and thus $\mathrm{rank}(G)$ can saturate at a value significantly smaller than q{q}.
+$\mathrm{rank}(G)\leq\min(\mathrm{rank}(R),\mathrm{rank}(\Phi))\leq\min({B},P,{q}).$ (30)
+For sufficiently diverse data, $\Phi$ typically achieves a rank near its maximum $\min({q},P)$, but can suffer from ill-conditioning since the features $\phi_{k}$ are in general not orthogonal.
+Accordingly, for large sample number $P$, the rank of the readout approximately follows the rank of the effective PTM:
+$\mathrm{rank}(G)\simeq\mathrm{rank}(R).$ (31)
+For random Hamiltonians, $\mathrm{rank}(R)$ typically grows proportionally to the measurement budget B{B} and then saturates near q{q} (see Fig.
+4).
+By contrast, structured Hamiltonians like the TFIM are not fully scrambling, resulting in $\mathrm{rank}(R)\ll{q}$.
+The achieved ranks indicate the quality of the temporal multiplexing scheme and thus of the training accuracy of the QELM (see Fig.
+10).
 IV.2 Krylov growth and Pauli weight spreading
-Temporal multiplexing effectively “reverses” the information scrambling caused by quantum evolution.
+Temporal multiplexing enables the regressor to “reverse” the information scrambling caused by quantum evolution.
 The latter is captured by the Krylov growth of an operator under conjugation [72, 64, 67], which we summarize here in our setting.
@@
 18] of a given Pauli observable $P_{k}$ satisfies
-$P_{k}^{\text{(H)}}(t)\equiv U(t)^{\dagger}P_{k}U(t)=e^{iHt}P_{k}e^{-iHt}=\sum_{\ell=0}^{\infty}\frac{(it)^{\ell}}{\ell!}\,\mathrm{ad}_{H}^{\ell}(P_{k})$ (31)
+$P_{k}^{\text{(H)}}(t)\equiv U(t)^{\dagger}P_{k}U(t)=e^{iHt}P_{k}e^{-iHt}=\sum_{\ell=0}^{\infty}\frac{(it)^{\ell}}{\ell!}\,\mathrm{ad}_{H}^{\ell}(P_{k}),$ (32)
 where $\mathrm{ad}_{H}(O)\equiv[H,O]$ and $\mathrm{ad}_{H}^{\ell}$ denotes $\ell$ nested commutators with $H$. The time-evolved observable therefore spans the (increasing) Krylov subspace
@@
 5, the Hadamard equivalence (6) implies that
-$(Z_{j_{1}}\cdots Z_{j_{k}})^{\text{H}}(t)\big|_{\text{zz-x}}=U_{\text{H}}^{\dagger}(X_{j_{1}}\cdots X_{j_{k}})^{\text{H}}(t)\big|_{\text{xx-z}}U_{\text{H}}$ (33)
+$(Z_{j_{1}}\cdots Z_{j_{k}})^{\text{H}}(t)\big|_{\text{zz-x}}=U_{\text{H}}^{\dagger}(X_{j_{1}}\cdots X_{j_{k}})^{\text{H}}(t)\big|_{\text{xx-z}}U_{\text{H}},$ (34)
 i.e., Pauli-$Z$ observables evolve under $H_{\text{TFIM}}^{\text{zz-x}}$ equivalently to Pauli-$X$ observables under $H_{\text{TFIM}}^{\text{xx-z}}$ up to a global unitary.
@@
 In the PTM approach, Krylov growth can be represented in terms of operator spreading in the Pauli basis [73]. Accordingly, expanding the time-evolved observable $P_{k}^{\text{(H)}}(t)$ gives
-$P_{k}^{\text{(H)}}(t)=\sum_{j=0}^{d^{2}-1}V_{kj}(t)P_{j}$ (34)
+$P_{k}^{\text{(H)}}(t)=\sum_{j=0}^{d^{2}-1}V_{kj}(t)P_{j},$ (35)
 where we used cyclicity of the trace to introduce $V(t)$ from Eq.
 19.
-Thus, for non-Clifford $H$, an initial weight fully localized on $P_{k}$ spreads to other (higher and lower order) Pauli strings as $t$ increases.
+Thus, for non-Clifford $H$, an initial weight fully localized on $P_{k}$ can spread to other (higher and lower order) Pauli strings as $t$ increases.
 A convenient diagnostic is the Pauli (Hamming) weight $\nu(P_{j})\in\{0,\ldots,n\}$, i.e., the number of non-identity factors in $P_{j}$, and its coefficient-weighted average
 $\bar{\nu}_{k}(t)=\sum_{j=0}^{d^{2}-1}\nu(P_{j})\,V_{kj}(t)^{2}.$ (36)
-For local (including nonintegrable/chaotic ones) Hamiltonians, the Lieb–Robinson bound [67, 68, 69, 70, 47] implies that an initially local operator can only develop support within distance $\lesssim v_{\mathrm{LR}}t$ until finite-size saturation:
+For local (including nonintegrable/chaotic ones) Hamiltonians, the Lieb–Robinson bound [74, 75, 76, 77, 50] heuristically implies that an initially local operator can only develop support within distance $\lesssim v_{\mathrm{LR}}t$ until finite-size saturation:
 $\bar{\nu}_{k}(t)\sim\begin{cases}1+\mathrm{const}\cdot v_{\mathrm{LR}}t&\text{(intermediate times)}\\ \mathcal{O}(n)&\text{(late times)}\end{cases}$ (37)
 with a typical late-time value $\bar{\nu}_{k}\approx\tfrac{3}{4}n$, corresponding to a uniform distribution of coefficients $V_{kj}^{2}$ over Pauli strings.
-From the definition (18) and the expansion (34), it straightforwardly follows that $\|P_{k}^{(H)}\|=d=d\sum_{j=0}^{d^{2}-1}\|V_{kj}\|^{2}$, and thus $\sum_{j=0}^{d^{2}-1}V_{kj}^{2}=1$ for all $t$.
+From the definition (18) and the expansion (35), it straightforwardly follows that $\|P_{k}^{(H)}\|^{2}=d=d\sum_{j=0}^{d^{2}-1}V_{kj}^{2}$, and thus $\sum_{j=0}^{d^{2}-1}V_{kj}^{2}=1$ for all $t$.
 Consequently, for Haar-random unitaries, one expects a typical late-time magnitude
-$|V_{kj}(t\to\infty)|=2^{-n}$ (37)
-This is the origin of the exponential concentration effect, where the output of the QELM essentially looses any information about the input for large qubit numbers $n\to\infty$ [51].
+$|V_{kj}(t\to\infty)|=2^{-n}.$ (38)
+This is the origin of the exponential concentration effect, which makes accessing information about the input infeasible for large qubit numbers $n\to\infty$ [53].
 Integrable models (such as the TFIM at its integrable points) typically show more structured behavior due to stable quasiparticle modes.
-An extreme case is a Clifford unitary, for which $V_{kj}=\mathrm{const}.$ By contrast, a Haar-random unitary on the full Hilbert space can generate rapid delocalization [71, 72, 73, 74, 75].
+An extreme case is a Clifford unitary, for which $V_{kj}$ is a signed permutation matrix (no spreading).
+By contrast, a Haar-random unitary on the full Hilbert space can generate rapid delocalization [78, 79, 80, 81, 82].
 The evolution of the Pauli weight average is illustrated in Fig.
@@
 With increasing evolution time (from top to bottom), these initial operators develop nonzero projection onto the full $d^{2}$-dimensional Pauli basis [Eq.
-34].
+35].
 In the TFIM case, the supporting basis exhibits a sparse pattern, whereas it is dense for a random unitary.
@@
 Following Eq.
-25, we denote the set of Pauli features available at the readout by $F_{k}=R_{kj}\phi_{j}$ ($k=1,\ldots,{m}$), which emerge by linear transformations of the input features $\phi_{j}$ ($j=1,\ldots,{q}\leq d^{2}$) via the effective PTM $R\in\mathbb{R}^{{m}\times{q}}$ [Eq.
+25, we denote the set of Pauli features available at the readout by $F_{k}=R_{kj}\phi_{j}$ ($k=1,\ldots,{B}$), which emerge by linear transformations of the input features $\phi_{j}$ ($j=1,\ldots,{q}\leq d^{2}$) via the effective PTM $R\in\mathbb{R}^{{B}\times{q}}$ [Eq.
 24], possibly including temporal multiplexing.
-We now ask for conditions that enable the regressor to isolate a given input feature $\phi_{r}(x)$ by finding weights $\mathbf{w}$ such that (approximately)
-$\mathbf{w}\cdot\mathbf{F}\propto\phi_{r}\quad\Leftrightarrow\quad\mathbf{w}^{\top}R=\omega\mathbf{e}_{r}^{\top}$ (38)
+We now ask for conditions that enable the regressor to isolate a given input feature $\phi_{r}(\mathbf{x})$ by finding weights $\mathbf{w}$ such that (approximately)
+$\mathbf{w}\cdot\mathbf{F}\propto\phi_{r}\quad\Leftrightarrow\quad\mathbf{w}^{\top}R=\omega\mathbf{e}_{r}^{\top},$ (39)
 where $\omega$ is a free parameter and $\mathbf{e}_{r}$ is the $r$-th basis vector of $\mathbb{R}^{{q}}$.
-If $R$ corresponds to the full unital channel, then $R=V(t)\in\mathbb{R}^{{q}\times{q}}$ is invertible and any feature can be decoded by
-$\mathbf{w}=\omega V(t)\mathbf{e}_{r}.\qquad\text{(unital channel)}$ (39)
+If $R$ corresponds to the full unitary channel, then $R=V(t)\in\mathbb{R}^{{q}\times{q}}$ is invertible and any feature can be decoded by
+$\mathbf{w}=\omega V(t)\mathbf{e}_{r}.\qquad\text{(unitary channel)}$ (40)
 If we select subsets of observables, then Eq.
@@
 The exact condition derived in Appendix B provides a Pauli feature decodability score:
-$\gamma_{r}^{2}\equiv(R^{+}R)_{rr},\qquad\sum_{r=1}^{d^{2}}\gamma_{r}^{2}=\mathrm{rank}(R)\leq d^{2}$ (40)
+$\gamma_{r}^{2}\equiv(R^{+}R)_{rr},\qquad\sum_{r=1}^{{q}}\gamma_{r}^{2}=\mathrm{rank}(R)\leq{q},$ (41)
 and associated decoding weights
-$\mathbf{w}=\omega\left((R^{+})_{r,:}\right)^{\top}.$ (41)
+$\mathbf{w}=\omega\left((R^{+})_{r,:}\right)^{\top},$ (42)
+in terms of the Moore–Penrose pseudoinverse $R^{+}$.
 These coincide with the minimum-norm least-squares solution.
 Exact decodability of the $r$-th feature is indicated by $\gamma_{r}^{2}=1$.
-A values $\gamma_{r}^{2}<1$ implies that the feature can only be reconstructed up to an error $1-\gamma_{r}^{2}$.
+A value $\gamma_{r}^{2}<1$ implies that the feature can only be reconstructed up to an error $\sim\sqrt{1-\gamma_{r}^{2}}\|\phi\|_{2}$.
 We take $\gamma_{r}^{2}\gtrsim 0.5$ as a useful decodability threshold.
-Note that $\gamma_{r}^{2}$ is a geometric notion of feature decodability and ignores statistical correlations (see Appendix C for further discussion).
+Note that $\gamma_{r}^{2}$ is a geometric notion and ignores statistical correlations or noise floor (see Appendix C for further discussion).
 Both score and weights preserve tensor product structure of the general PTM $R$: if $R=\otimes_{i=1}^{n}R_{i}$ then $\gamma_{r}^{2}=\prod_{i=1}^{n}\gamma_{r,i}^{2}$ and $\mathbf{w}=\otimes_{i=1}^{n}\mathbf{w}_{i}$.
-For $R=\mathcal{S}V(t)$, one obtains $\gamma_{r}^{2}=\sum_{k\in\mathcal{S}}V(t)^{2}_{kr}$ showing that feature $r$ is decodable only to the extent that the “spread” of that Pauli under $V(t)$ lands in the measured subset $\mathcal{S}$.
+For $R=SV(t)$, one obtains $\gamma_{r}^{2}=\sum_{k\in\mathcal{S}}V(t)^{2}_{kr}$, showing that feature $r$ is decodable only to the extent that the “spread” of that Pauli under $V(t)$ lands in the measured subset $\mathcal{S}$.
 [^2] If $R$ contains a $t=0$ block, then since $V(0)=\mathbbm{1}_{d^{2}}$ the corresponding rows of $R$ include $\mathbf{e}_{j}^{\top}$ for all $j\in\mathcal{S}$ (up to basis ordering), hence $\mathbf{e}_{j}\in\mathrm{row}(R)$.
@@
 Specifically, for Haar-random unitaries, late-time decodability decays exponentially with qubit number $n$ [see Eq.
-37],
-$\gamma_{r}^{2}\big|_{t\to\infty}\sim|\mathcal{S}|4^{-n}$ (42)
-Noisy channels generically have invertible PTMs [see Section D.2] and hence do not strictly reduce the accessible feature subspace.
-However, they are contracting and can effectively suppress high-weight Pauli features [see Eq.
-84], requiring large weight norms $\|\mathbf{w}\|$ and thereby degrading conditioning.
-Such non-CPTP maps are generally unstable under finite sampling and thus effectively lead to information loss.
+38],
+$\gamma_{r}^{2}\big|_{t\to\infty}\sim|\mathcal{S}|4^{-n}.$ (43)
+Noisy channels can have invertible PTMs [see Section D.2] and then do not strictly reduce the accessible feature subspace.
+If they are contracting, they effectively suppress high-weight Pauli features [see Eq.
+81], but require large weight norms $\|\mathbf{w}\|$ which degrades conditioning.
+Such maps are unstable under finite sampling and thus effectively lead to information loss.
 Figure 3: Projector $R^{+}R$ of the effective PTM [Eq.
@@
 Figure 4: Feature decodability score $\bar{\gamma}_{j}^{2}$ obtained from the effective PTM $R$ [Eq.
-40], averaged over each Pauli weight sector, of a $n=3$ qubit model as a function of temporal multiplexing iteration length L{L} [Eq.
+41], averaged over each Pauli weight sector, of a $n=3$ qubit model as a function of temporal multiplexing iteration length L{L} [Eq.
 24].
@@
 Since we now start with $t=1$, the selected low-weight Paulis are already spread and initial decodability is low on average  [^3] Including $t\approx 0$ in the temporal multiplexing scheme improves decodability at early iterations, but does not affect the final decodability for sufficiently large L{L}..
-Compared to structured Hamiltonians, like the TFIM, random Hamiltonians lead to a consistently higher rank of $R$ and a more even decodability over Pauli sectors for large L{L}.
+Structured Hamiltonians, like the TFIM [Eq.
+5], can have strong impact on the achievable decodability score depending on the variant or the specific observables used.
+Comparing Fig.
+4(c) and (g), Pauli-$Z$’s evolving under the Hamiltonian $H_{\text{TFIM}}^{\text{xx-z}}$ render only weak scores compared to the Hamiltonian $H_{\text{TFIM}}^{\text{zz-x}}$  [^4] By global Hadamard equivalence, the statement flips for Pauli-$X$ as initial operators..
+A more detailed analysis using the Jordan-Wigner/Majorana formalism [86, 87] shows that, under $H_{\text{TFIM}}^{\text{xx-z}}$, the Pauli strings generated from an initial $Z$ are mainly consisting of $Z$-strings with XX/$Y$ endpoints, of which there are $\sim\mathcal{O}(n^{2})$.
+By contrast, spreading under $H_{\text{TFIM}}^{\text{zz-x}}$ almost fully saturates the Pauli basis ($\mathcal{O}(4^{n})$ strings, up to sub-exponential factors).
+Random Hamiltonians, by contrast, typically populate the full Pauli basis.
+Under temporal multiplexing they thus consistently lead to a higher rank of $R$ and a more even decodability over Pauli sectors for large L{L}.
 In general, we find that temporal multiplexing with Pauli $Z$ and $ZZ$ observables and random unitaries requires ${L}\sim d^{2}/|\mathcal{S}|$ measurements to achieve high decoding scores in all Pauli sectors, along with a saturated rank $\mathrm{rank}(R)\sim|\mathcal{S}|{L}\sim d^{2}$.
-In practice, however, this exponential measurement budget is usually not needed, since important features can be statistically decoded with much fewer measurements (see Figs.
+In practice, however, this exponential measurement budget is rarely needed, since relevant features can be statistically decoded with much fewer measurements (see Figs.
 6 and 7 below).
 Summary.
-The decodability framework is essentially a reformulation of pseudo-inverse solution of the RC training and provides a handle on the available features [78, 79, 80, 52].The decodability score (40) is a data-independent property of the chosen measurement observables and quantum operations after the encoding.
+The decodability framework reinterprets the pseudo-inverse solution of the RC training in terms of individually decodable features [88, 89, 90, 54].
+The decodability score (41) is a data-independent property of the chosen measurement observables and quantum operations after the encoding.
 For moderate qubit numbers $n$, full Pauli feature decodability is guaranteed when one either performs a full measurement (${m}=d^{2}$) on an invertible quantum channel, or one employs a temporal multiplexing scheme leading to a full column rank of $R$, again requiring an exponential measurement budget ${m}{L}\sim d^{2}$.
@@
 We focused above on the Pauli features $\phi_{j}$, which are generated by the encoding and transform linearly under the action of the QELM.
-However, they are themselves nonlinear functions of the classical input $u$.
-For example, with amplitude encoding (8), a single Pauli-$Z$ generates a linear readout $\langle Z\rangle\propto u$, a single Pauli-$X$ can introduce significant nonlinearities since $\langle X\rangle=\sqrt{1-(2u-1)^{2}}=1-\frac{1}{2}(2u-1)^{2}-\frac{1}{8}(2u-1)^{4}+\mathcal{O}(u^{6})$ (cf.
+However, they are themselves nonlinear functions of the classical input $\mathbf{u}\in\mathbb{R}^{D}$.
+For example, with amplitude encoding (8), a single Pauli-$Z$ generates a linear readout $\langle Z\rangle\propto u$, a single Pauli-$X$ can introduce significant nonlinearities since $\langle X\rangle=\sqrt{1-(2u-1)^{2}}=1-\frac{1}{2}(2u-1)^{2}-\frac{1}{8}(2u-1)^{4}+\ldots$ (cf.
 Table 1).
 It is therefore important to understand the actual nonlinearities in $u$ present at the readout.
-This is a classical problem in reservoir computing, studied initially as memory capacity [80] and later generalized towards information processing capacity [52, 53, 81, 44].
+This is a classical problem in reservoir computing, studied initially as memory capacity [90] and later generalized in terms of information processing capacity [54, 91, 55, 47].
 We discuss this, as well as an alternative approach, next.
@@
 Raw monomials like $\{u^{j}\}$ are strongly correlated under typical distributions of $u$ and thus do not allow for clear identification of the nonlinear degree within this statistical approach.
-Following [52, 44], we therefore use orthogonal polynomials $p_{j}(u_{\alpha})$ (degree $j$) to construct orthogonal degree-$k$ targets as $y_{k}(\mathbf{u})=\prod_{i=1}^{n}p_{e_{i}}(\mathbf{u})$, with $\sum_{i=1}^{n}e_{i}=k$.
+Following [54, 47], we therefore use orthogonal polynomials $p_{j}(u_{\alpha})$ (degree $j$) to construct orthogonal degree-$k$ targets as $y_{\mathbf{e}}(\mathbf{u})=\prod_{\alpha=1}^{D}p_{e_{\alpha}}(u_{\alpha})$, with $\sum_{\alpha=1}^{D}e_{\alpha}=k$.
 Specifically, for amplitude encoding, we use (shifted) Legendre polynomials, which are orthonormal w.r.t.
 the uniform distribution [1,1][-1,1] or $[0,1]$; for angle encoding, we use Hermite polynomials, which are orthonormal w.r.t.
-the Gaussian distribution $\mathcal{N}(0,1)$.
+the Gaussian distribution $\mathcal{N}(0,1)$  [^5] Due to periodicity of rotational encoding [Eq.
+10], the achievable capacity can be slightly reduced when sampling $u\sim\mathcal{N}$..
 (a)Refer to caption  (b)Refer to caption
 Figure 5: Ideal nonlinear capacity score $\mathcal{R}^{2}(k)$ [Eq.
-43] for amplitude and rotational encodings [Eqs.
-8 and 10], indicating the ability of an ideal QELM $f(\mathbf{u})=\mathbf{w}^{\top}\boldsymbol{\upphi}(\mathbf{u})$ to generate degree-$k$ monomials of the input $u_{\alpha}$ ($\alpha=1,\ldots,n$).
+44] for amplitude and rotational encodings [Eqs.
+8 and 10], indicating the ability of an ideal QELM $f(\mathbf{u})=\mathbf{w}^{\top}\boldsymbol{\upphi}(\mathbf{u})$ to generate degree-$k$ monomials of the input $u_{\alpha}$ ($\alpha=1,\ldots,D$ and $n=D$).
 Unless indicated, only Pauli-$Z$ features are measured in (a).
@@
 The dash-dotted curve represents the theoretical score for $\mathcal{R}^{2}_{Z}(k)$ with amplitude encoding and $n=5$ input qubits [Eq.
-44], which agrees exactly with the numerical results.
+45], which agrees exactly with the numerical results.
 $\Sigma$ denotes the integrated capacity over all $k$.
-The curves drop with increasing $k$ is because the quantum model can only generate square-free monomials of the Pauli features $\phi_{j}$.
 We take the coefficient of determination [54, 55]
-$\mathcal{R}^{2}(k)=1-\frac{\mathbb{E}_{\mathbf{u}}(y_{k}(\mathbf{u})-\hat{f}(\mathbf{u}))^{2}}{\mathrm{Var}_{\mathbf{u}}y_{k}(\mathbf{u})},$ (43)
-to indicate whether the trained predictor $\hat{f}(\mathbf{u})$ can generate nonlinearities of degree $k$ in the input $u$  [^4] In Eq.
-43, $\mathrm{Var}_{\mathbf{u}}y(\mathbf{u})$ denotes the variance of $y(\mathbf{u})$ over the distribution of $\mathbf{u}$..
+$\mathcal{R}^{2}(k)=\frac{1}{N_{k}}\sum_{|\mathbf{e}|=k}\mathcal{R}^{2}[y_{\mathbf{e}}],\qquad\mathcal{R}^{2}[y_{\mathbf{e}}]=1-\frac{\mathbb{E}_{\mathbf{u}}(y_{\mathbf{e}}(\mathbf{u})-\hat{f}(\mathbf{u}))^{2}}{\mathrm{Var}_{\mathbf{u}}y_{\mathbf{e}}(\mathbf{u})},$ (44)
+averaged over all $N_{k}=\binom{D+k-1}{k}$ degree-$k$ monomials, to indicate whether the trained predictor $\hat{f}(\mathbf{u})$ can generate nonlinearities of degree $k$ in the $(D=n)$-dimensional input  [^6] In Eq.
+44, $\mathrm{Var}_{\mathbf{u}}y(\mathbf{u})$ denotes the variance of $y(\mathbf{u})$ over the distribution of $\mathbf{u}$..
 $\mathcal{R}^{2}$ generalizes the geometric score (41) by taking statistical correlations into account (see Appendix C).
-In fact, for whitened orthonormal targets, one has $\mathcal{R}^{2}(k)=\gamma_{k}^{2}$.
-In the case of amplitude encoding (8) with Pauli-$Z$ readout, the number of representable degree-$k$ targets is $\binom{n}{k}$, which exponentially decreases compared to the total number of degree-$k$ monomials $\binom{n+k-1}{k}$, giving an expected score
-$\mathcal{R}^{2}_{Z}(k)=\begin{cases}\binom{n}{k}/\binom{n+k-1}{k}&\text{for }k\leq n\\ 0&\text{for }k>n\end{cases}$ (44)
+In the special case where the targets are whitened orthonormal features, one has $\mathcal{R}^{2}[y_{r}]=\gamma_{r}^{2}$.
+In the case of amplitude encoding (8) with Pauli-$Z$ readout, the number of representable degree-$k$ targets is $\binom{D}{k}$, giving an averaged score
+$\mathcal{R}^{2}_{Z}(k)=\begin{cases}\binom{D}{k}/\binom{D+k-1}{k}&\text{for }k\leq D,\\ 0&\text{for }k>D\end{cases}$ (45)
+which exponentially decreases with $k$.
 In Fig.
@@
 5(b) we show the nonlinear capacity score $\mathcal{R}^{2}(k)$ obtained for a complete Pauli observable set.
-Notably, the presence of non-$Z$ Paulis significantly enhances the nonlinearity present in the readout, as becomes clear by comparison to the dash-dotted curve corresponding to $\mathcal{R}^{2}_{Z}(k)$ for $n=5$ [Eq.
-44].
+Notably, the presence of non-$Z$ Paulis significantly enhances the nonlinearity present in the readout, as becomes clear by comparison to the dash-dotted curve corresponding to $\mathcal{R}^{2}_{Z}(k)$ for $D=n=5$ [Eq.
+45].
 Moreover, angle and amplitude encodings show a similar behavior.
-The decay of $\mathcal{R}^{2}(k)$ with $k$ stems from the fact that the encodings used here can only generate square-free monomials of the Pauli features $\phi_{j}$, while we test against (the practically relevant case of) all possible degree-$k$ targets.
+We attribute the decay of $\mathcal{R}^{2}(k)$ with $k$ mainly to fact that the product state encoding only generates square-free monomials of the Pauli features $\phi_{j}$, while we test against (the practically relevant case of) all possible degree-$k$ targets  [^7] This argument is not fully rigorous, since features like $\phi_{x}^{(i)}(u_{i})$ have infinite Taylor expansion in $u_{i}$ and can thus correlate locally with raw monomials $u_{i}^{m}$..
 (a)Refer to caption  (b)Refer to caption
@@
 Figure 6: Nonlinear capacity score $\mathcal{R}^{2}(k)$ [Eq.
-43] as a function of the nonlinear order $k$ and the temporal multiplexing iteration length L{L} [Eq.
+44] as a function of the nonlinear order $k$ and the temporal multiplexing iteration length L{L} [Eq.
 24].
@@
 Figure 7: Relative nonlinear capacity score $\mathcal{R}^{2}(k)/(\mathcal{R}^{2}_{\text{all}}(k)+\delta)$ [Eq.
-43] as a function of the nonlinear order $k$ and the temporal multiplexing iteration length L{L} [Eq.
+44] as a function of the nonlinear order $k$ and the temporal multiplexing iteration length L{L} [Eq.
 24].
@@
 (a) Decodability score $\kappa_{j}^{2}$ according to Eq.
-45.
+46.
 (b) Projector matrix $A^{+}A$, where nonzero off-diagonals correspond to $\kappa_{j}^{2}<1$ and indicate decoding is possible only within a certain accuracy.
 The underlying Taylor expansion is performed up to order $r=4$ in each $u_{j}=\{x,y,z\}$ and changes are mild upon further increasing $r$.
-The scores and matrices are averaged over a uniform grid of expansion points $\mathbf{u}_{0}\in[0.1,0.9]^{3}$ for amplitude encoding and $\mathbf{u}_{0}\in[-1,1]^{3}$ for rotational encoding.
+The scores and matrices are averaged over a uniform grid of expansion points $\mathbf{u}_{0}\in[0.1,0.9]^{3}$ for amplitude encoding (with some margin to the singular boundaries) and $\mathbf{u}_{0}\in[0,1]^{3}$ for rotational encoding.
 For readability, monomial labels are incompletely displayed in (a) and replaced by column/row indices in (b).
 Raw monomials in $u_{\alpha}$ are often relevant features for practical applications, such as autoregressive learning of dynamical systems.
-Let thus $\mathcal{B}(\mathbf{u})=\{1,u_{1},\ldots,u_{1}u_{2},\ldots,\mathcal{O}(u^{r})\}$ be the vector of size $b=\binom{D+r}{r}$ of monomials of $\{u_{\alpha}\}_{1\leq\alpha\leq D}$ up to degree $r$.
-The order-$r$ Taylor expansion of the Pauli features $\{\phi_{j}(\mathbf{u})\}_{1\leq j\leq q}$ within a small region around some operating point $\mathbf{u}_{0}$ (e.g., data mean) can then be written as $\boldsymbol{\upphi}(\mathbf{u})=A\mathcal{B}(\mathbf{u})$ with a coefficient matrix $A\in\mathbb{R}^{q\times b}$.
+Let thus $\mathcal{B}(\mathbf{u})=\{1,u_{1},\ldots,u_{1}u_{2},\ldots,\prod_{\alpha}u_{\alpha}^{r}\}$ be the vector of size $b=(r+1)^{D}$ of monomials of $\{u_{\alpha}\}_{1\leq\alpha\leq D}$ up to degree $r$ (per coordinate).
+The order-$r$ Taylor expansion of the Pauli features $\{\phi_{j}(\mathbf{u})\}_{1\leq j\leq q}$ within a small region around some operating point $\mathbf{u}_{0}$ (e.g., data mean) can then be written as $\boldsymbol{\upphi}(\mathbf{u})\simeq A(\mathbf{u}_{0})\mathcal{B}(\mathbf{u}-\mathbf{u}_{0})$ with a coefficient matrix $A(\mathbf{u}_{0})\in\mathbb{R}^{q\times b}$.
 As before, we can define a geometric decodability score
@@
 This score depends on the operating point $\mathbf{u}_{0}$, but is independent of quantum properties of the reservoir, and ranges between 0 and 1 (no/exact construction possible).
-Intermediate values indicate that construction is possible only within an error $\sim\sqrt{1-\kappa_{j}^{2}}$ due to off-diagonal terms in the $j$-th row/column of the projector $A^{+}A$.
+Intermediate values indicate that construction is possible only within an error $\sim(1-\kappa_{j}^{2})^{1/2}$ due to off-diagonal terms in the $j$-th row/column of the projector $A^{+}A$.
 A somewhat artificial dependence of $\kappa_{j}^{2}$ on the Taylor order $r$ (owing to the rank constraint) can be mitigated by considering large $r$ and $b$.
@@
 These include monomials that arise directly from the tensor product structure of the encoding state.
-For amplitude encoding (8), these are all square-free monomials up to order $n$ (e.g., for $n=D=3$: $\mathcal{B}_{\text{stable}}=\{1,x,y,z,xy,xz,yz,xyz\}$).
+For amplitude encoding (8) including all $Z$-strings, these are all square-free monomials up to order $n$ (e.g., for $n=D=3$: $\mathcal{B}_{\text{stable}}=\{1,x,y,z,xy,xz,yz,xyz\}$).
 The typical behavior of $\kappa_{j}^{2}$ and of the projectors $A^{+}A$ (averaged over a range of expansion points) is illustrated in Fig.
@@
 This is associated with less leakage into off-diagonal terms of $A^{+}A$ [see Fig.
-8(b)], and can explain why rotational encoding often leads to robust performance of QELMs.
-IV.5 Summary
-Feature decodability is the underlying concept that determines the expressivity of a QELM.
-Pauli features are the fundamental “building blocks” of the QELM expressivity.
-Any other nonlinear feature required for the learning task must be constructible from Pauli features.
-For the encoding types considered here, expressing simple monomials is often only possible up to a certain approximation error ($\kappa_{j}^{2}<1$), which increases with the size of the data domain.
-Knowing the structure of the readout features allows to design feature engineering strategies to enhance the expressivity at low cost.
+8(b)], and contributes to the robust performance often observed for rotational encoding.
+IV.5 Summary and Discussion
+It is useful to briefly discuss the injectivity property of the feature map [49], which ensures that distinct inputs map to distinct readouts.
+A standard condition for local injectivity of $\mathbf{F}(\mathbf{u})=R\boldsymbol{\upphi}(\mathbf{u})$ [Eq.
+25] is a full-rank Jacobian:
+$\mathrm{rank}\,RJ_{\boldsymbol{\upphi}}(\mathbf{u})=D,\qquad J_{\boldsymbol{\upphi}}(\mathbf{u})=\frac{\partial\boldsymbol{\upphi}(\mathbf{u})}{\partial\mathbf{u}}\in\mathbb{R}^{{q}\times D}.$ (47)
+For the considered product state encodings it is straightforward to check that $\boldsymbol{\upphi}(\mathbf{u})$ is injective on the respective domains in $\mathbb{R}^{D}$ (possibly after removing one endpoint).
+Hence, local injectivity of $\mathbf{F}(\mathbf{u})$ is determined by whether $R$ is injective on the tangent space $\operatorname{range}\,J_{\boldsymbol{\upphi}}(\mathbf{u})$, i.e.
+whether $\ker R\cap\operatorname{im}J_{\boldsymbol{\upphi}}(\mathbf{u})=\{0\}$.
+In a generic setting, this condition is satisfied (at fixed $\mathbf{u}$) if $\mathrm{rank}\,R\geq D$, which is the case for adequate measurement schemes and sufficiently scrambling dynamics.
+Decodability of Pauli features can be understood as the underlying concept that determines the expressivity of a QELM with initial-state encoding and data-independent reservoir dynamics.
+High decodability scores are associated with high-rank effective PTMs $R$, which are achieved for random unitaries.
+Strongly mixing quantum channels can help expressivity and nonlinear processing in QELMs [44, 14, 50], and are thus often beneficial for training performance apart from concentration effects [53].
+By contrast, QRCs with internal memory states often work best at a point where random matrix behavior just begins to dominate (“edge of chaos”) (see, e.g., [41, 95, 96]).
+An alternative to increasing randomness is to devise task-aligned reservoir structures and measurement schemes [2, 43, 97, 50, 32].
+Knowing the connection between the readout and the encoding-generated features via the PTM framework can thus inform feature engineering strategies to enhance expressivity.
 V Interpretable learning of dynamical systems
@@
 The non-integrable nature and fractal attractors of chaotic systems make them natural benchmarks for system identification, which addresses the question to which extent the underlying data-generating process can be uniquely discovered from the data [98, 56].
-If the dynamics is Markovian, it can be modeled solely with a NG-RC or QELM based on an instantaneous feature map [84, 29]. Accordingly, we formulate the training problem of the QELM (25) as
-$\text{QELM:}\qquad\mathbf{u}_{t+1}=\mathbf{f}(\mathbf{u}_{t})=W^{\top}R\boldsymbol{\upphi}(\mathbf{u})$ (48)
+If the dynamics is Markovian, it can be modeled solely with a NG-RC or QELM based on an instantaneous feature map (see [32] for examples).
+Accordingly, we formulate the training problem of the QELM (25) as
+$\text{QELM:}\qquad\mathbf{u}_{t+1}=\mathbf{f}(\mathbf{u}_{t})=W^{\top}R\boldsymbol{\upphi}(\mathbf{u}_{t})$ (50)
 […]
@@
 The polynomial form of the vector field in (51) then suggests to use amplitude encoding (8) with $n=3$ qubits, leading to a library of the following 27 terms:
-$\boldsymbol{\upphi}_{\text{amp}}^{n=3}(\mathbf{u})=\Bigl\{$ $1,2\sqrt{(1-x)x},2x-1,2\sqrt{(1-y)y},2(2x-1)\sqrt{-((y-1)y)},4\sqrt{(x-1)x(y-1)y},2y-1,$ (50)
-$2\sqrt{-((x-1)x)}(2y-1),(2x-1)(2y-1),2\sqrt{(1-z)z},2(2x-1)\sqrt{-((z-1)z)},$ (51)
-$2(2y-1)\sqrt{-((z-1)z)},2(2x-1)(2y-1)\sqrt{-((z-1)z)},4\sqrt{(x-1)x(z-1)z},$ (52)
+$\boldsymbol{\upphi}_{\text{amp}}^{n=3}(\mathbf{u})=\Bigl\{$ $1,2\sqrt{(1-x)x},2x-1,2\sqrt{(1-y)y},2(2x-1)\sqrt{-((y-1)y)},4\sqrt{(x-1)x(y-1)y},2y-1,$
+$2\sqrt{-((x-1)x)}(2y-1),(2x-1)(2y-1),2\sqrt{(1-z)z},2(2x-1)\sqrt{-((z-1)z)},$
+$2(2y-1)\sqrt{-((z-1)z)},2(2x-1)(2y-1)\sqrt{-((z-1)z)},4\sqrt{(x-1)x(z-1)z},$
 $4(2y-1)\sqrt{(x-1)x(z-1)z},4\sqrt{(y-1)y(z-1)z},4(2x-1)\sqrt{(y-1)y(z-1)z},$ (52)
-$8\sqrt{-((x-1)x(y-1)y(z-1)z)},2z-1,2\sqrt{-((x-1)x)}(2z-1),(2x-1)(2z-1),$ (54)
-$2\sqrt{-((y-1)y)}(2z-1),2(2x-1)\sqrt{-((y-1)y)}(2z-1),4(2z-1)\sqrt{(x-1)x(y-1)y},$ (55)
-$(2y-1)(2z-1),2\sqrt{-((x-1)x)}(2y-1)(2z-1),(2x-1)(2y-1)(2z-1)\Bigr\}$ (56)
-We will also consider the amplitude encoding scheme of Eq.
-9, which is obtained by replacing in Eq.
-56 $\{x,y,z\}\to\{x^{2},y^{2},z^{2}\}$.
-Depending on the dynamical system, we will observe differences in the performance.
-In order to confine the trajectory to the interval $\mathbf{u}\in[0,1]^{3}$ or $1,1]^{3}$, as required by the encoding, we apply the transformation $\mathbf{u}\to\mathbf{u}^{\prime}=\boldsymbol{\alpha}\cdot(\mathbf{u}-\mathbf{m})$ with $\boldsymbol{\alpha},\mathbf{m}\in\mathbb{R}^{3}$ to the state.
+$8\sqrt{-((x-1)x(y-1)y(z-1)z)},2z-1,2\sqrt{-((x-1)x)}(2z-1),(2x-1)(2z-1),$
+$2\sqrt{-((y-1)y)}(2z-1),2(2x-1)\sqrt{-((y-1)y)}(2z-1),4(2z-1)\sqrt{(x-1)x(y-1)y},$
+$(2y-1)(2z-1),2\sqrt{-((x-1)x)}(2y-1)(2z-1),(2x-1)(2y-1)(2z-1)\Bigr\}$
+We will also consider the amplitude encoding scheme (9), which has a similar polynomial structure.
+In order to confine the trajectory to the interval $\mathbf{u}\in[0,1]^{3}$ or $1,1]^{3}$, as required by the encoding, we apply the transformation $\mathbf{u}\to\mathbf{u}^{\prime}=\boldsymbol{\alpha}\odot(\mathbf{u}-\mathbf{m})$ with $\boldsymbol{\alpha},\mathbf{m}\in\mathbb{R}^{3}$ to the state (we typically use the same $\alpha$ for all coordinates).
 This leads to the rescaled Lorenz-63 model:
@@
 $\dot{w}$ $=-aw-4\frac{\alpha_{z}}{\alpha_{x}}u-4\frac{\alpha_{z}}{\alpha_{y}}v-2\frac{\alpha_{z}m_{x}}{\alpha_{x}}u-\frac{\alpha_{z}}{\alpha_{x}^{2}}u^{2}+\alpha_{z}(-am_{z}-4m_{x}-4m_{y}-m_{x}^{2}).$
-This vector field-level transformation allows us to directly compare the learned coefficients with the true ones.
+This vector field-level transformation simplifies comparison between the learned coefficients and the true ones.
 V.2 Numerical results
@@
 In this setup, the reservoir unitary becomes immaterial [see Eq.
-29], except for possible numerical conditioning issues.
+29].
 We generate trajectories using a 5th-order Runge-Kutta integrator with a time step of $\Delta t=0.01$.
@@
 The RMS training error
-$\epsilon_{\text{train}}=\Big(\frac{1}{P}\sum_{t=1}^{P}\|\mathbf{u}_{t+1}-\hat{\mathbf{f}}(\mathbf{u}_{t})\|^{2}\Big)^{1/2},$ (59)
+$\epsilon_{\text{train}}=\Big(\frac{1}{P}\sum_{t=1}^{P}\|\mathbf{u}_{t+1}-\hat{\mathbf{f}}(\mathbf{u}_{t})\|^{2}\Big)^{1/2}$ (55)
 measures one-step accuracy of the trained predictor $\hat{\mathbf{f}}$.
-As a measure of multi-step prediction accuracy, the autonomously generated trajectory $\hat{\mathbf{u}}_{t+1}=\hat{\mathbf{f}}(\hat{\mathbf{u}}_{t})$ is compared to the true trajectory via the forecast horizon,
-$T_{\text{fch}}=\inf\left\{t:\|\mathbf{u}_{t}-\hat{\mathbf{u}}_{t}\|>\sigma\right\}$ (60)
+As a measure of multi-step prediction accuracy, the autonomously generated trajectory $\hat{\mathbf{u}}_{t+1}=\hat{\mathbf{f}}(\hat{\mathbf{u}}_{t})$ is compared to the true trajectory via the forecast horizon
+$T_{\text{fch}}/\Delta t=\inf\left\{t:\|\mathbf{u}_{t}-\hat{\mathbf{u}}_{t}\|>\sigma\right\},$ (56)
 where $\bar{\mathbf{u}}=\frac{1}{P}\sum_{t=1}^{P}\mathbf{u}_{t}$ and $\sigma^{2}=\frac{1}{P}\sum_{t=1}^{P}\|\bar{\mathbf{u}}-\mathbf{u}_{t}\|^{2}$ are the mean and variance of the trajectory, respectively.
@@
 Figure 9: QELM vs.
-its classical representation: Forecast obtained from a QELM [Eq.
-1, ‘QRC’] with amplitude encoding (8) and $n=3$ qubits, trained on the rescaled Lorenz-63 model.
-This is macthed exactly by the forecast of the iterated classical representation [Eq.
-25, ‘Surrogate’].
+its classical representation: The forecast obtained from a QELM [Eq.
+1] trained on the rescaled Lorenz-63 model is matched exactly by the forecast of its iterated classical (surrogate) representation [Eq.
+25].
 Both eventually diverge from the true trajectory due to the Lyapunov instability.
+The QELM employs amplitude encoding (8) with $n=3$ qubits.
 Ground-truth trajectories are generated with a resolution of $\Delta t=0.01$.
@@
 For the scaling transformation in Eq.
-57, we use parameters $\boldsymbol{\alpha}=0.004\cdot[1,1,1]$ and $\mathbf{m}=[-30,-30,-5]$.
+53, we use parameters $\alpha_{\kappa}=0.004$ and $\mathbf{m}=[-30,-30,-5]$.
 System type QELM setup $\alpha$ $\lambda$ $\epsilon_{\text{train}}$ $T_{\text{fch}}/T_{L}$
@@
 Table 2: Training RMS errors $\epsilon_{\text{train}}$ [Eq.
-59] and forecast horizons $T_{\text{fch}}$ [Eq.
-60] for the benchmark systems and (unitary-free) QELM setups discussed in the text.
+55] and forecast horizons $T_{\text{fch}}$ [Eq.
+56] for the benchmark systems and (unitary-free) QELM setups discussed in the text.
 We generally use $n=3$ input qubits (no delay states) and a time resolution of $\Delta t=0.01$.
@@
 The column ‘type’ indicates whether we use the standard [Eq.
-49] or rescaled dynamical system.
+51] or rescaled dynamical system.
 In the latter case, $\alpha$ gives the rescaling factor in Eqs.
-57 and 58 (same for all coordinates).
+53 and 54 (same for all coordinates).
 Results of training a QELM in various setups are summarized in Table 2.
@@
 Figure 10: Prediction performance (forecast horizon and RMS training error) under temporal multiplexing of a QELM trained on the Halvorsen model (54).
-We use amplitude encoding (9) with $n=3$ qubits, single-qubit Pauli-$Z$ observables, and (a) TFIM Hamiltonian (5a), (b) random Hamiltonian.
+We use amplitude encoding (9) with $n=3$ qubits, single-qubit Pauli-$Z$ observables, and (a) the TFIM Hamiltonian $H_{\text{TFIM}}^{\text{zz-x}}$ [Eq.
+5a] or (b) a random Hamiltonian.
 The top horizontal axis shows the numerically obtained rank of the readout matrix $G$ [Eq.
@@
 Forecast horizon is given in units of the Lyapunov time $T_{L}$; no additional features are added to the readout.
-When employing temporal multiplexing with a QELM, the rank of the readout matrix $G$ [Eqs.
-28 and 30] can serve as an indicator of the training accuracy, as illustrated in Fig.
+For the TFIM variant $H_{\text{TFIM}}^{\text{xx-z}}$ [Eq.
+5b], we obtain a maximum rank of $\mathrm{rank}(G)/3^{n}\simeq 0.22$, RMSE $\epsilon_{\text{train}}\simeq 7\times 10^{-4}$, and forecast horizon $T_{\text{fch}}\simeq 0$.
+When employing temporal multiplexing, the rank of the readout matrix $G$ [Eq.
+28] can serve as an indicator of the training accuracy, as illustrated in Fig.
 10.
 This rank is upper bounded by the number of nonzero Pauli features q{q}, which, for the encoding (9) used here, is ${q}=3^{n}$.
-When $\mathrm{rank}(G)\simeq{q}$, all Pauli features are available at the readout.
-This can, in principle, improve performance when conditioning is controlled and exponential concentration effects can be avoided [51, 45].
+When conditioning is controlled and exponential concentration effects are avoided [53, 48], achieving $\mathrm{rank}(G)\simeq{q}$ makes essentially all Pauli features available to the readout and can thus improve performance.
 Indeed, the QELM in Fig.
 10(b) achieves training error and forecast horizon essentially identical to the ones reported in Table 2 obtained with a complete Pauli observable set.
-On the other hand, as seen in Fig.
-10(a), performance is poor when the readout utilizes only a few effective observables.
+By contrast, Fig.
+10(a) shows that performance remains poor when the readout utilizes only a few effective observables.
 In the present setting, this is a consequence of using a structured Hamiltonian like the TFIM [Eq.
-5a], which spreads observables less effectively over the full Hilbert space (see Fig.
-4), such that $\mathrm{rank}(G)$ is capped by the rank of the PTM $R$.
+5a], which spreads observables less effectively over the full operator space, so that $\mathrm{rank}(G)\simeq\mathrm{rank}(R)\ll q$ [see Eqs.
+31 and 4].
+This effect is most pronounced when using the specific TFIM variant $H_{\text{TFIM}}^{\text{xx-z}}$ [Eq.
+5b] with $Z$-type observables.
+Numerical experiments confirm the close relation (31) between the readout and PTM rank.
 V.3 Interpretability
@@
 […]
+(a)Refer to caption (b)Refer to caption (c)Refer to caption (d)Refer to caption
 […]
-$\boldsymbol{\upphi}(\mathbf{u})=K\mathcal{B}(\mathbf{u}),$ (61)
+[…]
+$\boldsymbol{\upphi}(\mathbf{u})\simeq K\mathcal{B}(\mathbf{u}),$ (57)
 and insert this into the learned model (25),
@@
 […]
-For the dynamical systems considered here, we can simply take $\mathcal{B}=\mathcal{B}_{r}$ to be the set of monomials in $u_{j}$ up to order $r$, of which there are $N={D+r\choose r}$.
-Accordingly, $K$ follows from Taylor-expanding $\boldsymbol{\upphi}(\mathbf{u})$ around some operating point $\bar{\mathbf{u}}$ (usually the data mean).
-As shown in Section IV.4.2, low-order monomials are indeed well constructible from $\boldsymbol{\upphi}$.
+While there are several methods to determine $K$ (such as collocation), we perform here a Taylor expansion of $\boldsymbol{\upphi}(\mathbf{u})$ around some operating point $\bar{\mathbf{u}}$ (data mean).
+This is natural here, since, for the considered polynomial dynamical systems, we can take $\mathcal{B}=\mathcal{B}_{r}$ to be the set of monomials in $u_{j}$ up to order $r$, of which there are $N={D+r\choose r}$.
+Recall that low-order monomials are well constructible from the library $\boldsymbol{\upphi}$ [see Section IV.4.2].
 However, due to the tensor product structure of $\boldsymbol{\upphi}$ [see Eq.
-22] and the fact, that the derivatives of the single-qubit feature vector lie in a 2-dimensional space (vanishing Pauli-$Y$), $K$ typically has not full column rank when $N<q$.
-A consequence of this is that the expression of $\boldsymbol{\upphi}$ in terms of $\mathcal{B}$ is not unique.
+22] and the fact, that the derivatives of the single-qubit feature vector lie in a 2-dimensional space (vanishing Pauli-$Y$), $K$ typically has here not full column rank when $N<q$.
+A consequence of this is that only a $\mathrm{rank}(K)$-dimensional subspace of the library $\mathcal{B}$ is accessible through the encoding features $\boldsymbol{\upphi}$ and that hence some monomials in $\mathcal{B}$ are not independently identifiable.
 Nevertheless, deviations are small for the settings considered here  [^10] Deviations are small for $N<q$: in practice, for $n=D=3$ (thus $q=27$) and monomials of order $r=3$ (thus $N=19$ excluding the constant term), we get $\mathrm{rank}(K)=16$.
-Figure 11 shows the learned flow maps, as approximated in Eq.
-62, resulting from training on the (rescaled) Lorenz-63 and Halvorsen models (no feature engineering in the readout).
+Figure 11 shows the learned flow maps, resulting from Eq.
+58 after training a QELM with amplitude encoding [Eqs.
+8 and 9] on the (rescaled) Lorenz-63 and Halvorsen models (no feature engineering in the readout).
 Comparison to the theoretical expectation [Eq.
-47] reveals good agreement especially for low-order coefficients.
-By contrast, deviations can be significant for higher-order coefficients.
-However, since they are multiplied by increasing powers of $\Delta t$, their influence on the overall training error is small.
+49] reveals good agreement especially for low-order coefficients.
+While deviations can be large for higher-order coefficients, their influence on the overall training error is still small, since they are multiplied by increasing powers of $\Delta t$.
+[…]
 VI Summary
-We have described a memoryless quantum reservoir computer (quantum extreme learning machine, QELM) in the Pauli transfer matrix (PTM) formalism, building upon [33, 25, 43, 46].
+We have employed the Pauli transfer matrix (PTM) formalism to describe a memoryless quantum reservoir computer (quantum extreme learning machine, QELM) with initial-state encoding and data-independent reservoir dynamics.
 Writing states and channels in the Pauli basis (or any other orthogonal operator basis) turns the QELM into a sequence of real linear maps $T_{\mathcal{E}_{j}}$ acting on the vector of Pauli features $\boldsymbol{\upphi}(\mathbf{u})$ representing the encoded state $\rho_{\text{enc}}(\mathbf{u})$.
 These features carry the full dependence of the QELM on the encoded input $\mathbf{u}$ and determine the nonlinearities that the model can possibly express.
-Unitaries appear as orthogonal rotations on the traceless sector of $\boldsymbol{\upphi}(\mathbf{u})$, while general CPTP maps act affinely and typically contract the feature vector.
+Unitaries appear as orthogonal rotations on the traceless sector of $\boldsymbol{\upphi}(\mathbf{u})$, while general CPTP maps act affinely.
 Training a QELM reduces to classical linear regression over the components $\phi_{k}(\mathbf{u})$ and can be framed as a decoding problem of features mixed by quantum channels.
-While the PTM formalism has been previously been applied to QRCs [33, 25, 43, 46], we focused here on its implications for time-series prediction and model identification with QELMs.
-First, this approach allows one to cast the optimization of a QELM as an inverse problem, in which the desired target features imply an ideal encoding scheme and decodability from (preferably low-weight) Pauli features.
-Second, the PTM formalism provides a classical representation of the QELM [93, 94, 95, 47] and thereby renders it directly interpretable as a nonlinear vector (auto-)regression model (or next-generation RC [64]).
-These insights can help to construct reduced-order models in cases where the exact data-generating process is unknown and opens the door to further studies of black-box model identification [83, 96].
-We also discussed the problem of temporal multiplexing and operator spreading [1, 33, 44, 59, 60] within the PTM framework, which can enable a more precise control over these and similar measurement enhancement strategies.
+While the PTM formalism has been previously been applied to QRCs [36, 28, 46, 49], we focused here on its implications for time-series prediction and system identification with QELMs.
+First, this approach allows one to cast optimization of a QELM as an inverse problem: knowing the target features can point to an ideal encoding scheme as well as to constraints on the decodability paths.
+This goes beyond universal approximation theorems [26, 6, 27, 28, 29, 30] and allows one to gain more control over the training performance.
+Second, we have determined a classical and interpretable representation of the QELM in terms of a nonlinear vector (auto-)regression model (or “next-generation” RC [71]).
+The question of classical representations of QML models has been explored in numerous works [108, 109, 110, 50].
+These insights can help to construct reduced-order models in cases where the exact data-generating process is unknown and address the problem of black-box system identification [98, 111].
+Examining temporal multiplexing and operator spreading within the PTM framework demonstrated that greater reservoir randomness is advantageous under restricted measurement schemes.
 Using the Pauli basis is justified by the measurement process, but not the only possibility.
 For future work, it could thus be interesting to study other orthogonal operator sets and assess the generality of the present findings.
-Figure 12: Pauli weight average (a) of Heisenberg-evolved measurement observables can predict the trainig error (b) and forecast horizon (c) of a QELM.
+Figure 13: Pauli weight average (a) of Heisenberg-evolved measurement observables can predict the training error (b) and forecast horizon (c) of a QELM.
 We use a TFIM Hamiltonian (varying coupling parameters $J,h$, fixed $t=6$) with $n=3$ qubits and amplitude encoding (8) (no temporal multiplexing).
-The Pauli weight average is data-independent, while the QELM in (b,c) is trained on the Lorenz-63 model and employs 1-qubit Pauli-$Z$ observables in the readout, enhanced by monomials up to order 2 in order to increase sensitivity to nonlinear terms.
-It has previously been observed [97, 98, 99, 100, 101, 102, 103] that enhancing the quantum nature of a QML model often does not improve its performance on classical datasets.
-The present study adds further explanations of this finding from the viewpoint of learning dynamical systems with QELMs: first, a reservoir unitary mixes features and reversing this potentially detrimental effect requires complex measurement schemes (either high-order Pauli observables or temporal multiplexing), which, in turn, puts more demand on the linear decoder.
+The Pauli weight average is data-independent, while the QELM in (b,c) is trained on the Lorenz-63 model and employs 1-qubit Pauli-$Z$ observables in the readout, enhanced by monomials up to degree 2 to increase sensitivity to nonlinear terms.
+Numerous previous studies noted that simply enhancing the quantum nature of a QML model often does not improve its performance on practical datasets [112, 113, 114, 115, 116, 117, 118].
+The present study adds further explanations of this finding: first, a reservoir unitary mixes features and reversing this potentially detrimental effect requires complex measurement schemes (i.e., high-order Pauli observables, optimized pre-measurement unitaries, or temporal multiplexing schemes).
 We can demonstrate this mechanism directly by comparing the Pauli weight average of single-qubit observables to its training accuracy.
 As illustrated in Fig.
-12 for a minimal QELM without temporal multiplexing, large Pauli weight average strongly correlates with high training error and short forecast horizon.
-In conventional QML models, instead, controlling this potentially detrimental feature mixing must be achieved by expensive training of variational gates.
-Second, optimal training error scaling requires learning of the flow map by a structurally matching feature library [88, 84].
-Feature libraries provided by usual encoding schemes do not achieve this a priori and lead to surrogate models with significant errors (cf.
+13, for such a minimal QELM without temporal multiplexing, large Pauli weight average strongly correlates with high training error and short forecast horizon on a time-series task.
+In conventional QML models, instead, feature mixing must be controlled by training of variational gates.
+Second, learning of the flow map of a dynamical system benefits from a structurally aligned feature library [102, 105].
+If the encoding scheme does not provide this, the resulting surrogate model can exhibit significant errors (cf.
 Fig.
 11).
-The forecast horizon $T_{\text{fch}}$ can be understood here akin to generalization error for static learning tasks: a nonzero $T_{\text{fch}}$ is only possible when the model has learned the actual rule governing the data  [^8] This is in contrast to the one-step training error, which can be decreased even by linear delay features..
-Beside optimizing the measurement process [29], the present work thus points to the encoding as a central turning knob for performance of QELMs and, more generally, QML models (see, e.g., [105, 106] for insights along this line).
-How this surrogate-model argument gets modified when training on datasets without clear ground truths is an interesting avenue for future work [107].
-The $4^{n}$-dimensional space of Pauli features  [^9] Due to symmetries, the actual effective dimension is lower, and typically $\sim 3^{n}$ underlies the exponential frequency content found for QML models that employ rotational encoding [109, 49].
-These and similar properties stemming from the exponential Hilbert space dimensionality point to a possible quantum advantages, but require specific measurement strategies.
-Moreover, simply increasing qubit number $n$ does not generally improve QRC performance, as overfitting and exponential concentration effects can become prevalent [51, 12, 45, 47].
+In the case of chaotic systems, such errors are magnified by the Lyapunov instability and yield short forecast horizons $T_{\text{fch}}$  [^11] This is in contrast to the one-step training error, which can be decreased even by linear delay features..
+Beside optimizing the measurement process [45, 32] and carefully structuring the Hamiltonian [43, 97, 50], the present work thus points to the encoding as a central turning knob for performance of QELMs and, more generally, QML models (see, e.g., [120, 121] for further insights along this line).
+How this surrogate-model argument gets modified when training on datasets with unknown ground truths is an interesting avenue for future work [122].
+The $4^{n}$-dimensional space of Pauli features  [^12] Due to symmetries, the actual effective dimension is lower, and typically $\sim 3^{n}$ can be compared to the exponential frequency content of QML models that employ gate-based rotational encoding [51, 124, 52].
+These and similar properties stemming from the exponential Hilbert space dimensionality point to possible quantum advantages, but require specific measurement strategies.
+Moreover, simply increasing qubit number $n$ does not generally improve QRC performance, as ill-conditioning and exponential concentration effects can become prevalent [53, 12, 48, 50].
 In that regime, more sophisticated reservoir design strategies will be required to steer feature propagation through the quantum layers.
@@
 q{q} Number of nonzero Pauli features (${q}\leq d^{2}$)
-$\mathbf{x}$ Input data ($\mathbf{x}\in\mathbb{R}^{D}$)
+$\mathbf{x},\mathbf{u}$ Input data ($\mathbf{x},\mathbf{u}\in\mathbb{R}^{D}$)
 $y$ Target value ($y\in\mathbb{R}$)
@@
 $M_{k}$ Measurement operators
-$P_{k}$ $n$-qubit Pauli operators ($k=1,\ldots,d^{2}$)
+$P_{k}$ $n$-qubit Pauli operators ($k=0,\ldots,d^{2}-1$)
 $\sigma_{j}$ single-site Pauli matrices ($\mathbb{I}$, $X$, $Y$, $Z$)
@@
 $\mathcal{S}$ Selected observable set ($|\mathcal{S}|={m}$)
+$S$ Selector matrix corresponding to $\mathcal{S}$
 $\boldsymbol{\upphi}(\mathbf{x})$ Pauli feature vector ($\boldsymbol{\upphi}(\mathbf{x})\in\mathbb{R}^{d^{2}}$)
@@
 $\gamma$ Noise parameter
-Table 3: Summary of notation used in this work.
+Table 3: Summary of notation used in this work. A trained quantity is indicated by a hat (e.g., $\hat{\mathbf{w}}$).
 Appendix A Kernel representation
-Training over a complete basis of measurement operators $\{M_{k}\}_{k=1}^{d}$ results in a quantum kernel form of the QELM [29]:
-$f(\mathbf{x})=\sum_{j=1}^{P}\alpha_{j}^{*}K(\mathbf{x},\mathbf{x}_{j})=\mathrm{tr}(M^{*}\rho(\mathbf{x}))$ (63)
-with coefficients $\alpha_{j}$ and the kernel [28]
-$K(\mathbf{x},\mathbf{x}^{\prime})=\mathrm{tr}\big(\rho(\mathbf{x})\rho(\mathbf{x}^{\prime})\big)=\mathrm{tr}\big(\rho_{\text{enc}}(\mathbf{x})\rho_{\text{enc}}(\mathbf{x}^{\prime})\big)$ (64)
-Notably, any data-independent unitary such as $U_{R}$ in (4) drops out from Eq.
-63, implying that such a QELM is only dependent on the encoding state $\rho_{\text{enc}}(\mathbf{x})$.
-Furthermore, Eq.
-63 allows to identify an optimal measurement operator:
+Training over a complete basis of measurement operators $\{M_{k}\}_{k=1}^{d^{2}}$ results in a quantum kernel form of the QELM [32]:
+$f(\mathbf{x})=\sum_{j=1}^{P}\alpha_{j}^{*}K(\mathbf{x},\mathbf{x}_{j})=\mathrm{tr}(M^{*}\rho(\mathbf{x})),\qquad K(\mathbf{x},\mathbf{x}^{\prime})=\mathrm{tr}\big(\rho(\mathbf{x})\rho(\mathbf{x}^{\prime})\big)$ (59)
+For the least-squares loss, the optimal parameters have the explicit representation $\boldsymbol{\alpha}^{*}=(\mathbf{K}+\lambda\mathbbm{1})^{-1}\mathbf{y}$, with the Gram matrix $K_{ij}=K(\mathbf{x}_{i},\mathbf{x}_{j})$ and the sample vector $\mathbf{y}\in\mathbb{R}^{P}$.
+If the reservoir consists only of a data-independent unitary, it drops out from Eq.
+59,
+[…]
+showing that such a QELM only depends on the encoding state $\rho_{\text{enc}}(\mathbf{x})$. An optimal measurement operator can be identified from Eq. 59 as
 $M^{*}=\sum_{j=1}^{P}\alpha_{j}^{*}\rho(\mathbf{x}_{j}).$ (61)
-For the least-squares loss, the optimal parameters have the explicit representation $\boldsymbol{\alpha}^{*}=(\mathbf{K}+\lambda\mathbbm{1})^{-1}\mathbf{y}$, with the Gram matrix $K_{ij}=K(\mathbf{x}_{i},\mathbf{x}_{j})$.
+Compared to the equivalent primal formulation [Eq. 1], this representation is computationally more efficient when $P\ll d^{2}$.
 Appendix B Exact feature isolation criterion: null-space condition
-Equation 38 is equivalent to requiring $\mathbf{w}R=\omega e_{r}^{\top}$ with some scalar weight $\omega$, i.e., $e_{r}^{\top}\in\mathbb{R}^{q}$ belongs to the row-space of the effective PTM $R\in\mathbb{R}^{{m}\times{q}}$.
+Equation 39 is equivalent to requiring $\mathbf{w}^{\top}R=\omega e_{r}^{\top}$ with some scalar weight $\omega$, i.e., $e_{r}^{\top}\in\mathbb{R}^{q}$ belongs to the row-space of the effective PTM $R\in\mathbb{R}^{{m}\times{q}}$.
 This, in turn, is equivalent to
@@
 i.e., the $r$-th component of every (basis) vector of $\mathrm{null}(R)$ must be zero. If Eq. 62 is met, the solution for $\mathbf{w}$ is provided by the pseudoinverse $R^{+}$ via
-$\mathbf{w}=\omega\mathbf{e}_{r}^{\top}R^{+}$ (67)
+$\mathbf{w}^{\top}=\omega\mathbf{e}_{r}^{\top}R^{+}$ (63)
 up to a term $\propto h(\mathbbm{1}_{m}-RR^{+})$ with an arbitrary vector $h\in\mathbb{R}^{1\times{m}}$ accounting for the non-uniqueness of the solution if $R$ does not have full row rank.
@@
 Panel (a) shows the total number of isolatable features, while panel (b) shows the number of isolatable 1-qubit features.
-In contrast to statistical decodability [cf.
-Appendix C], isolatability is an exact geometric criterion independent of the data.
+In contrast to statistical decodability (cf.
+Appendix C), isolatability is an exact geometric criterion independent of the data.
 Consider the case where $R\in\mathbb{R}^{{m}\times{q}}$ is dense and quasi-random, arising, e.g., for a Haar-random unitary $U$.
 For m<q{m}<{q}, the null space is necessarily nonzero and in random orientation.
-Accordingly, the probability that all vectors of $\mathrm{null}(R)$ have the same component vanish is essentially zero, such that no features are isolatable.
+Accordingly, the probability that the same component vanishes for all vectors of $\mathrm{null}(R)$ is essentially zero, such that no features are isolatable.
 Once ${m}\geq{q}$, $R$ has full column rank, hence $\mathrm{null}(R)=0$ and the null-space condition in Eq.
 62 is trivially fulfilled, implying that all q{q} features become instantly isolatable for a random $R$.
-In this case one may define $W=R^{+}=R^{\top}$ and assign $\mathbf{w}=\Pi_{r}W$ with a $\Pi_{r}$ that projects onto the $r$-th coordinate.
-This $W$ fulfills the slightly more general condition of finding a matrix $W\in\mathbb{R}^{{m}\times{m}}$ such that $WR=[\mathbbm{1}_{{m}},0_{{q}-{m}}]$, thereby isolating all m{m} features.
-[^10] Writing the PTM as $R=[R_{1},R_{2}]$ with $R_{2}\in\mathbb{R}^{{m}\times({q}-{m})}$, this requires in general $R_{2}=0$, which is not fulfilled for typical (random) unitaries except if symmetries are present and the basis is ordered accordingly (e.g., the parity symmetry of the TFIM would allow a diagonal block structure of $R$ with m=q/2{m}={q}/2).
-For a Clifford unitary, which maps between elements of the Pauli group, $\mathrm{row}(R)$ simply consists of all $\mathbf{e}_{j}$, hence $\mathrm{rank}R$ features are isolatable even if $R$ is wide.
+In this case one may define $W=R^{+}=(R^{\top}R)^{-1}R^{\top}$ and assign $\mathbf{w}^{\top}=\mathbf{e}_{r}^{\top}W$.
+For a Clifford unitary, which maps between elements of the Pauli group, $\mathrm{row}(R)$ simply consists of a subset of signed standard basis vectors, hence $\mathrm{rank}(R)$ features are isolatable even if $R$ is wide.
 For unitaries based on structured Hamiltonians or at small evolution times, one may expect Eq.
@@
 This is confirmed numerically in Fig.
-13, where the behavior of the feature isolation criterion (66) is illustrated for various unitaries as a function of the number of observables m{m} (rows of $R$).
+14, where the behavior of the feature isolation criterion (62) is illustrated for various unitaries as a function of the number of observables m{m} (rows of $R$).
 Appendix C Decodability scores and coefficient of determination
 The decodability score $\gamma_{r}^{2}$ [Eq.
-40] tells to which extent a feature can be geometrically reconstructed from the readout.
+41] tells to which extent a feature can be geometrically reconstructed from the readout.
 It disregards statistical correlations and effects of noise, i.e., features with small singular values in the effective PTM $R$ could be considered decodable, although it would lead to huge weight norms.
-We present a refined $\mathcal{R}^{2}$-score below, but reserve a more detailed discussion of noise/regularization effects for a future work.
-Consider the regression solution in Section IV.1, assume all quantities are centered and define the target alignment vector
-$\mathbf{b}_{k}:=\mathbb{E}\!\left[\boldsymbol{\phi}(\mathbf{u})y_{k}(\mathbf{u})\right]\in\mathbb{R}^{q}.$ (68)
+A straightforward way to limit noise amplification is to replace the pseudoinverse by its Tikhonov regularized form
+$\gamma_{r,\lambda}^{2}=(R_{\lambda}^{+}R)_{rr}=\left[(R^{\top}R+\lambda\mathbbm{1})^{-1}R^{\top}R\right]_{rr}=\sum_{i=1}^{s}\frac{\sigma_{i}^{2}}{\sigma_{i}^{2}+\lambda}v_{ri}^{2},\qquad\gamma_{r,\lambda\to 0}^{2}\to\gamma_{r}^{2}$ (64)
+where $\sigma_{i=1,\ldots,s}$ are the singular values of $R$ and $\mathbf{v}_{i}$ the right-singular vectors.
+Directions with $\sigma_{i}^{2}\ll\lambda$ contribute little to $\gamma_{r,\lambda}^{2}$ and are thus considered undecodable.
+To account for data correlations within a refined $\mathcal{R}^{2}$-score, consider the regression solution in Section IV.1, assume all quantities are centered, and define the target alignment vector
+$\mathbf{b}_{k}:=\mathbb{E}\!\left[\boldsymbol{\upphi}(\mathbf{u})y_{k}(\mathbf{u})\right]\in\mathbb{R}^{q}.$ (65)
 Using Eq. 26, the mean-squared error satisfies
-$\mathbb{E}_{\mathbf{u}}\bigl[(y_{k}-f)^{2}\bigr]=\mathrm{Var}_{\mathbf{u}}(y_{k})-(R\mathbf{b}_{k})^{\top}\mathbf{w}^{*}.$ (69)
-Plugging this into the definition in Eq. 43 gives the general expression for the coefficient of determination
-$\mathcal{R}^{2}(k)=\frac{\mathbf{b}_{k}^{\top}R^{\top}(RCR^{\top})^{+}R\mathbf{b}_{k}}{\mathrm{Var}(y_{k})}.$ (70)
-All quantities herein are to be averaged over the input distribution of $\mathbf{u}$.
-Notably, it is possible to have $\gamma_{k}^{2}=0$ (feature direction not present in row-space of $R$), yet still $\mathcal{R}^{2}(k)>0$ if $\phi_{r}$ is statistically correlated with other $\phi_{j}$ that are present.
-Hence, the above $\mathcal{R}^{2}$ generally measures predictability from correlations and not just isolatability of a feature.
-The coefficient of determination $\mathcal{R}^{2}(k)$ can be related to the decodability score $\gamma_{r}^{2}$ [Eq.
-40]: In the special where the targets are $y_{k}(\mathbf{u})=\phi_{k}(\mathbf{u})$, and under the additional assumption of orthonormality,
-$\mathbb{E}[y_{k}y_{k^{\prime}}]=\delta_{kk^{\prime}}\implies C=\mathbbm{1}_{q},\quad\mathbf{b}_{k}=\mathbf{e}_{k},\quad\mathrm{Var}(y_{k})=1$ (71)
-Eq. 70 reduces to the decodability score
-$\mathcal{R}^{2}(k)=\mathbf{b}_{k}^{\top}R^{\top}(RR^{\top})^{+}R\mathbf{b}_{k}=\mathbf{e}_{k}^{\top}(R^{+}R)\mathbf{e}_{k}=\gamma_{k}^{2}.$ (72)
-In the generic non-orthogonal case, the exact relation breaks down and $\mathcal{R}^{2}$ depends on the full projector $R^{+}R$ and not just on its diagonal elements.
+$\mathbb{E}_{\mathbf{u}}\bigl[(y_{k}-f)^{2}\bigr]=\mathrm{Var}_{\mathbf{u}}(y_{k})-(R\mathbf{b}_{k})^{\top}\hat{\mathbf{w}}$. (66)
+Plugging this into the definition in Eq. 44 gives the expression for the coefficient of determination (in the unregularized case $\lambda=0$)
+$\mathcal{R}^{2}[y_{k}]=\frac{\mathbf{b}_{k}^{\top}R^{\top}(RCR^{\top})^{+}R\mathbf{b}_{k}}{\mathrm{Var}(y_{k})},$ (67)
+where $C=\mathbb{E}[\boldsymbol{\upphi}\boldsymbol{\upphi}^{\top}]$ is the feature covariance matrix. All quantities herein are to be averaged over the input distribution of $\mathbf{u}$.
+In the special where the targets are the Pauli features, $y_{r}(\mathbf{u})=\phi_{r}(\mathbf{u})$, and under the additional assumption of orthonormality (whitened features),
+$\mathbb{E}[y_{r}y_{r^{\prime}}]=\delta_{rr^{\prime}}\implies C=\mathbbm{1}_{q},\quad\mathbf{b}_{r}=\mathbf{e}_{r},\quad\mathrm{Var}(y_{r})=1,$ (68)
+the coefficient of determination $\mathcal{R}^{2}[y_{r}]$ in Eq. 67 reduces to the decodability score
+$\mathcal{R}^{2}[y_{r}]=\mathbf{b}_{r}^{\top}R^{\top}(RR^{\top})^{+}R\mathbf{b}_{r}=\mathbf{e}_{r}^{\top}(R^{+}R)\mathbf{e}_{r}=\gamma_{r}^{2}.$ (69)
+In the generic non-orthogonal case, this simple relation breaks down and $\mathcal{R}^{2}$ depends also on the covariance $C$ and target-feature alignment.
 Then, $\mathcal{R}^{2}$ measures the extent to which the target lies in the decodable subspace.
+As a consequence, it is then possible to have $\gamma_{r}^{2}=0$ (feature direction not present in row-space of $R$), yet still $\mathcal{R}^{2}[y_{r}]>0$ if $\phi_{r}$ is statistically correlated with other $\phi_{j}$ that are present.
+Hence, $\mathcal{R}^{2}$ in Eq.
+67 generally measures predictability from correlations and not just isolatability of a feature.
 Appendix D Pauli-Transfer Matrix Formalism: Further Details
@@
 Consider a separable reservoir Hamiltonian $H_{R}=\sum_{j=1}^{n}h_{R}^{(j)}$, such that the reservoir unitary factorizes as $U(t)=\bigotimes_{j=1}^{n}U_{j}(t)$.
-Such a structure emerges, e.g., in the limit of large transverse field $h/J\gg 1$ of the TFIM (5) This implies a Kronecker-product structure of the unitary PTM (73) (and similarly for any other separable quantum channel),
+Such a structure emerges, e.g., for a TFIM with $h/J\to\infty$ [Eq.
+5].
+This implies a Kronecker-product structure of the unitary PTM (70) (and similarly for any other product quantum channel),
 $T_{U}(t)=V^{(1)}(t)\otimes\cdots\otimes V^{(n)}(t),$ (72)
@@
 For example, for a local $Z$-field $h_{k}=Z$,
-$U_{k}(t)^{\dagger}XU_{k}(t)=\cos(2t)\,X+\sin(2t)\,Y,\qquad U_{k}(t)^{\dagger}YU_{k}(t)=\cos(2t)\,Y-\sin(2t)\,X,$ (78)
-so that $\phi_{y}^{{}^{\prime}(k)}(u_{k})=-\sin(2t)\,\phi_{x}^{(k)}(u_{k})$ and, correspondingly, $\phi^{\prime}_{\ldots y_{k}\ldots}(\mathbf{x})=-\sin(2t)\,\phi_{\ldots x_{k}\ldots}(\mathbf{x})$, while $Z$-features remain invariant.
+$U_{k}(t)^{\dagger}XU_{k}(t)=\cos(2t)\,X-\sin(2t)\,Y,\qquad U_{k}(t)^{\dagger}YU_{k}(t)=\cos(2t)\,Y+\sin(2t)\,X,$ (75)
+so that $\phi_{y}^{{}^{\prime}(k)}(u_{k})=\sin(2t)\,\phi_{x}^{(k)}(u_{k})$ and, correspondingly, $\phi^{\prime}_{\ldots y_{k}\ldots}(\mathbf{x})=\sin(2t)\,\phi_{\ldots x_{k}\ldots}(\mathbf{x})$, while $Z$-features remain invariant.
 This shows how unitaries act to make specific features accessible to the readout when the latter uses a restricted set of observables.
@@
 $T_{\mathcal{E}}=\begin{pmatrix}1&\mathbf{0}^{\top}\\ \mathbf{t}&\Omega\end{pmatrix},\qquad\mathbf{s}^{\prime}=\mathbf{t}+\Omega\mathbf{s},$ (77)
-where $\mathbf{t}=\mathbf{0}$ for unital channels and $\Omega$ is a contracting map (singular values $\leq 1$).
-Such a map can improve numerical conditioning by damping high-weight Pauli components that are often noise sensitive.
+where $\mathbf{t}=\mathbf{0}$ for unital channels.
+For a single qubit, $\Omega$ is a contracting map (singular values $\leq 1$).
+Such a map can improve numerical robustness by damping high-weight Pauli components that are often noise sensitive.
 Non-unitality produces a translation $\mathbf{t}$ of the Bloch part, e.g., a maximally mixed state would be shifted away from the center of the Bloch sphere.
-This becomes a learnable bias once combined with the identity component of an observable.
 Consider the following single-qubit PTMs for non-unitary noise channels [126, 127, 128]:
@@
     $T_{\mathrm{dep}}(\lambda)=\mathrm{diag}(1,\lambda,\lambda,\lambda).$ (78)
-    For an $n$-qubit isotropic depolarizing channel on dimension $d=2^{n}$, $T_{\mathrm{dep}}^{(n)}(\lambda)=\mathrm{diag}\big(1,\lambda I_{d^{2}-1}\big)$.
+    For a global isotropic depolarizing channel on dimension $d=2^{n}$, $T_{\mathrm{dep}}^{(n)}(\lambda)=\mathrm{diag}\big(1,\lambda I_{d^{2}-1}\big)$.
   • $Z$-dephasing with probability $p$:
@@
     $T_{\mathrm{AD}}(\gamma)=\begin{pmatrix}1&0&0&0\\ 0&\sqrt{1-\gamma}&0&0\\ 0&0&\sqrt{1-\gamma}&0\\ z_{*}\gamma&0&0&1-\gamma\end{pmatrix}$. (80)
-Modeling $n$-qubit noise as tensor products of single-qubit channels $\mathcal{E}^{(n)}=\bigotimes_{i=1}^{n}\mathcal{E}^{(i)}$ is often used as a baseline in QRC models [115, 18, 21, 115, 46].
+Modeling $n$-qubit noise as tensor products of single-qubit channels $\mathcal{E}^{(n)}=\bigotimes_{i=1}^{n}\mathcal{E}^{(i)}$ is often used as a baseline in QRC models [129, 19, 22, 49].
 Since the PTMs (78) and (79) are Pauli-diagonal, they do not lead to Pauli spreading and hence do not enlarge the accessible feature subspace.
 Amplitude damping [Eq.
-83] produces Pauli-I/Z-mixing, which causes operator weight decay.
+80] produces Pauli-I/Z-mixing, which causes operator weight decay.
 A particularly transparent characterization emerges if the noise channel is approximately diagonal.
 For instance, for the $n$-qubit depolarizing channel $\mathcal{D}^{(n)}=\otimes_{i=1}^{n}\mathcal{D}_{\lambda}$ [see Eq.
-81], any Pauli string $P$ of weight $\nu(P)$ is an eigenoperator with
-$\mathcal{D}^{(n)}(P)=\lambda^{\nu(P)}P$ (84)
+78], any Pauli string $P$ of weight $\nu(P)$ is an eigenoperator with
+$\mathcal{D}^{(n)}(P)=\lambda^{\nu(P)}P.$ (81)
 Accordingly, high-weight Pauli strings are exponentially suppressed by near-diagonal noise channels.
 In conjunction with unitary-induced operator spreading, this can increase information loss under finite sampling.
-On the other hand, amplitude damping noise was observed to improve QRC performance as the associated superoperator spreads the state’s probability weight into more Pauli directions, thereby enriching the feature map [18].
-This has been related to the ’coherence influx’ present for amplitude damping but not depolarization [43].
+On the other hand, amplitude damping noise is often observed to improve QRC performance as the associated superoperator spreads the state’s probability weight into more Pauli directions, thereby enriching the feature map [19].
+This has been related to the ‘coherence influx’ present for amplitude damping but not depolarization [46].